A Complete Bibliography of Publications in the Bulletin for the History of Chemistry

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

20 August 2024 Version 1.17

Title word cross-reference

$15.00 [571].

1 [631]. 100 [483]. 13th [385]. 1493-1541 [770]. 1676 [762, 473]. 16th [412]. 17a [206]. 17th [145]. 17th-century [145]. 1860s [501]. 1870s [501]. 1893 [31]. 18th [3, 23, 40, 51]. 1914 [236]. 1918 [583]. 1930S [491]. 1970* [317]. 1974 [511]. 1987 [4, 35]. 1988 [27]. 1989 [38, 60]. 1990 [93]. 1991 [132]. 1992 [157]. 1993 [169, 770]. 1994 [190]. 1995 [203]. 1996 [212]. 1997 [219]. 1998 [227]. 1999 [247]. 19th [218, 85, 299, 286, 412, 281, 147, 703]. 19th-Century [299, 286, 281, 147].

2 [646]. 2000 [270]. 2001 [280]. 2002 [278, 298]. 2003 [313]. 2004 [329]. 2005 [348]. 2006 [363]. 2007 [360, 374, 381]. 2008 [399]. 2009 [397, 417]. 20th [41, 444]. 250th [19].

300th [441].

5-Triphenyltetrazolium [481]. 500th [770].

978-0-9916503-0-9 [571].

A. [590]. Aaron [260, 261, 263, 262, 264]. Abridged [546]. Absorption [323]. Academia [223]. Academic [348, 454]. Acceptance [311]. According [467]. account [732]. Accum [613]. Acetylene [366]. Achievement [76, 199, 748]. achievements [744, 226]. Acid [137, 570, 411, 533, 689]. ACS [633, 656, 31, 62, 318]. Act [490, 486]. Action [298, 234]. Ada [252]. Adam [459]. Adams [341]. Added [287, 700]. Addendum [562]. Address [190, 157, 219, 203, 4, 169, 27, 132, 348, 313, 212, 329, 417, 227, 363, 247, 298, 270, 93, 60]. Adept [236, 736]. adjuste [99]. Ado [450, 705]. Adolphe [192]. Advances [385]. advent [108]. Adventure [461, 722, 106]. Aerations [494]. Affinities [227]. ’Afinity [418]. African [485, 694, 456, 647]. After [182, 579, 481]. Age [451, 691]. Agency [490]. Agent [413]. Agricola [132]. Agriculture [491]. Ahead [375]. Ain’t [357]. Airs [523]. Alberta [85]. Alchemical [236, 306, 736, 738]. Alchemy [473, 356, 244, 145, 467, 737, 762, 507]. Alchymia [244]. Aldersey [466]. Aldersey-Williams [466]. Aleksandr [534, 67, 172, 433]. Alexander [587, 302, 602, 612, 758, 509]. Alfred [199, 384, 588, 494, 569]. Alkali [293]. Alkaline [293]. Alkylidenehydrazines [547, 548]. Alkyls [285]. Almer [204]. almost [62]. Along [581]. Alumina [171]. Alyea [104]. am [211, 770]. amateur [106]. America [473, 762, 295, 237, 195, 213, 86, 22]. American [110, 485, 647, 5, 271, 287, 694, 456, 85, 341, 299, 20, 95, 455, 503, 767, 730, 174, 268, 148, 196, 213, 443, 524, 106, 35]. Americium [393]. Ami [232]. Amines [186]. Ammonia [334]. among [734, 147, 106]. Ana [484]. Anabuse [392]. Analysis [228, 99, 268, 274, 350, 714, 148, 72]. Andres [302]. Anglo [455]. Anglo-American [455]. Aniline [107]. Anna [205, 436]. Anne [461]. Anne-Marie [461]. Annie [197]. Anniversaries [367]. Anniversary [441, 19]. Annotated [226, 750]. antecedents [84]. Antiquity [447, 745]. Antozone [404]. Apparatus [226, 87, 578, 54, 750, 53]. application [72, 134]. Applications [156]. Approach [526, 510, 698, 508, 710]. Approaches [679, 681, 664]. approche [698, 508]. Aproximacion [99]. Arbuzov [590]. arcane [12]. Archaeological [659]. Archaeology [218, 723]. Archeological [11]. Argand [232]. Arkel [149]. Arndt [303]. Arrhenius [99, 696, 382, 751]. Arrhenius. [211]. Arsenic [466, 383, 687]. Art [125, 233, 104]. Arthur [275, 309, 401]. Article [641]. Artifacts [87, 660, 21, 435, 436, 6, 74]. artist [107]. Aspiring [236, 736]. Assaying [174]. Atomic [641, 614, 642, 471, 20]. Atomism [165]. Atoms [464, 709, 605, 464]. Atoxyl [222]. Atti [685, 137]. Attitudes [137, 742]. audience [701]. Audiences [657]. August [407, 409]. Augustus [249]. Auroral [430]. Aurum [499]. Author [79, 201, 258, 245, 346, 437, 83, 71, 64]. Authors [141, 155, 388, 398, 497, 514, 530, 543, 559, 573, 598, 584, 595, 608, 619, 634, 650]. autobiographical [732]. automated [108]. Avowed [196]. Award [278, 280, 219, 203, 169, 348, 313, 212, 329, 498, 227, 363, 247, 405, 298, 270, 9, 215, 328, 343, 360, 397, 190, 9, 18, 29, 42, 365, 399, 611, 157, 381, 618]. Award-Winning [405]. Awarded [280]. Awards [139, 153, 11, 24, 46, 68, 101, 362, 630].

B [571, 629, 505]. B. [349]. Babo [91]. Bache [162]. Back [370, 660, 635, 651, 675]. Bacon [306]. Baekeland [483, 768]. Bailey [243]. Baker [552, 480, 136]. Baking [494, 578]. Balance [84]. Baltimore [283]. Baptiste [83]. Barnes [537]. barrier [40]. Bartlett [206]. Bascom [76, 261, 718]. Base [186, 502, 96]. Base-Promoted [502]. Basel [770]. Basis [719]. Baumeister [211]. Baur [549]. Bayer [171]. Be [518, 223, 232, 350, 54]. Became [525]. Becker [84]. Beecham [239]. Beehives [91]. Before [20, 442, 107, 701, 108]. Beginnings [285, 184]. beguiled [33]. Beitrag [734]. Beiträge [770]. Benedict [139]. Benfey [143]. Benjamin [443, 136]. Benzene [133, 309]. Benzidine [188]. Bernard [342]. Berolinensis [441]. Bertereau [467]. Bertheim [222]. Better [453]. Between [463, 717, 73, 399, 323, 83, 764]. Beyond [491, 393, 664, 709, 768, 464, 483]. Bibles [120]. Bibliography [43, 66, 98, 226, 750]. bicentennial [86, 109]. Biochemical [137, 706]. Biochemistry [377, 616, 410]. Biocolloidal [628]. Biogenesis [628]. Biographers [116]. Biographical [58, 130, 143]. Biographies [630]. Biography [415, 365, 758, 67, 509]. Biology [377]. Biomolecular [628]. biotechnology [689, 137]. Bird [494]. Birth [451, 268, 691]. birthday [770]. Bischoff [551]. Black [417, 293, 159]. Blanc [209]. Blazer [378]. Bleaching [293, 413]. blinds [63]. Blood [344]. Blowpipe [40, 273]. Blue [389, 414, 441, 500, 562]. Blum [451]. Bobierre [192]. Bochard [40]. Bodies [642]. Body [523]. Bohning [635]. Bolton [635]. Bombastus [770]. Bonaparte [234]. Bond [149, 298, 191]. Bond-Type [149]. Bonding [70]. Bones [34, 19]. Book [10, 22, 35, 44, 76, 67, 110, 99, 151, 137, 175, 211, 218, 226, 236, 246, 257, 269, 279, 291, 297, 304, 312, 319, 327, 336, 345, 354, 361, 368, 380, 387, 396, 406, 416, 424, 434, 459, 461, 464, 450, 451, 466, 447, 449, 462, 460, 463, 465, 485, 484, 472, 474, 476, 483, 473, 477, 496, 495, 506, 521, 540, 555, 571, 572, 601, 622, 623, 624, 638, 654, 511, 594, 607, 510, 631, 646, 648, 649, 448, 593, 618, 583, 507, 632, 513, 512, 509, 508]. book [86]. Books [303, 28, 39, 55]. Border [399]. Borodin [67, 509, 758]. Bosch [432]. both [107]. Böttger [420]. bourgeoisie [50]. Boyle [236, 463, 717, 736]. Bragg [110, 755]. brainchild [135]. Breaking [40]. Breakthrough [589, 498]. Breakthroughs [326]. Bridging [491]. Brief [238, 48, 102]. Bring [660]. Britain [714, 657, 494, 99]. British [602, 221, 311, 412, 517, 94, 410]. Brock [477]. Brotherhood [254]. Brothers [220]. Brown [485]. Bryant [308]. Buckingham [189, 289]. Bull [211]. Bulletin [369, 426, 531, 544, 560, 574, 599, 585, 596, 609, 620, 636, 652, 676]. Bunsen [468]. Burr [169]. Butlerov [534, 74]. Byzantium [745, 447].

C [476, 483, 448, 401]. C. [180, 178, 181]. Call [215, 278, 328, 343, 360, 397]. Cambodia [324]. Can [660, 350, 34, 3, 23]. Canada [19]. candle [700, 701, 702, 474]. Cane [7]. Capillary [333, 315]. Capsule [606]. Carbide [580]. Carbohydrate [395]. Carbon [494, 578, 7]. Career [565, 502, 83, 61]. Carey [301, 603]. Carl [432]. Carmen [464]. Carotene [403]. Carothers [211, 715]. Carrington [635]. Case [477, 693, 290, 119, 538, 254, 145]. Cases [535, 534]. Caspar [414]. Catalyst [492]. Catalytic [547, 548, 204]. Catholic [300]. Cation [256]. Caught [51]. Caveman [110, 226, 744]. Cavendish [226]. Celebrated [56]. Celebrating [456, 86, 109]. Celebration [522, 104]. Celebrity [270]. Cell [249]. Cells [549]. Centennial [656, 173]. Center [31]. Central [238]. centuries [7, 625]. Century [99, 218, 613, 493, 299, 95, 286, 412, 281, 502, 5, 3, 23, 135, 85, 703, 40, 714, 41, 33, 105, 145, 738, 147, 742, 51, 137, 236, 468, 494, 385, 444, 467]. Certain [535, 534]. Certainly [227]. Chainomatic [84]. Chair [11, 24, 37, 68, 112, 353]. Chairs [439]. Chandler [21]. Change [227]. Changing [657, 236, 577, 625, 738]. Chaos [219]. Chapter [137, 689, 133]. Characters [522]. Charles [349]. Charlotte [163]. Chastelet [467]. chauvinism [147]. cheap [106]. Check [130]. Checklist [58]. Chem [211, 206]. Chemical [22, 35, 90, 137, 218, 326, 449, 474, 571, 572, 87, 62, 271, 287, 135, 203, 28, 39, 55, 316, 96, 295, 335, 21, 594, 501, 190, 125, 409, 457, 470, 565, 6, 282, 568, 212, 724, 577, 417, 498, 318, 605, 364, 413, 533, 298, 458, 355, 538, 385, 126, 196, 350, 567, 157, 661, 122, 234, 86, 134, 338, 74, 686, 692, 56, 700, 701, 702, 703, 706, 27, 38, 105, 728, 729, 767, 734, 70, 148, 72, 106, 136, 57, 766, 492, 175]. Chemicals [237, 524]. Chemie [44, 211, 726, 236]. Chemiker [236]. Chemische [44]. chemischen [734]. Chemist [67, 110, 226, 236, 549, 613, 683, 85, 170, 197, 124, 519, 602, 583, 645, 386, 247, 518, 148, 200, 213, 308, 210, 158, 5, 83, 50, 41, 721, 744, 761, 110]. Chemist-Composers [386]. Chemist-Historians [683]. Chemistries [97, 17, 86, 64, 109, 136, 167, 193]. Chemistry [657, 22, 35, 76, 160, 175, 211, 236, 464, 465, 484, 556, 552, 627, 679, 670, 682, 684, 475, 526, 180, 690, 493, 400, 693, 303, 660, 666, 288, 179, 697, 511, 703, 491, 169, 375, 607, 206, 510, 658, 606, 478, 479, 428, 299, 718, 119, 57, 263, 348, 445, 442, 565, 671, 719, 95, 194, 372, 468, 282, 520, 313, 430, 320, 525, 536, 329, 618, 310, 577, 285, 296, 590, 639, 662, 490, 286, 453, 300, 318, 363, 731, 669, 365, 659, 293, 663, 522, 630, 418, 165, 580, 664, 412, 454, 444]. Chemistry [517, 665, 281, 184, 147, 270, 231, 399, 93, 350, 443, 749, 373, 582, 672, 317, 513, 754, 366, 381, 489, 668, 109, 167, 488, 698, 508, 5, 3, 43, 144, 4, 708, 709, 710, 711, 82, 20, 720, 721, 723, 726, 107, 735, 738, 73, 741, 743, 747, 49, 753, 17, 86, 35, 76, 137, 211, 218, 656, 640, 718, 769, 175, 218, 236, 484, 477, 631, 646, 10, 587, 512, 64, 719]. Chemists [35, 551, 627, 370, 456, 205, 576, 227, 453, 405, 221, 311, 591, 667, 358, 647, 401, 34, 694, 33, 767, 147, 51, 31, 485]. Chemisty [373]. Chemotherapy [330, 222, 358]. Chemurgy [491]. chimica [685, 137]. chimie [698, 508, 39, 55]. Chimique [712, 28]. Chinese [8]. Chittick [578]. Chloride [481]. Chlorination [309]. Chlorine [379, 413]. Chlorofluorocarbon [357]. Christlieb [230]. Christopher [189, 248]. Chuit [99, 704]. chymiae [735]. Chymistry [230, 104]. Circumstances [226, 351, 744]. Citation [498]. Citations [326]. clamp [102]. Clark [95]. Clarke [127, 273]. Classes [489]. Classic [457]. Classical [632]. Classification [229, 276]. Clear [379]. Clever [505]. Clinical [231]. Club [519]. CO [626]. Cobalt [186, 199]. Coerulei [441]. Coleridge [339]. Collection [564, 482]. College [288, 178, 492, 288, 335]. Colleges [491, 489]. Colloidal [377]. Colonial [96]. Colorants [107]. Colorimetry [267]. Columbia [21]. Column [436]. Combustion [93]. come [69, 81, 92, 103]. Comfortable [675]. Coming [668]. Comment [527, 570, 526, 510]. Commentary [277, 286, 165, 276, 582]. comments [75]. Commercial [204]. Commode [185]. Common [576]. Communicating [372, 488]. Communication [505]. Communist [196]. community [135, 70]. Company [241]. Compelling [553]. Complex [464, 32, 709]. Complexes [199]. Component [294]. Composer [509, 758]. Composers [386]. Compounds [605]. compromise [62]. Computational [665]. Conant [308]. Concept [282, 294, 181]. concepts [729, 137]. Concerning [45]. concluding [61]. Conducting [539]. Conductive [504]. Conferences [394]. Confidence [488]. Congress [31]. Connection [216]. Connections [340]. connoisseurs [17]. Conservation [146, 93, 52]. Conservative [172]. Consideration [516]. considered [50]. Conspiracy [211, 743]. Constitutional [726]. Construction [472, 733]. Consultant [158]. Consulting [645]. consummate [53]. Contemporary [172, 668]. Content [577]. Context [526, 510, 119, 710]. Contextualizing [667]. Continental [62]. continues [98]. continuing [9, 18, 29, 42]. Contraband [479]. Contrasting [468]. Contrasts [706, 212, 137]. Contribution [457, 516, 458, 734]. Contributions [243, 614, 263, 286, 267, 391, 644, 281, 274, 443, 275, 770]. Control [515, 108]. Controversies [377]. Controversy [211, 248, 273, 256, 61, 743]. convegno [685, 137]. Conversations [577, 587]. Cooke [442]. Cooking [493]. Cooper [339]. Cope [401]. copper [45]. Corporate [661]. Corpuscular [145]. Correspondence [587, 294]. Cosmochemistry [549]. Course [552, 206, 373, 700, 701]. Courtois [342]. Cover [678]. Cracking [204]. Cradle [607]. Creation [473, 762, 649]. Credit [176]. credits [678]. Criteria [265]. Critique [276]. Croatian [382, 538]. Crookes [135]. Crossovers [207]. Crucial [391, 422]. Crystallization [31]. Crystals [249]. Cultural [466, 687]. Culture [473, 503, 270, 762]. cum [735]. Cuprene [581]. Curie [460, 739, 740, 196]. curiosities [134]. Curiosity [581]. curious [83]. Curriculum [625]. Cutting [286, 281]. Cutting-Edge [286, 281].

Cycle [240]. Cyrus [274].

D [206, 231]. D. [255, 282]. Dalton [226, 464, 709, 750]. Dangers [383]. Daniel [127]. darstellung [726, 44]. Darzens [168]. Dasein [734]. Data [290]. David [450]. Davies [334]. Davis [241]. Dawn [233]. Days [300]. deals [61]. Debate [295]. Deborah [451]. decade [18, 29, 42]. December [770]. Decline [94]. Decomposition [547, 548, 344, 502]. Decorating [499]. Découverte [236]. Deed [400]. Deep [329]. Defense [295, 505]. del [685, 137, 302, 7]. Delhuyar [220]. Deliberations [505]. delivered [701]. della [685, 137]. demise [62]. Demonstrate [104]. demonstration [3, 104]. Denison [445]. density [7]. Denunciation [256]. Deoxygenation [546]. Department [76, 288, 697, 511, 718, 178]. Derivatives [391]. describes [28, 39, 55]. Describing [604]. Deserves [518]. Designation [256]. designer [53]. Designs [333]. Desk [78, 1, 13, 25]. Desreux [551]. detailed [133]. Detector [393]. Determination [325]. Deutscher [236]. Deutsches [211]. Development [137, 243, 149, 577, 639, 227, 267, 663, 418, 569, 431, 181, 729, 732, 72]. Dexter [219, 203, 4, 169, 27, 132, 190, 212, 227, 247, 270, 93, 157, 60, 215, 280, 362, 9, 18, 29, 42, 365, 630]. Dezember [770]. d’Histoire [39]. Diabetes [520]. diagrams [149]. Dialog [223]. Dialogue [736]. Dictionary [698, 508]. Dictionnaire [698, 508]. did [52, 8, 146]. Different [535, 534]. Digestive [520]. Digressions [5, 63, 30, 75, 16]. Dioxide [494, 578, 7]. Directory [11, 37, 68, 89, 112, 139, 153]. disappeared [65]. Discharge [334]. Discipline [348]. Disciplines [663, 675]. Discourse [122]. Discourses [97]. discover [8]. Discovered [302, 33]. Discoverer [307, 321]. Discovering [342]. Discovery [626, 272, 393, 389, 562, 563, 252, 314, 403, 344, 429]. Discussions [179]. Diseases [520]. Dispute [649]. Disputes [164]. Dissociation [408, 323, 407]. Distillation [233]. Distorting [132]. Distribution [383]. Disulfiram [392]. Divergence [359]. Diversions [5, 63, 30, 75, 16]. Diversity [684, 456]. Division [656, 9]. Divisional [11, 24, 46, 37, 68, 77, 89, 101, 112, 139, 153, 12]. Do [227, 453, 15]. Dobereiner [235]. Doctoral [192]. Documents [547, 548, 535, 642, 408, 605, 458, 166]. Does [671]. Domestic [444]. Donald [231]. Donnan [323]. Doors [311]. downs [95]. doyen [106]. Dr. [104, 320]. Draper [200]. Dream [554, 351, 15]. Dreams [611]. Drug [447, 240, 392, 298, 745]. Drugs [462, 615, 725]. Dry [390]. Duche [57]. Dulong [290]. Dumas [83]. During [587, 444, 358]. Dyer [56]. Dyes [284, 107]. Dyestuffs [94].

E. [483, 243]. Early [170, 205, 607, 606, 241, 411, 448, 503, 300, 293, 208, 391, 355, 539, 644, 517, 200, 317, 403, 108, 224, 502, 134, 401, 4, 30, 764, 444]. Earth [226, 763]. earths [6]. Eclectic [143, 565]. eclecticism [144]. Economic [86]. ecuaciones [99]. Ed [464, 474]. Edelgeborne [244]. Edelstein [360, 397, 348, 313, 329, 417, 363, 298, 630, 399, 381, 278, 328, 343, 362, 365]. Edgar [289, 106]. Edge [286, 281]. Edinburgh [607]. Edition [474]. Editions [236]. Editor [154, 197, 78, 250, 352, 426, 440, 452, 1, 13, 25, 376]. Editorial [531, 544, 560, 574, 599, 585, 596, 609, 620, 636, 652, 676]. Eds [483, 448]. EDTA [592]. Educating [118]. Education [456, 335, 444, 489, 82, 567]. Educator [180, 178, 509, 758]. Edward [211, 127, 743]. Effect [385, 480, 334, 696, 211]. Effects [187, 750, 226]. Ehregott [230]. Eighty [370]. Eighty-Five [370]. Eilhard [211, 211, 747]. Eirenaeus [145]. Eka [429]. Eka-Iodine [429]. Elbs [331]. Election [117, 11, 37]. Electrical [334]. Electricity [248]. Electrochemical [129]. electrochemistry [72, 67]. Electrode [214]. Electrogravimetry [72]. Electrolysis [128]. Electrolytic [209, 323]. Electromagnetic [310]. Electrometer [315]. electronics [94]. Electrons [129]. Elegant [185]. Elemens [39]. Element [302, 349, 359]. Élémentaire [55]. Elemental [419]. Elements [466, 763, 541, 229, 253, 324, 282, 649, 498, 446, 469, 569, 276, 746, 687, 226, 496, 719]. Elixir [615]. Ellen [461, 722, 252]. Elucidation [632]. Embedded [180]. emergence [31, 4]. Emergency [358]. Emigré [341]. Emil [549, 564]. Empire [381]. encyclopedia [695, 218]. Endowed [95]. ends [57]. Energy [614, 294, 146]. Engage [658]. Engagement [666]. Engineering [493, 492]. engineers [767, 35]. England [473, 762, 211, 519, 444, 743]. English [211, 546, 604, 641, 407, 457, 468, 534, 454]. English-Speaking [468]. Enjoyable [630]. enlightened [50]. Enlightenment [556]. Enough [715, 211]. entertainments [753, 513]. Environment [419]. Environmental [200]. Enzyme [344]. Episode [445]. Episodes [763, 226]. Epitomizing [657]. Equilibria [355, 224]. Equilibrium [234]. Equity [684]. Equivalency [129]. Era [647]. Erasmus [356]. Eric [496]. Erlich [222]. Ernest [307]. Errata [529]. Erratum [176, 305]. Erzherzogs [110]. Essay [413]. Essays [86, 769, 10, 22]. Esteem [340]. Estimating [479]. Ethics [568]. Etymological [698, 508]. étymologique [698, 508]. Eureka [420]. Europe [236, 764, 721, 467, 448]. European [475, 244, 147, 420]. Evan [488]. Even [149, 96, 66]. Events [24, 46, 37, 68, 77, 89, 101, 112, 139, 153]. Everitt [500]. evolucion [99]. Evolution [129]. Evolving [664]. Ewell [97]. Example [553, 7]. Examples [373]. exemplified [145]. existence [734]. Expansive [679]. expected [32]. Experience [489]. Experiment [536]. Experimentalist [53]. Experiments [224, 53]. Expertise [448, 764]. Explanation [144]. Explanations [535, 534]. Exploration [663]. explores [9, 18, 29, 42]. Exploring [660]. Explosives [499]. extraordinary [761, 110]. Eye [567]. eyes [5]. Eyring [423, 428, 212].

F [136]. F. [391]. fabric [49]. Fabrication [290]. Faces [503]. Factory [572, 731]. Fahs [106]. Failed [536]. Faith [120]. Fajans [425, 41, 61]. Fall [444, 21, 633]. Fallen [613]. fallout [739, 740, 460]. Fame [405]. Famed [342]. Family [643]. famous [83, 75, 84, 51]. Faraday [110, 474, 113, 131, 118, 124, 179, 117, 123, 119, 125, 417, 121, 120, 127, 756, 115, 129, 122, 116]. Farewell [11]. fates [51]. Father [590, 301]. favorite [7]. FDA [615]. Feigl [173]. Female [606, 340]. Ferdinand [592]. Festival [513, 753]. Festschrift [142]. Feuding [433]. Field [525]. Fifty [362, 618, 365]. Fight [311]. Figure [602]. Film [435]. Final [244]. Finally [525]. Finding [429]. Fine [237]. finitur [735]. Firmenich [99, 704]. First [462, 604, 237, 419, 648, 95, 199, 441, 539, 213, 350, 314, 39, 41, 18, 725, 16, 72, 86, 516, 221]. Fischer [564]. Fisher [450]. fission [30]. Five [370, 588]. Flack [618]. Florida [482]. Flour [494, 578]. Fluorescence [289]. Fluorine [123]. Foci [680]. Folin [231]. follow [48]. fondamenti [685, 137]. Food [243, 107, 210]. foremost [105]. Forensic [451, 691]. Foreword [655]. Forgotten [606, 445, 457, 667]. Formation [271, 287]. Formeln [44, 726]. Former [382]. formulas [726]. Foundation [567]. Foundations [618]. founder [12]. Four [288, 253, 324, 446]. Fourcroy [39]. Fourteen [167]. France [455]. Francis [306, 293]. Francois [232]. Francois-Pierre [232]. Frank [474]. Frankland [211, 743]. Franklin [162]. Frederick [613, 323]. French [239, 83, 505, 84, 234, 698, 508]. Friedlieb [333]. Fritz [236, 303, 173]. Fritzsche [644]. Fuel [549]. Fulhame [56]. Fulminans} [499]. Fundament [211]. Future [69, 81, 92, 103, 113, 670, 682, 375, 348, 671, 662, 659, 11, 24, 46, 37, 68, 77, 89, 101, 112, 139, 153].

G [449]. Gabriel [315]. Ganz [459]. Gap [137, 688]. Gas [536, 379, 273]. gases [705, 450]. gatekeeper [692, 218]. Gateway [660]. Gavroglu [484]. Geber [140]. Geburtstag [770]. Gel [390]. Gellert [230]. genannt [770]. Genealogical [265]. Genealogy [90]. General [391]. Generality [525]. Generation [472, 582, 733]. generators [91]. Genesis [72]. Genius [110, 417, 756]. Geoffrey [565]. Geologist [602]. George [323, 136, 249, 344]. Georges [168]. Georgius [132]. German [5, 38, 726, 734, 770]. Germany [468]. Giant [237]. Giants [288]. Gibber [140]. Gibbs [294]. Gilbert [518]. Girl [454, 517]. Giunta [464, 276]. Glass [385]. Gleditsch [461, 722, 252]. Glitters [76, 760]. Gmelin [402]. God [463, 717]. Good [576]. Gordon [394]. gourmet [107]. Grant [493, 491, 492, 486, 489, 488]. Granville [330]. graphical [726]. Graphischer [44, 726]. Great [542]. Greece [447, 745]. Green [430, 379, 527]. Greenhouse [211, 696]. Griep [465]. Grips [668]. Grotta [7]. Grotthuss [378]. Groups [137, 706]. Grow [164, 98]. Guide [79, 558]. Gustavus [225].

H [477, 243, 248]. Haber [236]. Hackh [445]. Half [95]. Hall [405, 194]. Hampshire [6]. Hand [56]. Handbook [451, 691]. Handbuch [402]. Happened [7, 20, 33, 65]. Happy [656]. Hare [273, 162]. Harkins [200]. Harry [75, 401]. Harvard [537]. Head [178]. Health [121]. Heart [332]. Heat [214]. Henry [423, 428, 105, 494, 635]. Henrys [218, 703]. Her [461, 163, 722]. Herbert [132, 569]. heritage [38]. Hermann [175, 741]. high [20, 75]. Higher [444]. Hill [76, 261, 718]. Him [226, 750]. himself [50]. Hinduism [446]. Hinrichs [225]. Hints [122]. Hispanic [478]. Hist [211, 326, 371, 532, 545, 561, 427, 373, 575, 586, 597, 600, 610, 621, 637, 653, 677, 611, 353, 374]. Histoire [236]. Historian [180, 43, 12]. Historians [683, 453, 105, 106]. historica [99]. Historical [76, 137, 218, 526, 523, 510, 478, 479, 171, 149, 515, 663, 418, 581, 664, 698, 508, 695, 710, 104, 729, 760, 157]. Historiography [449, 161, 364, 661, 728]. historique [698, 508]. History [10, 22, 35, 76, 137, 160, 175, 218, 226, 236, 450, 466, 449, 484, 474, 476, 550, 558, 572, 627, 679, 670, 656, 682, 684, 673, 180, 690, 370, 427, 660, 666, 395, 335, 763, 375, 658, 132, 241, 9, 18, 29, 42, 263, 348, 671, 372, 411, 520, 389, 500, 562, 329, 587, 662, 238, 453, 107, 294, 363, 731, 669, 247, 365, 379, 630, 539, 644, 664, 665, 266, 292, 325, 390, 404, 373, 582, 672, 188, 317, 403, 661, 686, 687, 689, 700, 701, 702, 705, 708, 30, 713, 27, 718, 149]. history [720, 721, 727, 728, 133, 73, 749, 60, 769, 757, 102, 82]. Hitchins [252]. Hoff [75]. Hofmann [80]. Hohenheim [770]. Homberg [33]. Home [293]. Homer [332]. honor [104]. Hoover [132]. Hopkins [349]. Horovitz [252]. Horstmann [407, 409]. Howard [537]. Hubert [104]. Hugh [466]. Hughes [334, 187]. Hugo [83]. Hulett [249]. human [96]. Humanity [522]. Humboldt [302]. Hundred [174, 674]. Hungarian [198, 242]. Hunter [463]. Hydrating [283]. Hydrazones [502]. Hydrocarbon [233]. Hydrocarbons [547, 274]. Hydrogen [70, 344, 235, 47]. Hydrolysis [186]. Hypotheses [422].

Iconoclast [168]. Iconoclastic [308]. Ideas [720, 175]. Identifying [589, 478]. if [54]. Ignoring [283]. Ihde [260, 261, 263, 262, 264]. Ii [645, 427, 39, 18, 292, 358, 481]. III [137, 685, 186, 55, 29]. Illinois [395, 318]. illustrate [7]. Imidazole [391]. Immediately [481]. impact [148]. Importance [669]. Important [613, 135, 41, 9, 52, 72]. Improved [273]. Impurity [99, 714]. Inc. [240]. including [736]. Inclusivity [684]. independent [56]. Index [201, 202, 258, 259, 346, 347, 437, 438]. Industrial [565, 108, 489]. industrialist [95]. Industry [218, 557, 96, 491, 364, 94, 223, 524, 661, 686, 703, 752, 76, 137]. Influence [180, 605, 275]. Influences [373]. Information [657, 546, 547, 548, 641, 629, 505, 520, 612, 675, 499]. Infrastructure [394]. Ingenious [56]. Ingeniously [505]. Ingold [183, 180, 182, 187, 178, 181]. Inorganic [525, 107]. Insights [420, 108]. installment [41, 61]. Institute [520, 629, 660]. Institution [110, 701, 756, 115]. Instruction [489]. Instructions [155, 388, 398, 497, 514, 530, 543, 559, 573, 598, 584, 595, 608, 619, 634, 650]. Instrument [227]. Instrumentation [157]. Instruments [695, 54, 218]. Integrity [553]. Intellectual [557, 481]. Interest [24, 46, 37, 68, 77, 89, 101, 112, 139, 153]. Interesting [294, 94]. International [137, 505, 587, 686, 31]. Interpretation [449, 626, 640, 533, 332, 728]. Intersection [664]. intertwinned [49].

Interview [554]. Interviewing [672]. Introducing [43]. Introduction [142, 546, 604, 641, 114, 177, 407, 457, 534, 48, 487]. Inventi [441]. Invention [357, 392, 273]. Inventions [592]. Inventor [211, 168, 405, 715]. Investments [567]. Invited [156]. Involved [344]. Iodine [342, 429]. Ionic [211, 696]. Ions [283]. Ipatieff [643]. Iron [174, 332]. ironic [51]. ironworks [19]. Irving [435]. ISBN [571]. Isn’t [536]. Isolation [325]. Isomerism [535, 534]. Isotopes [589, 252]. Israeli [242]. Issue [369, 32, 113]. Issues [69, 81, 92, 103, 113]. Iuliia [205]. IV [42, 325].

J [464, 474, 635]. J. [552, 400, 229]. Jabber [140]. Jaffa [210]. Jahrbuch [110]. James [474, 618, 349, 413, 308, 635, 109]. Jane [587]. Japanese [99, 759]. Japp [391]. Jazz [451, 691]. Jean [83, 467]. Jean-Baptiste [83]. Jeannette [485]. Jefferson [341]. Jeffrey [353]. Jeopardy [117]. Jo [472]. Johannes [195, 165, 166]. John [260, 473, 762, 226, 447, 449, 462, 414, 480, 750, 334, 86, 64]. Johnson [320]. Johnston [64]. Joint [384]. Jonas [645]. Jones [75, 494]. Joseph [76, 626, 71, 417, 340, 293, 159, 337, 748]. Josiah [442]. Journal [538, 567]. Journals [350]. Journey [630]. Jr [473, 136, 762]. Jr. [357, 443]. Jude [236]. Judgment [118]. Julius [516]. jumping [688, 137]. June [459]. Jungfer [244]. Just [140, 50]. Justus [218, 692]. juvenile [701].

Kamlet [645, 482]. Kampf [734]. Kansas [243]. Karl [331]. Karrer [384]. Kasimir [425, 41, 61]. Kaye [536]. Kazan’ [285, 640, 563, 74]. Keeping [375]. Kein [448]. Kekulé [554, 351, 75]. Key [224, 678]. Khemye [316]. Kidney [520]. Kinetic [470, 515]. Kinetics [457, 224]. Kizhner [546, 502, 579]. Klaus [563]. Kline [239]. Knowledge [501]. Known [386, 136, 770]. knows [54]. Knoxville [35, 749]. Koji [651]. Kolbe [175, 73, 741]. Konstitutions [44, 726]. Konstitutions-Formeln [44, 726]. Kopp [73]. Kostas [484]. Kubanek [461].

L [474]. Lab [648]. Labillardiere [267]. Laboratories [482, 239]. Laboratory [419, 227, 731, 489, 395, 318, 764, 572]. Lactic [137, 689]. Lafayette [335]. L’Affaire [196]. Lagerkvist [495]. Lamp [431]. Land [493, 491, 492, 486, 489, 488]. Land-Grant [493, 492, 486, 489]. Landmark [318]. Landmarks [515]. Langmuir [435]. Language [219, 303, 505]. larger [66]. largest [47]. Lars [471]. Last [365, 674, 32, 66]. late [85]. Latent [214]. Later [199]. Laura [85]. Lauren [460]. Lavoisier [55, 50, 53, 54, 52, 49, 51, 225]. Law [516]. Lawrence [110, 755]. Laws [243, 97, 128]. Lea [436, 301, 603]. Lead [566]. learn [34]. Learned [350]. least [20]. Lecithin [314]. Lecture [160, 150, 161, 82, 3, 700, 104, 38]. Lecturer [118]. Lectures [552, 700, 701]. Legacy [110, 566, 49, 755]. Leicester [105]. Leipzig [283]. Leo [483, 768]. Leopold [457, 110, 470]. Leprosy [330]. Lermontova [205]. Lesch [462]. Less [461, 722]. Lesser [386]. Let [232]. Letter [250, 352, 439, 440, 452, 376, 446, 162]. Letters [2, 14, 26, 88, 100, 111, 138, 152, 414, 382, 127]. level [31]. Levinstein [223]. Lewis [268]. lexicon [28]. Liebig [218, 692, 468]. Life [211, 461, 143, 261, 240, 423, 716, 320, 724, 732, 580, 616, 342, 83, 722, 571, 67]. Lifetime [211, 715]. Light [232, 323]. Lighter [235]. liked [106]. Limits [556]. Line [430]. Linton [85]. Linus [247]. Lippmann [315]. Liquid [249]. Liquids [428]. Lisbon [501]. List [130, 226, 750]. literal [21]. literary [17]. Literature [316, 98]. little [136]. little-known [136]. Lived [716, 211]. Lively [56]. Lives [405]. Lloyd [170]. Loeb [213, 251]. London [351, 591, 178]. Long [670]. look [97, 133, 147, 54, 136]. Looking [370, 375]. loose [57]. Lore [447, 745]. Loschmidt [216]. Lost [541, 435, 436, 736]. Love [460, 739, 740]. Lowry [191].

M [630]. M. [301, 603]. Made [110, 699]. Madison [718, 76]. Magic [639]. Magician [651]. Makers [433]. Making [137, 236, 313, 308, 689, 721]. Malley [476]. Man [110, 756]. Manchester [218, 571, 703]. manifesto [55]. Manipulation [125, 417]. Manipulator [126]. Manque [124]. Manual [719, 55, 64]. Manufacturing [385]. Manuscript [516]. Many [675]. Marcet [612, 587, 602]. March [651]. Marie [460, 461, 739, 740]. Marjorie [465, 476]. Mark [465]. market [764]. Markovnikov [604, 172, 433]. Markovnikov’s [604]. Marshall [105, 160]. Martha [197]. Martin [565]. Martine [467]. Mary [472]. Mass [489, 234, 146, 20]. master [71]. Mastering [614]. Materials [448, 764]. mathematics [4]. Matter [556, 572, 146, 97, 731]. Matveevich [502, 579]. Mauve [367]. Max [209]. May [651]. Mayer [146, 516]. Mcafee [204]. McCoy [569]. Mcduffie [204]. McEvoy [449]. Means [658]. Measurement [108]. Mechanics [428]. Mechanisms [185]. mechanization [145]. Mechanochemistry [301]. medals [6]. Medical [557]. Medicinal [663]. Medicine [451, 591, 691, 4, 725, 462]. Medium [187]. Meeting [633]. Meetings [11, 24, 46, 37, 68, 77, 89, 101, 112, 139, 153]. Meets [567, 75]. Mellor [71]. Melville [616]. Members [11, 24, 46, 37]. Memorabilia [106]. Memorial [255]. Memoriam [139]. Memories [264]. Men [137, 630, 742]. Mendeleev [641, 658, 282, 498]. Mentored [401]. Mentors [468]. Merck [237]. Mercury [478, 431, 33]. Mesomerism [181]. Message [11, 37, 353]. Metallurgic [230]. Metallurgical [332]. metallurgy [713, 757, 760, 218, 76]. Metals [421, 569, 736]. Metamorphosis [239]. Meter [207, 108]. Method [547, 414]. Méthode [712, 28]. Methodology [455]. Methods [632]. metodos [99]. Meyer [641]. Micelles [187]. Michael [110, 463, 472, 474, 119, 125, 417, 733, 208, 120, 756, 115]. microcrith [20]. Microreactors [187]. Mid [236, 412, 489, 738]. Mid-19th [412]. Mid-Twentieth [236, 738]. Mid-Western [489]. Midgley [357]. Mighty [164]. Mikasen [465]. Mikhailovich [172, 433]. Military [119]. Minds [542]. Mining [556, 467]. Miracle [462, 725]. Mirrored [179]. Mirrors [132, 120]. Missed [495, 765]. Mission [371]. Mistaken [422]. Misunderstandings [164]. Mitchell [57]. Mitscherlich [211, 747]. Mixed [191]. Mixter [518]. Mixtures [275]. Model [423]. Models [133]. Modern [175, 557, 626, 448, 363, 222, 647, 429, 4, 764, 766, 16, 49]. Mole [129]. Molecular [351, 732, 377]. molecules [734]. Molekülen [734]. Moment [582]. Monopoly [557]. Montana [200]. Monument [225]. Moors [274]. Morianen [419]. Morrill [194, 490, 486]. Morris [213, 251]. morte [735]. Moscow [283, 579]. Mosher [401]. most [9]. Motion [748, 76]. movies [711, 465]. Moving [664, 582]. Mrs. [56]. Much [705, 54, 450]. Mulliken [228]. Multidisciplinary [681, 664]. Multiple [191]. Munchen [211]. Munitions [227]. Münz [592]. Murder [451, 691]. Museum [211, 198, 21, 432, 134, 87, 74]. musings [27, 60]. Mutal [605]. Mutual [669]. My [735, 175]. Myer [210]. mysteries [17]. Mysterious [476, 727]. Mythical [404].

Nadezhda [205]. Naef [99, 704]. Naïveté [611]. Nakanishi [651]. Name [256]. Napoleon [234]. nascent [65, 95]. Nathan [480]. National [198, 505, 318, 358, 520, 405]. Natural [295, 337, 632]. Naturelle [39]. nazionale [685, 137]. nearly [135, 7]. Nebraska [170]. nec [735]. Necessarily [357]. Neglected [602, 41, 133]. Neither [708, 484]. Nernst [251, 275]. Network [501]. Neumann [414]. Newlands [229, 276]. News [11, 24, 46, 37, 68, 77, 89, 101, 112, 139, 153, 135]. Next [674]. NIDDK [520]. Niepce [272]. NIH [358]. Nikolai [284, 502, 579]. Nineteenth [99, 613, 493, 742, 714, 137, 468, 494]. Nineteenth-Century [613, 742, 137, 494]. nisi [735]. Nitrate [383, 45]. Nitric [411]. No [211, 229]. Nobel [495, 631, 646, 765]. Nobelpreisträger [236]. Noble [450, 536, 705]. Nomenclature [28, 712]. Nomination [215]. Nominations [278, 328, 343, 360, 397]. Nonclassical [256]. Nonmetallic [569]. nor [484, 708]. Norms [417]. Norris [618]. North [571]. Note [287, 626, 159]. Notebook [122]. Notes [10, 22, 35, 44, 76, 67, 110, 99, 151, 137, 175, 211, 218, 226, 236, 549, 411, 11, 24, 46, 37]. Nothing [450, 722, 705, 461]. Notice [141]. Notitia [441]. novel [134]. novels [106]. Novum [306]. Noyes [318, 275]. Nuclear [707, 503, 30, 459]. Number [251]. Numbers [201, 202]. Nuper [441]. Nutrition [210]. Nye [472]. nylon [715, 211].

Oak [310]. Oberlin [288]. Oblivion [613]. Obscurity [667]. Observations [124, 422]. oddities [134]. Odyssey [306]. Oesper [38]. Officer [68]. Officers [532, 545, 561, 575, 586, 597, 600, 610, 637, 653, 677, 11, 37, 89, 112, 139, 153, 621]. often [41]. Ohio [456]. Old [97, 17, 86, 64, 109, 136, 167, 193]. Older [536]. oldest [19]. Once [65]. One [211, 648, 421, 715, 105, 54]. only [53, 54]. Oparin [580, 628]. Opposition [287]. Oppposition [271]. orbital [732]. Order [219, 199]. Ores [174]. Organic [228, 10, 175, 552, 526, 182, 124, 179, 206, 710, 445, 618, 285, 639, 662, 185, 539, 184, 381, 401, 726, 107, 741, 769, 510]. Organischen [44, 726]. Organophosphorus [590]. Organum [306]. Origin [580, 569, 480]. Originally [236]. Origins [472, 557, 203, 409, 470, 222, 84, 150, 9, 733].

Oruro [479]. Osborne [169]. Ostwald [442, 294, 213]. other [218, 63, 703, 106]. Otto [143, 231]. Our [456, 113, 38]. Ouroboros [519]. overview [48]. Own [648]. oxidizing [45]. Oxozone [404]. oxygen [8]. Oxyhydrogen [40]. Ozohydrogen [404]. Ozone [266, 292, 325, 390, 404, 390].

p70 [211]. Painting [436]. Palace [189]. palate [107]. palette [107]. Pan [469]. Pan-Slavism [469]. Pantisocracy [339]. Paper [604, 399, 381]. Papers [546, 384, 443]. Paracelsus [770, 226, 770]. paradoxes [51]. parallels [83, 94, 51]. Paris [185, 625, 342]. Parisian [270]. Parke [241]. Parsons [228, 442]. Part [548, 50, 41, 28, 39, 55, 9, 18, 29, 42]. Parting [12, 47, 91, 80, 140, 102]. Partington [400]. Passion [594]. Past [67, 218, 682, 386, 659, 582, 34, 723, 11]. patent [84]. Path [504, 581]. Paths [478]. Pathway [128]. Pathways [611]. Patterns [449, 728]. Paul [206, 384]. Pauling [247]. Payen [267]. Penington [86]. Penn [488]. Pennsylvania [339]. People [295]. Peoria [395]. Perchlorate [383]. Percy [537]. Perfection [76, 748]. Period [266, 70]. Periodic [687, 495, 633, 415, 541, 471, 498, 469, 150, 765, 466]. Periodicity [229]. Perkins [330]. Peroxide [344]. Personal [226, 364, 188, 750, 115]. Perspective [523, 171, 82, 413, 49]. Perspectives [150, 161]. Petersburg [283]. Petit [290]. Petrochemicals [752, 76]. Pfaff [248]. Pfaundler [457, 470]. Pfizer [238]. pH [108, 207]. Pharmaceutical [557, 237]. Pharmacist [299]. Pharmacological [625]. Pharmacology [418]. Pharmacy [745, 447, 419]. Phenomenon [242]. Philalethes [145]. Phillipines [330]. Philosopher [337]. Philosophical [276]. Philosophy [593, 165]. Phlogistion [338]. Phospholipid [314]. Physical [549, 552, 400, 182, 179, 206, 442, 618, 298, 184, 325, 213]. Physician [602]. Physics [484, 203, 399, 708]. pictorial [104]. Picture [189, 5]. Pierre [460, 739, 740, 232, 192]. Pigment [499]. Pioneer [228, 204, 206]. Pioneering [546, 552, 456, 210]. Pioneers [231, 754, 512]. Piperidine [217]. Place [523, 247, 756, 110]. Plain [97]. Planet [523]. Plans [339]. Plastics [483, 768]. Platinum [421, 700]. Played [224]. Plurality [669]. Plus [185, 185]. Pm [349]. Poisoner [451, 691]. Poisonous [477, 693]. Polanyi [472, 733, 208]. Polarization [209]. Polish [410]. Political [51]. Politics [50, 341, 538, 524]. Pollution [478]. Polyacetylene [504, 581]. Polyaniline [644]. Polymer [539]. Polypyrrole [539]. Polytechnic [501]. Pool [379]. Poor [668]. Popular [313]. Porcelain [420]. Porfir’evish [67]. portrait [143, 85]. possible [699, 110]. Post [295]. Post-WWI [295]. Posthumouly [280]. Posthumous [631, 646]. Potashes [268]. Potential [660, 486, 214]. Potosí [479]. Pounding [311]. Powder [494, 578]. pp [571]. Practical [203, 517]. Practically [450, 705]. Practice [669]. precede [146]. Predecessors [464, 709]. Preface [656]. Prejudice [147]. Preparadores [501]. Preparation [547, 414]. Present [682, 673, 386, 659, 67]. Press [571]. preview [69, 81, 92, 103, 113]. Previous [436]. Pride [147]. Priestley [76, 626, 341, 340, 337, 748, 51, 339]. Primary [547, 548, 535, 642, 408, 605, 458, 166]. Prince [211, 747]. Principle [52, 134]. Principles [526, 510, 57, 282, 710]. Priority [429]. Private [517]. Prize [631, 646, 765, 495]. Probation [427]. Problem [217]. Problems [121, 27, 658]. Process [171, 481]. Product [207]. Production [171]. Products [661, 632]. Prof. [527, 528]. profession [148]. Professors [606]. Profiles [611]. program [730, 11, 24, 37, 110]. Progress [179, 385]. Project [310]. promethium [349]. Promise [486]. Promoted [502]. Proof [287]. Properties [97, 325, 45]. Property [557, 481]. proposed [16]. Prospero [473, 762]. proves [32]. Prussian [211, 389, 414, 441, 500, 562, 747]. Public [295, 488]. Publication [589, 441, 587]. Publications [612]. published [236]. Pugh [488]. Pure [243, 325]. Purple [499]. pursuit [144, 147]. puzzle [32]. pyridine [16]. pyrophorus [23, 33]. Pyrotartaric [570, 533]. Pythagoreanism [150].

Qualitative [228]. quanticule [61]. Quantum [484, 428, 754, 512, 708]. quarters [135]. Queries [8, 15, 36]. quest [736, 236]. Question [605]. Questions [8, 15, 36]. Quiet [175, 741]. quimicas [99]. Quinine [289].

R [400, 229, 391]. R. [629, 505]. Rachel [170]. Radical [359]. Radioactive [739, 740, 460]. Radioactivity [272, 727, 476]. Radiochemistry [614]. Radium [503]. Radon [307, 321]. Ramberg [528]. Rare [226, 763, 6]. Rasmussen [483]. rather [107]. Rationality [340]. Raymond [139]. Re [640]. Re-Interpretation [640]. Reaction [711, 185, 465]. Reactions [707, 470, 459]. Readings [76, 760]. Rearrangement [188]. Reception [70, 770]. recipe [23]. Recipients [630]. Reciprocal [275]. Recollections [137, 425, 688]. Records [93]. Rediscovering [570, 533]. Redniss [460]. Reflect [582]. Reflections [180, 542, 674, 755]. Refrigerants [357]. Refugee [410]. Regard [340]. Regina [410]. Reinventing [80]. Reinvention [420]. Reinvestigating [421]. Rejected [516]. Relations [643]. Relationship [223, 323]. Relationships [683]. Relevant [375]. Remarkable [405]. Remembered [629, 518]. Remembering [64]. Remote [672]. Removal [383]. Rennie [413]. Reply [322]. Report [24, 11, 37]. representation [726]. represented [31]. Reputation [117]. Research [110, 137, 629, 395, 394, 478, 479, 568, 405, 706, 730, 238]. Researcher [173]. Resolution [199]. resource [43]. Respect [669, 340]. Respectable [525]. Responds [245]. Response [528, 570, 526, 277, 436, 617, 70, 510]. Results [11, 37]. Returning [189]. Revealed [606]. Revealing [419, 414]. Reverie [554]. Review [459, 461, 464, 450, 451, 466, 447, 449, 462, 460, 463, 465, 485, 484, 472, 474, 476, 483, 473, 477, 496, 495, 571, 572, 526, 511, 594, 607, 510, 631, 646, 648, 649, 448, 593, 618, 583, 507, 617, 632, 513, 512, 509, 508, 211]. Reviews [246, 257, 269, 279, 291, 297, 304, 312, 319, 327, 336, 345, 354, 361, 368, 380, 387, 396, 406, 416, 424, 434, 506, 521, 540, 555, 601, 622, 623, 624, 638, 654].

Revisited [644]. Revisiting [419]. Revoltion [51]. Revolution [175, 449, 28, 39, 55, 190, 579, 728, 741, 57, 54]. Revolutionaries [51]. Revolutionary [59, 247]. Revolutions [338]. Rezeption [770, 226]. Rice [296]. Riddle [430]. Ridge [310]. Rights [557, 56]. Rigueur [571]. Rio [302]. Ripe [663]. Rise [76, 444, 94, 21, 752]. Rising [494, 578]. Rivalry [603]. Road [284]. Robert [236, 516, 110, 146, 537, 736, 761, 162]. Roberts [163]. Robinson [110, 183, 761]. Role [310, 365, 355, 224]. Roles [236, 645, 4, 738]. Rome [447, 745]. Room [673]. Root [455]. Roots [265]. Roscoe [223]. Royal [110, 701, 117, 756, 115]. Rubber [110, 730]. Ruhemann [254]. Rule [604, 433, 134]. ruled [107]. Rules [134]. Runge [333]. Russia [205, 590]. Russian [286, 644, 281]. Ruthenium [563]. Rutherford [307, 321].

S [243, 401]. S.A. [704, 99]. Sahyun [616]. Saint [272]. Saint-Victor [272]. Salt [500, 533, 275]. Salts [293, 289]. Saltzman [245]. same [56]. Sample [564]. Samuel [320, 228]. Sanitizer [379]. satire [63]. satorical [63]. saved [62]. say [52]. Scarborough [447]. Scerri [496, 277]. Scheibler [578]. Schoental [410]. Schonbein [266]. School [640, 629, 185, 591, 20, 501, 285]. Schools [454, 517]. Science [99, 175, 218, 558, 673, 493, 198, 161, 648, 717, 125, 577, 602, 208, 242, 538, 254, 695, 714, 720, 722, 727, 728, 733, 147, 741, 742, 60, 759, 99, 137, 175, 660, 461, 449, 463, 472, 476, 593]. Sciences [175, 210, 706, 766, 137]. Scientific [137, 553, 626, 394, 295, 505, 612, 365, 422, 51, 706]. Scientist [509, 732, 758, 67]. Scientists [542, 593, 234, 617]. Scope [187]. Search [123]. second [29, 548]. Secret [414]. Secrets [218, 507, 723, 737]. Section [413]. See [211]. Seeman [353]. seen [5, 54]. Seized [481]. selections [755]. Selectivity [662]. Self [494, 578]. Self-Rising [494, 578]. Semantics [122]. Semesters [175, 735]. Seminal [582]. Senter [344]. Series [160, 28, 39, 55, 9, 18, 29, 42]. Serrin [84]. served [135]. Service [490, 467]. Sesquicentenary [474]. Seth [483]. Seven [593, 617]. Seventeenth [467]. Seventeenth-Century [467]. Seymour [139]. Shadow [541]. Shaped [615]. Shifting [417]. Short [670]. Shots [12, 47, 91, 80, 140, 102]. Should [223]. Shumova [205]. Side [541]. Sidgwick [212]. Sidney [630]. Siegried [254]. Significant [428]. Silent [334]. Silica [390]. Silicon [419]. Silliman [443, 136]. Silver [220, 479, 33]. Simões [484]. Simple [642]. Since [364]. Sir [110, 189, 755]. Sister [668]. six [700]. Slator [309]. Slavism [469]. Slide [134]. Slyke [231]. Small [164, 30]. Smeaton [280]. Smith [349, 239, 106]. Smoke [393]. Snapshot [577]. Social [236, 472, 721, 733]. Societies [350, 96]. Society [62, 179, 233, 271, 287, 117, 196]. Socks [477, 693]. Sodium [80]. solid [94]. Solubility [275]. Solvent [187]. Some [688, 456, 515, 411, 286, 281, 84, 51, 106, 134, 12, 137, 83, 4, 27, 38, 94, 108, 60, 17]. Sons [417]. South [482]. Southey [339]. Soviet [579]. Spanish [220]. Spary [448]. Spawn [404]. Speaking [468]. special [113]. Specific [680]. Spiral [415]. Spoon [80]. Spring [603]. Spurious [421]. St. [283]. Staff [531, 544, 560, 574, 599, 609, 620, 636, 652, 676, 585, 596]. Stage [244].

Standard [249]. Stanley [255]. Stantienite [550]. staple [65]. State [374, 350, 65, 94, 488]. State-of-the-Union [374]. Statement [371]. States [61, 379, 184, 366]. Statics [458, 734, 457]. Statik [734]. Statistical [428]. Steel [332, 167]. Stefanie [252]. Stereochemical [199, 133]. Stereochemistry [163, 133]. Sterling [240]. Steroids [699, 110]. Steve [675]. still [54, 49]. Stones [34, 19]. Stools [73]. storia [685, 137]. Stories [606]. Story [526, 510, 649, 635, 651, 675, 710]. Strom [483]. Structural [639, 217]. Structure [550, 428, 445, 471, 632, 16]. struggle [734]. strung [75]. Student [382, 213]. Students [666, 658, 468, 262]. Studien [44]. Studies [355, 209, 703, 735, 175, 218]. studium [735]. Study [578, 538, 664, 41, 145]. Style [137, 212, 706]. Subject [202, 259, 347, 438, 52]. Subscribers [141]. Subsection [11]. Suffrage [648]. Sugar [224]. Sulfa [462, 725]. Sulfanilamide [615]. sulfide [47]. Superheavy [649]. Supplemental [657, 546, 547, 548, 641, 629, 505, 520, 612, 675, 499]. Supreme [651]. surrounding [12]. Survey [627]. Surviving [226, 750]. Svante [99, 382, 751]. Swimming [379]. symbolism [12]. Symbols [445]. Symposium [114, 770]. Symposiums [770]. Synergy [373]. Synonyms [359]. Synthesis [629, 391, 492, 481]. Synthetic [110, 284, 285, 569, 524, 730]. System [229, 471, 498, 469, 334, 109].

Table [495, 633, 541, 150, 765]. Tagebuch [110]. Taken [284]. Takeoff [505]. Tale [460, 496, 593, 746, 617, 739, 740]. Tales [466, 477, 693, 687]. Tangled [349]. Tarbell [255]. Tartar [533]. Teachers [658, 82]. Teaching [180, 445]. Tech [317]. Technological [207]. Technologies [661]. Technology [366, 467, 27, 198]. Ted [143]. Telling [526, 510, 710]. temperature [40]. Tennessee [35, 749]. Terminology [182]. Terpene [643]. test [47, 102]. Testing [293]. Tetraethyl [566]. Texas [317]. text [136]. Textbook [163, 39, 71, 7, 65, 149]. Textbooks [348, 442]. Texts [625, 20, 17]. Thalidomide [615]. theatricals [106]. Their [236, 370, 405, 385, 223, 350, 738]. Them [283]. Theme [505]. Theodor [143, 378]. Theophrastus [770]. Theories [298, 355]. theorist [53]. Theory [211, 187, 253, 428, 408, 470, 565, 639, 580, 628, 338, 696, 61, 732, 145, 407]. There [673, 232, 20]. Thermodynamic [515]. Thermodynamics [409, 516]. These [350, 134]. Theses [192]. things [106]. Third [552, 42]. Third-Year [552]. Thirteenth [625]. Thomas [483, 169, 357, 97, 57, 500, 191]. thought [56]. thoughts [4, 38]. Three [174, 422, 135]. three-quarters [135]. Thriving [684]. tie [63]. Time [606, 488, 56]. Timeline [131]. Times [236, 429, 71, 738]. Timetable [59]. Titrimetry [268]. Took [455]. Tool [666, 478, 479]. Tools [551, 681]. Top [542]. Topic [663]. Topics [680]. Toronto [11]. Toxic [523]. Toxicology [410]. Traces [419, 723, 218]. Trade [293]. Trail [378]. Traité [55]. Transfer [366]. Transference [251]. Transformations [643]. Transformed [462, 725]. transitional [136]. Translation [211, 546, 604, 641, 407, 457, 534, 516, 166]. Translations [3, 23, 45, 32]. transmutation [736]. transmutational [145]. transmute [33]. Transnational [501]. Transuranic [649]. Transylvania [87]. Traveler [558]. Trends [161]. Triangle [149]. Tribute [255, 71, 105]. Triphenyltetrazolium [481]. True [307, 321]. Truths [104]. trying [33]. tube [102]. Tungsten [220]. Turkish [303]. turn [5]. turn-of-the-century [5]. Turner [719]. Twelfth [625]. Twentieth [236, 493, 738]. Twentieth-Century [493]. Twice [302]. Twice-Discovered [302]. twist [30]. Two [606, 414, 503, 73, 41, 7, 51]. two-part [41]. Tying [57]. Type [149]. Tytler [109].

Ubbelohde [569]. UCL [697, 511]. Ulf [495]. Undergraduates [552]. Underrepresented [666]. Unintended [366]. Union [374]. Unit [358, 20]. United [61, 379, 184, 366]. Universität [770]. University [35, 76, 607, 718, 486, 770, 95, 160, 87, 537, 6, 95, 482, 300, 318, 625, 178, 749, 317, 74]. Unpublished [127]. unravel [3, 23]. Unraveling [218, 723]. unravels [12]. Update [90]. ups [95]. Ursula [448]. USDA [395]. Use [505, 499]. Used [413, 7]. UTK [35, 749].

vain [75]. Value [488]. van’t [75]. Vasil’evich [433]. Vedic [446]. Vegard [471]. Versus [377, 515, 298, 661]. Vestium [421]. VI [390]. Victor [83, 272, 84]. Victorian [211, 743]. View [115]. Viewed [76, 718]. VII [404]. Vindicating [56]. Virginia [96]. Viscosimeter [551]. visiting [5]. Vladimir [433, 643]. Volatile [533]. Volkova [205]. Voltaic [248]. Volume [642, 631, 646]. Volumes [641, 350].

W [473, 552, 71]. Wallace [211, 715]. Walter [473]. Walther [251, 275]. Walthère [603]. War [493, 583, 645, 227, 148, 358, 524, 481, 221, 648]. Warfare [295]. Warren [274]. Was [190, 247, 217, 53, 7, 20, 75]. Water [99, 243, 383, 714]. Waters [379]. Way [144]. Weapons [614, 568]. Web [349]. weekly [135]. Weeks [167]. Weidler [461]. Weight [52]. Weininger [675]. weiterer [734]. Welcome [369]. Werk [226, 770]. Werner [199, 384, 588]. Western [99, 489, 759]. Whatever [7, 20, 33, 65]. Whence [662]. Where [662, 54]. which [32, 700]. while [33, 11]. Whiteley [197]. Whither [662]. Who [588, 518, 401, 16]. wilderness [96]. William [480, 280, 477, 135, 268, 518, 200]. Williams [466]. Winners [365]. Winning [405]. Winstein [183]. Winthrop [473, 762]. Wisconsin [76, 718]. Wislicenus [195, 165, 166]. Wissenschaftliches [110]. Wiswesser [216]. within [49]. Wöhler [468]. Woman [410, 85]. Women [485, 475, 493, 205, 588, 405, 630, 738, 221, 311, 412, 591, 667, 647, 489, 401, 694, 56, 236, 591]. Wonderful [381]. Woodward [473, 183, 629, 505, 414]. Word [400]. work [143, 49, 770]. Worked [588]. Works [226, 750, 467]. World [31, 648, 645, 221, 47, 96, 478, 148, 358, 481, 766, 175]. Write [453]. Writing [670, 587, 363, 27, 60]. WWI [295].

Yale [518]. Ye [590]. Year [552, 66]. Years [362, 483, 633, 370, 607, 237, 199, 536, 618, 639, 318, 365, 174, 208, 592, 674, 486, 95, 768]. Yiddish [316]. York [691, 451]. Yost [536]. Young [195].

Zaitsev [433]. Zaitzev [172]. Ziber [205]. Ziber-Shumova [205]. Zinc [285, 687, 466]. Zinin [284, 644]. Zoroaster [253]. zum [770]. zur [734].

References

Jensen:1988:EDa

[1] William B. Jensen. From the Editor’s desk. Bulletin for the History of Chemistry, (1):3, Spring 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num1/num1-p3.pdf.

Anonymous:1988:La

[2] Anonymous. Letters. Bulletin for the History of Chemistry, (1):3, Spring 1988. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/~mainzv/HIST/bulletin_open_access/num1/num1-p3.
pdf.

Anonymous:1988:TCYa

[3] Anonymous. Translations. Can you unravel the chemistry of this 18th century lecture demonstration? Bulletin for the History of Chemistry, (1):3, Spring 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num1/num1-p3.pdf.

Debus:1988:DAS

[4] Allen Debus. The 1987 Dexter Address. Some thoughts on the roles of mathematics and medicine in the emergence of early modern chemistry. Bulletin for the History of Chemistry, (1):4–5, Spring 1988. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num1/num1-p4-5.pdf.

Anonymous:1988:DDP

[5] Anonymous. Diversions and digressions. A picture of turn-of-the-century American chemistry as seen through the eyes of a visiting German chemist. Bulletin for the History of Chemistry, (1):6–8, Spring 1988. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num1/num1-p6-8.pdf.

Jones:1988:CAR

[6] Paul Jones. Chemical artifacts. Rare earths and “medals” at the University of New Hampshire. Bulletin for the History of Chemistry, (1): 8–10, Spring 1988. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num1/
num1-p8-10.pdf.

Jensen:1988:WHLa

[7] William Jensen. Whatever happened to … ? Used to illustrate the density of carbon dioxide, the Grotta del Cane was a favorite textbook example for nearly two centuries. Bulletin for the History of Chemistry, (1):10–12, Spring 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num1/num1-p10-12.pdf.

Anonymous:1988:QQDa

[8] Anonymous. Questions and queries. Did the Chinese discover oxygen? Bulletin for the History of Chemistry, (1):12–13, Spring 1988. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num1/num1-p12-13.pdf.

Ihde:1988:HDAa

[9] Aaron Ihde. The history of the Dexter Award. Part I of this continuing series explores the origins of the Division’s most important award. Bulletin for the History of Chemistry, (1):13–14, Spring 1988. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num1/num1-p13-14.pdf.

Anonymous:1988:BNEa

[10] Anonymous. Book notes: Essays on the History of Organic Chemistry. Bulletin for the History of Chemistry, (1):14–15, Spring 1988. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num1/num1-p14-15.pdf.

Anonymous:1988:DNM

[11] Anonymous. Divisional news: Message from the Chair. Farewell from the Past Chair. Report of the Program Chair. Report of the Archeological Subsection. Election results. Awards. Notes from members. While in Toronto. Future meetings. 1988 Officers directory. Bulletin for the History of Chemistry, (1):15–16, Spring 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num1/num1-p15-16.pdf.

Bohning:1988:PSN

[12] James Bohning. Parting shots. The new divisional historian unravels some arcane symbolism surrounding the divisional founder. Bulletin for the History of Chemistry, (1):20–22, Spring 1988. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num1/num1-p20-22.pdf.

Jensen:1988:EDb

[13] William B. Jensen. From the Editor’s desk. Bulletin for the History of Chemistry, (2):3, Fall 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num2/num2%20p3.pdf.

Anonymous:1988:Lb

[14] Anonymous. Letters. Bulletin for the History of Chemistry, (2): 3, Fall 1988. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num2/
num2%20p3.pdf.

Anonymous:1988:QQDb

[15] Anonymous. Questions and queries. Do you have a dream? Bulletin for the History of Chemistry, (2):3, Fall 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num2/num2%20p3.pdf.

Rocke:1988:DDW

[16] Alan J. Rocke. Diversions and digressions. Who first proposed the modern structure for pyridine? Bulletin for the History of Chemistry, (2): 4–6, Fall 1988. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num2/
num2%20p4-6.pdf.

Williams:1988:OCS

[17] William D. Williams. Old chemistries. Some literary mysteries for connoisseurs of old chemistry texts. Bulletin for the History of Chemistry, (2):6–10, Fall 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num2/num2%20p6-10.pdf.

Ihde:1988:HDAb

[18] Aaron Ihde. The history of the Dexter Award. Part II of this continuing series explores the Award’s first decade. Bulletin for the History of Chemistry, (2):11–14, Fall 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num2/num2%20p11-14.pdf.

Habashi:1988:BSA

[19] Fathi Habashi. Bones and stones. The 250th anniversary of Canada’s oldest ironworks. Bulletin for the History of Chemistry, (2):14–16, Fall 1988. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num2/num2%20p14-16.pdf.

Jensen:1988:WHLb

[20] William Jensen. Whatever happened to … ? before the atomic mass unit there was the microcrith — at least in American high school chemistry texts. Bulletin for the History of Chemistry, (2):16–19, Fall 1988. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num2/num2%20p16-19.pdf.

Fine:1988:CAR

[21] Leonard Fine. Chemical artifacts. The rise and (literal) fall of the Chandler Museum at Columbia. Bulletin for the History of Chemistry, (2):19–21, Fall 1988. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num2/num2%20p19-21.pdf.

Anonymous:1988:BNEb

[22] Anonymous. Book notes: Essays in Chemical History and Chemistry in America. Bulletin for the History of Chemistry, (2):22, Fall 1988. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num2/num2%20p22.pdf.

Anonymous:1988:TCYb

[23] Anonymous. Translations. Can you unravel this 18th century recipe for a pyrophorus? Bulletin for the History of Chemistry, (2):22–23, Fall 1988. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num2/num2%20p22-23.pdf.

Anonymous:1988:DNR

[24] Anonymous. Divisional news: Report of the Program Chair. Awards. Notes from members. Events of interest. Future meetings. Bulletin for the History of Chemistry, (2):23–24, Fall 1988. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num2/num2%20p23-24.pdf.

Jensen:1989:ED

[25] William B. Jensen. From the Editor’s desk. Bulletin for the History of Chemistry, (3):3, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p3.pdf.

Anonymous:1989:L

[26] Anonymous. Letters. Bulletin for the History of Chemistry, (3):3, Spring 1989. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num3/
num3%20p3.pdf.

Haber:1989:DAS

[27] Lutz Haber. The 1988 Dexter Address. Some musings on the problems of writing on the history of chemical technology. Bulletin for the History of Chemistry, (3):4–7, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p4-7.pdf.

Chastain:1989:BCRa

[28] Ben Chastain. Books of the chemical revolution. Part I of this new series describes the lexicon of the revolution, the Méthode de Nomenclature Chimique of 1787. Bulletin for the History of Chemistry, (3):7– 11, Spring 1989. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num3/
num3%20p7-11.pdf.

Ihde:1989:HDAa

[29] Aaron Ihde. The history of the Dexter Award. Part III of this continuing series explores the Award’s second decade. Bulletin for the History of Chemistry, (3):11–15, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p11-15.pdf.

Habashi:1989:DDS

[30] Fathi Habashi. Diversions and digressions. A small twist in the early history of nuclear fission. Bulletin for the History of Chemistry, (3):15–16, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p15-16.pdf.

Bohning:1989:CCW

[31] James Bohning. A center of crystallization. The 1893 World’s Congress of Chemists represented the emergence of the ACS at the international level. Bulletin for the History of Chemistry, (3):16–21, Spring 1989. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num3/num3%20p16-21.pdf.

Anonymous:1989:TWL

[32] Anonymous. Translations. In which last issue’s puzzle proves more complex than expected. Bulletin for the History of Chemistry, (3): 21, Spring 1989. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num3/
num3%20p21.pdf.

Jensen:1989:WHL

[33] William Jensen. Whatever happened to … ? Discovered while trying to transmute mercury into silver: Homberg’s pyrophorus beguiled chemists for more than a century. Bulletin for the History of Chemistry, (3):21–24, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p21-24.pdf.

Allen:1989:BSW

[34] Ralph Allen. Bones and stones. What chemists can learn from the past. Bulletin for the History of Chemistry, (3):24–26, Spring 1989. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num3/num3%20p24-26.pdf.

Anonymous:1989:BNA

[35] Anonymous. Book notes: American Chemists and Chemical Engineers; Chemistry at UTK: a History of Chemistry at the University of Tennessee–Knoxville from 1794–1987. Bulletin for the History of Chemistry, (3):26, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p26.pdf.

Anonymous:1989:QQ

[36] Anonymous. Questions and queries. Bulletin for the History of Chemistry, (3):26–27, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p26-27.pdf.

Anonymous:1989:DNM

[37] Anonymous. Divisional news: Message from the Chair. Report of the Program Chair. Election results. Notes from members. Events of interest. Future meetings. 1989 Officers directory. Bulletin for the History of Chemistry, (3):27–28, Spring 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num3/num3%20p27-28.pdf.

Jones:1989:OLS

[38] Paul R. Jones. The 1989 Oesper Lecture. Some thoughts on our German chemical heritage. Bulletin for the History of Chemistry, (4):3– 7, Fall 1989. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num4/
num4%20p3-7.pdf.

Chastain:1989:BCRb

[39] Ben Chastain. Books of the chemical revolution. Part II of this series describes the first textbook of the revolution, Fourcroy’s Elemens d’Histoire Naturelle et de Chimie of 1786. Bulletin for the History of Chemistry, (4):8–11, Fall 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num4/num4%20p8-11.pdf.

Gonzalez:1989:BOB

[40] Elsa Gonzalez. Bochard and the oxyhydrogen blowpipe: Breaking the temperature barrier in the 18th century. Bulletin for the History of Chemistry, (4):11–15, Fall 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num4/num4%20p11-15.pdf.

Holmen:1989:KFF

[41] Reynold Holmen. Kasimir Fajans. The first installment of a two-part study of an important, but often neglected, 20th century chemist. Bulletin for the History of Chemistry, (4):15–23, Fall 1989. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num4/num4%20p15-23.pdf.

Ihde:1989:HDAb

[42] Aaron Ihde. The history of the Dexter Award. Part IV of this continuing series explores the Award’s third decade. Bulletin for the History of Chemistry, (4):23–26, Fall 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num4/num4%20p23-26.pdf.

Anonymous:1989:BIN

[43] Anonymous. The 1988 bibliography. Introducing a new resource for the historian of chemistry. Bulletin for the History of Chemistry, (4):27– 29, Fall 1989. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num4/
num4%20p27-29.pdf.

Anonymous:1989:BNC

[44] Anonymous. Book notes: Chemische Studien; Konstitutions-Formeln der Organischen Chemie in Graphischer Darstellung. Bulletin for the History of Chemistry, (4):29–31, Fall 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num4/num4%20p29-31.pdf.

Anonymous:1989:TCO

[45] Anonymous. Translations. Concerning the oxidizing properties of copper nitrate. Bulletin for the History of Chemistry, (4):31, Fall 1989. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num4/num4%20p31.pdf.

Anonymous:1989:DNA

[46] Anonymous. Divisional news: Awards. Notes from members. Events of interest. Future meetings. Bulletin for the History of Chemistry, (4): 31–32, Fall 1989. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num4/
num4%20p31-32.pdf.

Jensen:1989:PSW

[47] William B. Jensen. Parting shots. The world’s largest hydrogen sulfide test. Bulletin for the History of Chemistry, (4):34, Fall 1989. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num4/num4%20p34.pdf.

Miller:1989:IBO

[48] Jane A. Miller. Introduction. A brief overview of what is to follow and why. Bulletin for the History of Chemistry, (5):3, Winter 1989. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num5/num5%20p3.pdf.

Smeaton:1989:LLP

[49] William A. Smeaton. The legacy of Lavoisier. A perspective on how Lavoisier’s work is still intertwinned within the fabric of modern chemistry. Bulletin for the History of Chemistry, (5):4–10, Winter 1989. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num5/num5%20p4-10.pdf.

Donovan:1989:LPM

[50] Arthur L. Donovan. Lavoisier’s politics. More than just a chemist, Lavoisier considered himself as part of the enlightened bourgeoisie. Bulletin for the History of Chemistry, (5):10–14, Winter 1989. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num5/num5%20p10-14.pdf.

White:1989:SRC

[51] J. Edmund White. Scientific revolutionaries caught in political revoltion: Priestley and Lavoisier: Some ironic parallels and paradoxes in the scientific and political fates of two famous 18th century chemists. Bulletin for the History of Chemistry, (5):15–18, Winter 1989. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num5/num5%20p15-18.pdf.

Siegfried:1989:LCW

[52] Robert Siegfried. Lavoisier and the conservation of weight principle: What Lavoisier did and did not say on this important subject. Bulletin for the History of Chemistry, (5):18–24, Winter 1989. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num5/num5%20p18-24.pdf.

Holmes:1989:LEL

[53] Frederic L. Holmes. Lavoisier the experimentalist: Lavoisier was not only a theorist but a consummate designer of experiments and apparatus. Bulletin for the History of Chemistry, (5):24–31, Winter 1989. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num5/num5%20p24-31.pdf.

Schwartz:1989:IRL

[54] A. Truman Schwartz. Instruments of the Revolution: Lavoisier’s apparatus: Much of it is still to be seen, if one only knows where to look. Bulletin for the History of Chemistry, (5):31–34, Winter 1989. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num5/num5%20p31-34.pdf.

Chastain:1989:BCRc

[55] Ben B. Chastain. Books of the chemical revolution. Part III of this series describes the manual and manifesto of the revolution, Lavoisier’s Traité Élémentaire de Chimie of 1789. Bulletin for the History of Chemistry, (5):34–37, Winter 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num5/num5%20p34-37.pdf.

Davenport:1989:ILC

[56] Derek A. Davenport and Kathleen M. Ireland. The ingenious, lively and celebrated Mrs. Fulhame and the dyer’s hand: Vindicating the rights of women and independent chemical thought at the same time. Bulletin for the History of Chemistry, (5):37–42, Winter 1989. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num5/num5%20p37-42.pdf.

Jensen:1989:TDM

[57] William B. Jensen. Thomas Duche Mitchell and the chemistry of principles: Tying up the loose ends of the Chemical Revolution. Bulletin for the History of Chemistry, (5):42–48, Winter 1989. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num5/num5%20p42-48.pdf.

Anonymous:1989:BC

[58] Anonymous. A biographical checklist. Bulletin for the History of Chemistry, (5):48–49, Winter 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num5/num5%20p48-49.pdf.

Anonymous:1989:RT

[59] Anonymous. A revolutionary timetable. Bulletin for the History of Chemistry, (5):49–50, Winter 1989. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num5/num5%20p49-50.pdf.

Tarbell:1990:DAS

[60] D. Stanley Tarbell. The 1989 Dexter Address. Some musings on the writing of history of science. Bulletin for the History of Chemistry, (6): 3–7, Spring 1990. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num6/
num6%20p3-7.pdf.

Holmen:1990:KFC

[61] Reynold Holmen. Kasimir Fajans. This concluding installment deals with Fajans’ career in the United States and the controversy over quanticule theory. Bulletin for the History of Chemistry, (6):7–15, Spring 1990.

CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num6/num6%20p7-15.pdf.

Bohning:1990:CCS

[62] James J. Bohning. The continental chemical society. The almost demise of the ACS and the compromise that saved it. Bulletin for the History of Chemistry, (6):15–21, Spring 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num6/num6%20p15-21.pdf.

Doheny:1990:DDT

[63] James Doheny. Diversions and digressions. The tie that blinds and other satorical satire. Bulletin for the History of Chemistry, (6):21– 23, Spring 1990. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num6/
num6%20p21-23.pdf.

Williams:1990:OCR

[64] William D. Williams. Old chemistries. Remembering John Johnston’s Manual of Chemistry and its author. Bulletin for the History of Chemistry, (6):23–26, Spring 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num6/num6%20p23-26.pdf.

Jensen:1990:WHO

[65] William B. Jensen. Whatever happened to … ?. Once a textbook staple, the nascent state has disappeared — or has it? Bulletin for the History of Chemistry, (6):26–36, Spring 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num6/num6%20p26-36.pdf.

Anonymous:1990:BEL

[66] Anonymous. The 1989 bibliography. Even larger than last year. Bulletin for the History of Chemistry, (6):37–41, Spring 1990. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num6/num6%20p37-41.pdf.

Anonymous:1990:BNL

[67] Anonymous. Book notes: Life of a Scientist; Aleksandr Porfir’evish Borodin, a Chemist’s Biography; Electrochemistry, Past and Present. Bulletin for the History of Chemistry, (6):41–43, Spring 1990. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num6/num6%20p41-43.pdf.

Anonymous:1990:DNC

[68] Anonymous. Divisional news: From the Chair. Awards. Events of interest. Future meetings. 1990 Officer’s directory. Bulletin for the History of Chemistry, (6):43–44, Spring 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num6/num6%20p43-44.pdf.

Anonymous:1990:FIPa

[69] Anonymous. Future issues. A preview of what’s to come. Bulletin for the History of Chemistry, (6):47, Spring 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num6/num6%20p47.pdf.

Quane:1990:RHB

[70] Denis Quane. The reception of hydrogen bonding. The response of the chemical community in the period 1920–1937. Bulletin for the History of Chemistry, (7):3–13, Fall 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num7/num7%20p3-13.pdf.

Habashi:1990:JWM

[71] Fathi Habashi. Joseph W. Mellor. A tribute to the master textbook author of all times. Bulletin for the History of Chemistry, (7):13–16, Fall 1990. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num7/num7%20p13-16.pdf.

Stock:1990:GED

[72] John T. Stock. The genesis of electrogravimetry. On the development of the first important application of electrochemistry to chemical analysis. Bulletin for the History of Chemistry, (7):17–19, Fall 1990. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num7/num7%20p17-19.pdf.

Rocke:1990:BTS

[73] Alan J. Rocke. Between two stools. Kopp, Kolbe and the history of chemistry. Bulletin for the History of Chemistry, (7):19–24, Fall 1990. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num7/num7%20p19-24.pdf.

Wotiz:1990:CAB

[74] John H. Wotiz. Chemical artifacts. The Butlerov Museum at the University of Kazan. Bulletin for the History of Chemistry, (7):24–26, Fall 1990.

CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num7/num7%20p24-26.pdf.

Jensen:1990:DDK

[75] William B. Jensen. Diversions and digressions. Kekulé was vain and van’t Hoff was high strung — Harry Jones meets the famous and comments. Bulletin for the History of Chemistry, (7):26–33, Fall 1990. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num7/num7%20p26-33.pdf.

Anonymous:1990:BNA

[76] Anonymous. Book notes: All That Glitters. Readings in Historical Metallurgy; Petrochemicals: the Rise of an Industry; Motion Toward Perfection: the Achievement of Joseph Priestley, Chemistry as Viewed from Bascom Hill. A History of the Chemistry Department at the University of Wisconsin in Madison. Bulletin for the History of Chemistry, (7):33–35, Fall 1990. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num7/num7%20p33-35.pdf.

Anonymous:1990:DNEa

[77] Anonymous. Divisional news: Events of interest. Future meetings. Bulletin for the History of Chemistry, (7):35, Fall 1990. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num7/num7%20p35.pdf.

Anonymous:1990:ED

[78] Anonymous. From the Editor’s desk. Bulletin for the History of Chemistry, (7):36–37, Fall 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num7/num7%20p36-37.pdf.

Anonymous:1990:AG

[79] Anonymous. Author’s guide. Bulletin for the History of Chemistry, (7):37, Fall 1990. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num7/
num7%20p37.pdf.

Jensen:1990:PSR

[80] William B. Jensen. Parting shots. Reinventing the Hofmann sodium spoon. Bulletin for the History of Chemistry, (7):38–39, Fall 1990. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num7/num7%20p38-39.pdf.

Anonymous:1990:FIPb

[81] Anonymous. Future issues. A preview of what’s to come. Bulletin for the History of Chemistry, (7):39, Fall 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num7/num7%20p39.pdf.

Ihde:1990:PLH

[82] Aaron J. Ihde. Perspective lecture. History of chemistry and the education of teachers. Bulletin for the History of Chemistry, (8):3–8, Winter 1990. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num8/num8%20p3-8.pdf.

Chastain:1990:JBD

[83] Ben B. Chastain. Jean-Baptiste Dumas. Some curious parallels between the life and career of this famous French chemist and that of the famous French author, Victor Hugo. Bulletin for the History of Chemistry, (8):8–12, Winter 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num8/num8%20p8-12.pdf.

Stock:1990:VSO

[84] John T. Stock. Victor Serrin and the origins of the chainomatic balance: Some French antecedents of the famous Becker patent. Bulletin for the History of Chemistry, (8):12–15, Winter 1990. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num8/num8%20p12-15.pdf.

Creese:1990:LAL

[85] Mary R. S. Creese and Thomas Creese. Laura Alberta Linton: an American chemist: a portrait of a woman chemist in the late 19th century. Bulletin for the History of Chemistry, (8):15–18, Winter 1990. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num8/num8%20p15-18.pdf.

Williams:1990:OCC

[86] William D. Williams and Wyndham D. Miles. Old chemistries. Celebrating the bicentennial of America’s first chemistry book: John Penington’s chemical and economic essays. Bulletin for the History of Chemistry, (8):18–22, Winter 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num8/num8%20p18-22.pdf.

Bodner:1990:CAA

[87] George M. Bodner. Chemical artifacts. The Apparatus Museum at Transylvania University. Bulletin for the History of Chemistry, (8):22–27, Winter 1990. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num8/
num8%20p22-27.pdf.

Anonymous:1990:L

[88] Anonymous. Letters. Bulletin for the History of Chemistry, (8):30– 31, Winter 1990. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num8/
num8%20p30-31.pdf.

Anonymous:1990:DNEb

[89] Anonymous. Divisional news: Events of interest. Future meetings. 1991 Officers directory. Bulletin for the History of Chemistry, (8):31–32, Winter 1990. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num8/
num8%20p31-32.pdf.

Anonymous:1990:CGU

[90] Anonymous. Chemical genealogy update. Bulletin for the History of Chemistry, (8):33, Winter 1990. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num8/num8%20p33.pdf.

Jensen:1990:PSB

[91] William B. Jensen. Parting shots. Of beehives and Babo generators. Bulletin for the History of Chemistry, (8):34–35, Winter 1990. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num8/num8%20p34-35.pdf.

Anonymous:1990:FIPc

[92] Anonymous. Future issues. A preview of what’s to come. Bulletin for the History of Chemistry, (8):35, Winter 1990. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num8/num8%20p35.pdf.

Russell:1991:DAR

[93] Colin A. Russell. The 1990 Dexter Address. Records of chemistry: Combustion or conservation? Bulletin for the History of Chemistry, (9):3–7, Spring 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num9/num9%20p3-7.pdf.

Saltzman:1991:RDB

[94] Martin D. Saltzman and Alan L. Kessler. The rise and decline of the British dyestuffs industry. Some interesting parallels for the solid state electronics industry. Bulletin for the History of Chemistry, (9):7–15, Spring 1991. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num9/
num9%20p7-15.pdf.

Jones:1991:FHC

[95] Paul R. Jones. The first half century of chemistry at Clark University: Endowed by an industrialist in 1887, the first 50 years of chemistry at this nascent American university had its ups and downs. Bulletin for the History of Chemistry, (9):15–19, Spring 1991. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num9/num9%20p15-19.pdf.

DeLoach:1991:CIC

[96] Will S. DeLoach. Chemical industry in Colonial Virginia. All human societies have a chemical base, even in the wilderness of the New World. Bulletin for the History of Chemistry, (9):19–22, Spring 1991. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num9/num9%20p19-22.pdf.

Goldsmith:1991:OCL

[97] Robert H. Goldsmith. Old chemistries. A look at Thomas Ewell’s plain discourses on the laws or properties of matter. Bulletin for the History of Chemistry, (9):22–25, Spring 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num9/num9%20p22-25.pdf.

Anonymous:1991:BLC

[98] Anonymous. The 1990 bibliography. The literature continues to grow. Bulletin for the History of Chemistry, (9):25–32, Spring 1991. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num9/num9%20p25-32.pdf.

Anonymous:1991:BNS

[99] Anonymous. Book notes: A Science of Impurity: Water Analysis in Nineteenth Century Britain; Aproximacion a la evolucion historica de los metodos de adjuste de las ecuaciones quimicas; The Japanese and Western Science; Svante Arrhenius; From Chuit & Naef to Firmenich S.A. Bulletin for the History of Chemistry, (9):32–34, Spring 1991. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num9/num9%20p32-34.pdf.

Anonymous:1991:La

[100] Anonymous. Letters. Bulletin for the History of Chemistry, (9):34, Spring 1991. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num9/
num9%20p34.pdf.

Anonymous:1991:DNA

[101] Anonymous. Divisional news: Awards. Events of interest. Future meetings. Bulletin for the History of Chemistry, (9):34–35, Spring 1991. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num9/num9%20p34-35.pdf.

Williams:1991:PSB

[102] William D. Williams. Parting shots. A brief history of the test tube clamp. Bulletin for the History of Chemistry, (9):37–39, Spring 1991. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num9/num9%20p37-39.pdf.

Anonymous:1991:FIPa

[103] Anonymous. Future issues. A preview of what’s to come. Bulletin for the History of Chemistry, (9):39, Spring 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num9/num9%20p39.pdf.

Jensen:1991:DTC

[104] William B. Jensen. To demonstrate the truths of “chymistry”. An historical and pictorial celebration of the art of the lecture demonstration in honor of Dr. Hubert Alyea. Bulletin for the History of Chemistry, (10):3–15, Fall 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num10/num10%20p3-15.pdf.

Kauffman:1991:HML

[105] George B. Kauffman. Henry Marshall Leicester. A tribute to one of this century’s foremost chemical historians. Bulletin for the History of Chemistry, (10):15–21, Fall 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num10/num10%20p15-21.pdf.

Williams:1991:SEF

[106] William D. Williams. Some Edgar Fahs Smith memorabilia. Among other things, the doyen of American chemical historians liked cheap adventure novels and amateur theatricals. Bulletin for the History of Chemistry, (10):21–25, Fall 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num10/num10%20p21-25.pdf.

McKone:1991:HFC

[107] Harold T. McKone. The history of food colorants before aniline dyes. When inorganic rather than organic chemistry ruled both the artist’s palette and the gourmet’s palate. Bulletin for the History of Chemistry, (10):25–31, Fall 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num10/num10%20p25-31.pdf.

Stock:1991:EIP

[108] John T. Stock. Early industrial pH measurement and control. Some insights into automated pH control before the advent of the pH meter. Bulletin for the History of Chemistry, (10):31–34, Fall 1991. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num10/num10%20p31-34.pdf.

Williams:1991:OCC

[109] William D. Williams. Old chemistries. Celebrating the bicentennial of James Tytler’s system of chemistry. Bulletin for the History of Chemistry, (10):34–39, Fall 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num10/num10%20p34-39.pdf.

Anonymous:1991:BNC

[110] Anonymous. Book notes: From Caveman to Chemist; Robert Robinson: Chemist Extraordinary; Michael Faraday and the Royal Institution: the Genius of Man and Place; The Legacy of Sir Lawrence Bragg; Wissenschaftliches Jahrbuch 1990 Das Tagebuch des Erzherzogs Leopold; The American Synthetic Rubber Research Program; Steroids Made It Possible. Bulletin for the History of Chemistry, (10):39–45, Fall 1991. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num10/num10%20p39-45.pdf.

Anonymous:1991:Lb

[111] Anonymous. Letters. Bulletin for the History of Chemistry, (10):45, Fall 1991. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/~mainzv/HIST/bulletin_open_access/num10/num10%20p45.
pdf.

Anonymous:1991:DNC

[112] Anonymous. Divisional news: From the Chair. Events of interest. Future meetings. 1992 Officers directory. Bulletin for the History of Chemistry, (10):45–46, Fall 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num10/num10%20p45-46.pdf.

Anonymous:1991:FIPb

[113] Anonymous. Future issues. A preview of our special Faraday issue. Bulletin for the History of Chemistry, (10):47, Fall 1991. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num10/num10%20p47.pdf.

Davenport:1991:SI

[114] Derek A. Davenport. Symposium introduction. Bulletin for the History of Chemistry, (11):3–4, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p3-4.pdf.

Thomas:1991:RIM

[115] John Meurig Thomas. The Royal Institution and Michael Faraday: a personal view. Bulletin for the History of Chemistry, (11):4–9, Winter 1991. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num11/num11%20p4-9.pdf.

Williams:1991:FHB

[116] L. Pearce Williams. Faraday and his biographers. Bulletin for the History of Chemistry, (11):9–17, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p9-17.pdf.

Fullmer:1991:FER

[117] June Z. Fullmer and Melvyn C. Usselman. Faraday’s election to the Royal Society: a reputation in jeopardy. Bulletin for the History of Chemistry, (11):17–28, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p17-28.pdf.

Cantor:1991:EJF

[118] Geoffrey Cantor. Educating the judgment: Faraday as a lecturer. Bulletin for the History of Chemistry, (11):28–36, Winter 1991. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num11/num11%20p28-36.pdf.

James:1991:MCC

[119] Frank A. J. L. James. The military context of chemistry: the case of Michael Faraday. Bulletin for the History of Chemistry, (11):36–40, Winter 1991. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num11/num11%20p36-40.pdf.

Pratt:1991:MFB

[120] Herbert T. Pratt. Michael Faraday’s Bibles as mirrors of his faith. Bulletin for the History of Chemistry, (11):40–47, Winter 1991. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num11/num11%20p40-47.pdf.

OBrien:1991:FHP

[121] James F. O’Brien. Faraday’s health problems. Bulletin for the History of Chemistry, (11):47–50, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p47-50.pdf.

Tweney:1991:FCH

[122] Ryan D. Tweney. Faraday’s 1822 chemical hints notebook and the semantics of chemical discourse. Bulletin for the History of Chemistry, (11):51–55, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p51-55.pdf.

Goldwhite:1991:FSF

[123] Harold Goldwhite. Faraday’s search for fluorine. Bulletin for the History of Chemistry, (11):55–60, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p55-60.pdf.

Davenport:1991:OFO

[124] Derek A. Davenport. Observations on Faraday as organic chemist manque. Bulletin for the History of Chemistry, (11):60–65, Winter 1991. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num11/num11%20p60-65.pdf.

Jensen:1991:MFA

[125] William B. Jensen. Michael Faraday and the art and science of chemical manipulation. Bulletin for the History of Chemistry, (11):65–76, Winter 1991. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num11/num11%20p65-76.pdf.

Ross:1991:CM

[126] Sydney Ross. The chemical manipulator. Bulletin for the History of Chemistry, (11):76–79, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p76-79.pdf.

Ross:1991:ULF

[127] Sydney Ross. Unpublished letters of Faraday and others to Edward Daniel Clarke. Bulletin for the History of Chemistry, (11):79–86, Winter 1991. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num11/num11%20p79-86.pdf.

Stock:1991:PLE

[128] John T. Stock. The pathway to the laws of electrolysis. Bulletin for the History of Chemistry, (11):86–92, Winter 1991. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num11/num11%20p86-92.pdf.

Towns:1991:EEM

[129] Marcy Hamby Towns and Derek A. Davenport. From electrochemical equivalency to a mole of electrons: the evolution of the Faraday. Bulletin for the History of Chemistry, (11):92–100, Winter 1991. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num11/num11%20p92-100.pdf.

Anonymous:1991:BCL

[130] Anonymous. A biographical check list. Bulletin for the History of Chemistry, (11):100–101, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p100-101.pdf.

Anonymous:1991:FT

[131] Anonymous. A Faraday timeline. Bulletin for the History of Chemistry, (11):101–104, Winter 1991. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num11/num11%20p101-104.pdf.

Hannaway:1992:DAH

[132] Owen Hannaway. The 1991 Dexter Address. Herbert Hoover and Georgius Agricola: the distorting mirrors of history. Bulletin for the History of Chemistry, (12):3–10, Fall 1992. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num12/num12%20p3-10.pdf.

Paoloni:1992:SMB

[133] Leonello Paoloni. Stereochemical models of benzene, 1869–1875. A detailed look at a neglected chapter in the history of stereochemistry. Bulletin for the History of Chemistry, (12):10–24, Fall 1992. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num12/num12%20p10-24.pdf.

Williams:1992:SEC

[134] William D. Williams. Some early chemical slide rules. An early and novel application of the principle of the slide rule, these chemical oddities are now museum curiosities. Bulletin for the History of Chemistry, (12):24–29, Fall 1992. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num12/num12%20p24-29.pdf.

Brock:1992:CNB

[135] William H. Brock. The Chemical News, 1859–1932. The brainchild of William Crookes, this important chemical weekly served the chemical community for nearly three-quarters of a century. Bulletin for the History of Chemistry, (12):30–35, Fall 1992. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num12/num12%20p30-35.pdf.

Williams:1992:OCL

[136] William D. Williams. Old chemistries. A look at a little-known transitional chemical text by Benjamin Silliman, Jr. and George F. Baker. Bulletin for the History of Chemistry, (12):35–38, Fall 1992. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num12/num12%20p35-38.pdf.

Anonymous:1992:BNH

[137] Anonymous. Book notes: A History of the International Chemical Industry; Chemistry; The Historical Development of Chemical Concepts; Nineteenth-Century Attitudes: Men of Science; A History of Lactic Acid Making: a Chapter in the History of Biotechnology; Contrasts in Scientific Style: Research Groups in the Chemical and Biochemical Sciences; Atti del III convegno nazionale di storia e fondamenti della chimica; Some Recollections of Gap Jumping. Bulletin for the History of Chemistry, (12):39–45, Fall 1992. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num12/num12%20p39-45.pdf.

Anonymous:1992:La

[138] Anonymous. Letters. Bulletin for the History of Chemistry, (12):45–46, Fall 1992. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/~mainzv/HIST/bulletin_open_access/num12/num12%20p45-
46.pdf.

Anonymous:1992:DNAa

[139] Anonymous. Divisional news: Awards. Events of interest. In memoriam: Raymond Benedict Seymour. Future meetings. Officers directory. Bulletin for the History of Chemistry, (12):46–47, Fall 1992. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num12/num12%20p46-47.pdf.

Jensen:1992:PSG

[140] William B. Jensen. Parting shots. Gibber, jabber, or just Geber? Bulletin for the History of Chemistry, (12):49–51, Fall 1992. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num12/num12%20p49-51.pdf.

Anonymous:1992:NAS

[141] Anonymous. Notice to all subscribers and authors. Bulletin for the History of Chemistry, (12):51, Fall 1992. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num12/num12%20p51.pdf.

Anonymous:1992:FI

[142] Anonymous. Festschrift introduction. Bulletin for the History of Chemistry, (13–14):3, Winter–Spring 1992–1993. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num13-14/num13-14%20p3.pdf.

Bohning:1992:ELO

[143] James J. Bohning. The eclectic life of Otto Theodor Benfey. A biographical portrait of the life and work of Ted Benfey. Bulletin for the History of Chemistry, (13–14):4–16, Winter–Spring 1992–1993. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num13-14/num13-14%20p4-16.pdf.

Benfey:1992:WEE

[144] Otto Theodor Benfey. By way of explanation. on eclecticism and the pursuit of chemistry. Bulletin for the History of Chemistry, (13–14): 16–19, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num13-14/num13-14%20p16-19.pdf.

Newman:1992:CAC

[145] William Newman. Corpuscular alchemy. A case study of the mechanization of 17th-century alchemy as exemplified by the transmutational theory of Eirenaeus Philalethes. Bulletin for the History of Chemistry, (13– 14):19–27, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num13-14/num13-14%20p19-27.pdf.

Caneva:1992:RMC

[146] Kenneth L. Caneva. Robert Mayer and the conservation of matter. Did conservation of energy precede conservation of mass? Bulletin for the History of Chemistry, (13–14):27–29, Winter–Spring 1992–1993. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num13-14/num13-14%20p27-29.
pdf.

Rocke:1992:PPC

[147] Alan J. Rocke. Pride and prejudice in chemistry. A look at chauvinism and the pursuit of science among 19th-century European chemists. Bulletin for the History of Chemistry, (13–14):29–40, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num13-14/
num13-14%20p29-40.pdf.

Rhees:1992:CWA

[148] David J. Rhees. The chemist’s war. An analysis of the impact of World War I on the American chemical profession. Bulletin for the History of Chemistry, (13–14):40–47, Winter–Spring 1992–1993. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num13-14/num13-14%20p40-47.pdf.

Jensen:1992:HDV

[149] William B. Jensen. The historical development of the van Arkel bondtype triangle: Even textbook diagrams have a history. Bulletin for the History of Chemistry, (13–14):47–59, Winter–Spring 1992–1993. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num13-14/num13-14%20p47-59.
pdf.

Benfey:1992:PLP

[150] Otto Theodor Benfey. Perspectives lecture. On Pythagoreanism and the origins of the periodic table. Bulletin for the History of Chemistry, (13– 14):60–66, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num13-14/num13-14%20p60-66.pdf.

Anonymous:1992:BN

[151] Anonymous. Book notes. Bulletin for the History of Chemistry, (13– 14):66–72, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num13-14/num13-14%20p66-72.pdf.

Anonymous:1992:Lb

[152] Anonymous. Letters. Bulletin for the History of Chemistry, (13–14): 72–73, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num13-14/num13-14%20p72-73.pdf.

Anonymous:1992:DNAb

[153] Anonymous. Divisional news: Awards. Events of interest. Future meetings. Officers directory. Bulletin for the History of Chemistry, (13– 14):73–75, Winter–Spring 1992–1993. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num13-14/num13-14%20p73-75.pdf.

Anonymous:1994:E

[154] Anonymous. From the editor. Bulletin for the History of Chemistry, (15–16):i, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20pi.pdf.

Anonymous:1994:IA

[155] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, (15–16):ii, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20pii.pdf.

Anonymous:1994:AI

[156] Anonymous. Applications invited. Bulletin for the History of Chemistry, (15–16):ii, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20pii.pdf.

Stock:1994:DAA

[157] John T. Stock. The 1992 Dexter Award address. Historical chemical instrumentation. Bulletin for the History of Chemistry, (15–16):1–8, ???? 1994. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num15-16/
num15-16%20p1-8.pdf.

Webb:1994:CC

[158] George E. Webb. The chemist as consultant. Bulletin for the History of Chemistry, (15–16):9–14, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20p9-14.pdf.

Palter:1994:NJB

[159] Robert Palter. A note on Joseph Black. Bulletin for the History of Chemistry, (15–16):15–17, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20p15-17.pdf.

Anonymous:1994:HCL

[160] Anonymous. History of chemistry lecture series. Marshall University. Bulletin for the History of Chemistry, (15–16):18, ???? 1994. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num15-16/num15-16%20p18.pdf.

Gillispie:1994:PLR

[161] Charles C. Gillispie. Perspectives lecture: Recent trends in historiography of science. Bulletin for the History of Chemistry, (15–16):19–26, ???? 1994. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num15-16/
num15-16%20p19-26.pdf.

Williams:1994:LFB

[162] William D. Williams. A letter from Franklin Bache to Robert Hare. Bulletin for the History of Chemistry, (15–16):27–29, ???? 1994. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num15-16/num15-16%20p27-29.
pdf.

Creese:1994:CRH

[163] Mary R. S. Creese and Thomas M. Creese. Charlotte Roberts and her textbook on stereochemistry. Bulletin for the History of Chemistry, (15–16):31–36, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20p31-36.pdf.

Saltzman:1994:SMM

[164] Martin D. Saltzman. From small misunderstandings mighty disputes grow. Bulletin for the History of Chemistry, (15–16):37–44, ???? 1994. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num15-16/num15-16%20p37-44.
pdf.

Ramberg:1994:CJW

[165] Peter J. Ramberg. Commentary: Johannes Wislicenus, atomism, and the philosophy of chemistry. Bulletin for the History of Chemistry, (15–16):45–51, ???? 1994. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num15-16/num15-16%20p45-51.pdf.

Ramberg:1994:PDT

[166] Peter J. Ramberg. Primary documents: a translation of Johannes Wislicenus. Bulletin for the History of Chemistry, (15–16):52–54, ???? 1994. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num15-16/
num15-16%20p52-54.pdf.

Williams:1994:OCS

[167] William D. Williams. Old chemistries: Steel’s fourteen weeks in chemistry. Bulletin for the History of Chemistry, (15–16):55–58, ???? 1994. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num15-16/num15-16%20p55-58.
pdf.

Laszlo:1994:GDI

[168] Pierre Laszlo. Georges Darzens (1867–1954): Inventor and iconoclast. Bulletin for the History of Chemistry, (15–16):59–64, ???? 1994. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num15-16/num15-16%20p59-64.
pdf.

Fruton:1995:DAA

[169] Joseph S. Fruton. The 1993 Dexter Award Address. Thomas Burr Osborne and chemistry. Bulletin for the History of Chemistry, (17–18):1–8, ???? 1995. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num17-18/
num17-18%20p1-8.pdf.

Creese:1995:RLE

[170] Mary R. S. Creese and Thomas M. Creese. Rachel Lloyd: Early Nebraska chemist. Bulletin for the History of Chemistry, (17–18):9–14, ???? 1995. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num17-18/
num17-18%20p9-14.pdf.

Habashi:1995:BPA

[171] Fathi Habashi. Bayer’s process for alumina production: a historical perspective. Bulletin for the History of Chemistry, (17–18):15–19, ???? 1995. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num17-18/
num17-18%20p15-19.pdf.

Lewis:1995:AMZ

[172] David E. Lewis. Aleksandr Mikhailovich Zaitzev (1841–1910). Markovnikov’s conservative contemporary. Bulletin for the History of Chemistry, (17–18):21–30, ???? 1995. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num17-18/num17-18%20p21-30.pdf.

Espinola:1995:FFC

[173] Aida Espinola, Mario Abrantes da Silva Pinto, and Claudio Costa Neto. Fritz Feigl (1891–1971). The centennial of a researcher. Bulletin for the History of Chemistry, (17–18):31–39, ???? 1995. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num17-18/num17-18%20p31-39.pdf.

Olsen:1995:THY

[174] Kevin K. Olsen. Three hundred years of assaying American iron and iron ores. Bulletin for the History of Chemistry, (17–18):41–56, ???? 1995. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/~mainzv/HIST/bulletin_open_access/num17-18/
num17-18%20p41-56.pdf.

Anonymous:1995:BNC

[175] Anonymous. Book notes: Chemical Sciences in the Modern World; Ideas in Chemistry. A History of the Science; My 132 Semesters of Chemistry Studies; The Quiet Revolution: Hermann Kolbe and the Science of Organic Chemistry. Bulletin for the History of Chemistry, (17–18):57–64, ???? 1995. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num17-18/num17-18%20p57-64.pdf.

Anonymous:1995:CE

[176] Anonymous. Credit and erratum. Bulletin for the History of Chemistry, (17–18):64, ???? 1995. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num17-18/num17-18%20p64.pdf.

Ingold:1996:I

[177] K. U. Ingold. Introduction. Bulletin for the History of Chemistry, (19):1, ???? 1996. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num19/num19%20p1.pdf.

Roberts:1996:CKI

[178] Gerrylynn K. Roberts. C. K. Ingold at University College London: Educator and department head. Bulletin for the History of Chemistry, (19):2–12, ???? 1996. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num19/num19%20p2-12.pdf.

Davenport:1996:PPO

[179] Derek A. Davenport. The progress of physical organic chemistry as mirrored in the Faraday Society Discussions of 1923, 1937, and 1941. Bulletin for the History of Chemistry, (19):13–18, ???? 1996. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num19/num19%20p13-18.pdf.

Benfey:1996:TCE

[180] Theodor Benfey. Teaching chemistry embedded in history: Reflections on C. K. Ingold’s influence as historian and educator. Bulletin for the History of Chemistry, (19):19–24, ???? 1996. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num19/num19%20p19-24.pdf.

Saltzman:1996:CKI

[181] Martin D. Saltzman. C. K. Ingold’s development of the concept of mesomerism. Bulletin for the History of Chemistry, (19):25–32, ???? 1996. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num19/num19%20p25-32.pdf.

Bunnet:1996:POT

[182] Joseph F. Bunnet. Physical organic terminology, after Ingold. Bulletin for the History of Chemistry, (19):33–42, ???? 1996. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num19/num19%20p33-42.pdf.

Barton:1996:IRW

[183] Derek H. R. Barton. Ingold, Robinson, Winstein, Woodward, and I. Bulletin for the History of Chemistry, (19):43–47, ???? 1996. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num19/num19%20p43-47.pdf.

Roberts:1996:BPO

[184] John D. Roberts. The beginnings of physical organic chemistry in the United States. Bulletin for the History of Chemistry, (19):48–56, ???? 1996. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num19/num19%20p48-56.pdf.

Nye:1996:PCP

[185] Mary Jo Nye. “Plus commode et plus elegant”: the Paris school of organic reaction mechanisms in the 1920’s and 1930’s. Bulletin for the History of Chemistry, (19):58–65, ???? 1996. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num19/num19%20p58-65.pdf.

Basolo:1996:BHC

[186] Fred Basolo. Base hydrolysis of cobalt (III) amines. Bulletin for the History of Chemistry, (19):66–71, ???? 1996. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num19/num19%20p66-71.pdf.

Bunton:1996:MEM

[187] Clifford A. Bunton. Medium effects of micelles as microreactors and the scope of the Hughes–Ingold solvent theory. Bulletin for the History of Chemistry, (19):72–76, ???? 1996. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num19/num19%20p72-76.pdf.

Shine:1996:PHB

[188] Henry J. Shine. A personal history of the benzidine rearrangement. Bulletin for the History of Chemistry, (19):77–92, ???? 1996. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num19/num19%20p77-92.pdf.

Anonymous:1996:PSC

[189] Anonymous. Picture of Sir Christopher returning from Buckingham Palace. Bulletin for the History of Chemistry, (19):93, ???? 1996. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num19/num19%20p93.pdf.

Holmes:1997:DAA

[190] Frederic L. Holmes. The 1994 Dexter Award address. What was the chemical revolution about? Bulletin for the History of Chemistry, (20):1–9, ???? 1997. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num20/num20%20p1-9.pdf.

Saltzman:1997:TML

[191] Martin D. Saltzman. Thomas Martin Lowry and the mixed multiple bond. Bulletin for the History of Chemistry, (20):10–17, ???? 1997. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num20/num20%20p10-17.pdf.

Stock:1997:DTP

[192] John T. Stock. The doctoral theses of Pierre Adolphe Bobierre. Bulletin for the History of Chemistry, (20):18–23, ???? 1997. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num20/num20%20p18-23.pdf.

Williams:1997:OC

[193] William D. Williams. Old chemistries. Bulletin for the History of Chemistry, (20):24–26, ???? 1997. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num20/num20%20p24-26.pdf.

Jones:1997:MHC

[194] Paul R. Jones. A Morrill Hall for chemistry. Bulletin for the History of Chemistry, (20):27, ???? 1997. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num20/num20%20p27.pdf.

Jones:1997:YJW

[195] Paul R. Jones. The young Johannes Wislicenus in America. Bulletin for the History of Chemistry, (20):28–32, ???? 1997. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num20/num20%20p28-32.pdf.

Rossiter:1997:SAC

[196] Margaret W. Rossiter. “But she’s an avowed Communist!” L’Affaire Curie at the American Chemical Society. Bulletin for the History of Chemistry, (20):33–41, ???? 1997. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num20/num20%20p33-41.pdf.

Creese:1997:MAW

[197] Mary R. S. Creese. Martha Annie Whiteley: Chemist and editor. Bulletin for the History of Chemistry, (20):42–45, ???? 1997. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num20/num20%20p42-45.pdf.

Caswell:1997:HNM

[198] Lyman R. Caswell. The Hungarian National Museum for Science and Technology. Bulletin for the History of Chemistry, (20):46–49, ???? 1997. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num20/num20%20p46-49.pdf.

Kauffman:1997:SAF

[199] George B. Kauffman. A stereochemical achievement of the first order: Alfred Werner’s resolution of cobalt complexes, 85 years later. Bulletin for the History of Chemistry, (20):50–59, ???? 1997. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num20/num20%20p50-59.pdf.

Rice:1997:WDH

[200] Richard E. Rice and George B. Kauffman. William Draper Harkins: an early environmental chemist in Montana. Bulletin for the History of Chemistry, (20):60–67, ???? 1997. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num20/num20%20p60-67.pdf.

Anonymous:1997:AIN

[201] Anonymous. Author index, numbers 1–20. Bulletin for the History of Chemistry, (20):73, ???? 1997. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num20/num20%20p73.pdf.

Anonymous:1997:SIN

[202] Anonymous. Subject index, numbers 1–20. Bulletin for the History of Chemistry, (20):74–86, ???? 1997. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num20/num20%20p74-86.pdf.

Brock:1998:DAA

[203] William H. Brock. The 1995 Dexter Award Address. The chemical origins of practical physics. Bulletin for the History of Chemistry, (21):1–11, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p1-11.pdf.

Buonora:1998:AMM

[204] Paul T. Buonora. Almer Mcduffie Mcafee (1886–1972): Commercial catalytic cracking pioneer. Bulletin for the History of Chemistry, (21):12–18, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p12-18.pdf.

Creese:1998:EWC

[205] Mary R. S. Creese. Early women chemists in Russia: Anna Volkova, Iuliia Lermontova and Nadezhda Ziber-Shumova. Bulletin for the History of Chemistry, (21):19–24, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p19-24.pdf.

Gortler:1998:PDB

[206] Leon B. Gortler and Martin D. Saltzman. Paul D. Bartlett’s Chem 17a: a pioneer course in physical organic chemistry. Bulletin for the History of Chemistry, (21):25–31, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p25-31.pdf.

Moore:1998:PMP

[207] Carl E. Moore and Bruno Jaselskis. The ph meter, a product of technological crossovers. Bulletin for the History of Chemistry, (21):32–37, ???? 1998. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num21/num21%20p32-37.pdf.

Pallo:1998:MPE

[208] Gabor Pallo. Michael Polanyi’s early years in science. Bulletin for the History of Chemistry, (21):39–43, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p39-43.pdf.

Stock:1998:MBS

[209] John T. Stock. Max Le Blanc’s studies on electrolytic polarization. Bulletin for the History of Chemistry, (21):44–50, ???? 1998. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num21/num21%20p44-50.pdf.

Swan:1998:MJP

[210] Patricia B. Swan and Kenneth J. Carpenter. Myer E. Jaffa: Pioneering chemist in the food and nutrition sciences. Bulletin for the History of Chemistry, (21):51–57, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p51-57.pdf.

Anonymous:1998:BNE

[211] Anonymous. Book notes: Enough for One Lifetime: Wallace Carothers, Inventor of Nylon; Arrhenius. from Ionic Theory to the Greenhouse Effect; Edward Frankland: Chemistry, Controversy and Conspiracy in Victorian England; A Life as Lived; Eilhard Mitscherlich. Prince of Prussian Chemistry [Translation in English of Eilhard Mitscherlich: Baumeister am Fundament der Chemie, Deutsches Museum, Munchen, 1992. See Bull. Hist. Chem. 1992–93, No. 13–14, p70 for review]. Bulletin for the History of Chemistry, (21):58–61, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num21/num21%20p58-61.pdf.

Laidler:1998:DAA

[212] Keith J. Laidler. The 1996 Dexter Award Address: Contrasts in chemical style: Sidgwick and Eyring. Bulletin for the History of Chemistry, (22):1–9, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num22/num22%20p1-9.pdf.

Saltzman:1998:MLO

[213] Martin D. Saltzman. Morris Loeb: Ostwald’s first American student and America’s first physical chemist. Bulletin for the History of Chemistry, (22):10–15, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num22/num22%20p10-15.pdf.

Stock:1998:LHE

[214] John T. Stock. Latent heat and electrode potential. Bulletin for the History of Chemistry, (22):16–19, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num22/num22%20p16-19.pdf.

Anonymous:1998:CND

[215] Anonymous. Call for nomination for Dexter Award for 1999. Bulletin for the History of Chemistry, (22):20, ???? 1998. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num22/num22%20p20.pdf.

Bader:1998:WLC

[216] Alfred Bader. The Wiswesser–Loschmidt connection. Bulletin for the History of Chemistry, (22):21–28, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num22/num22%20p21-28.pdf.

Warnhoff:1998:WPW

[217] Edgar Warnhoff. When piperidine was a structural problem. Bulletin for the History of Chemistry, (22):29–34, ???? 1998. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/num22/num22%20p29-34.pdf.

Anonymous:1998:BNT

[218] Anonymous. Book notes: Traces of the Past: Unraveling the Secrets of Archaeology through Chemistry; Justus von Liebig: the Chemical Gatekeeper; A History of Metallurgy; Instruments of Science: an Historical Encyclopedia; Chemistry and the Chemical Industry in the 19th Century: the Henrys of Manchester and other Studies. Bulletin for the History of Chemistry, (22):35–40, ???? 1998. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num22/num22%20p35-40.pdf.

Bensaude-Vincent:1999:DAA

[219] Bernadette Bensaude-Vincent. The 1997 Dexter Award Address: a language to order the chaos. Bulletin for the History of Chemistry, (23):1–10, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num23/num23%20p1-10.pdf.

Caswell:1999:DBT

[220] Lyman R. Caswell and Rebecca W. Stone Daley. The Delhuyar brothers, tungsten, and Spanish silver. Bulletin for the History of Chemistry, (23):11–19, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num23/num23%20p11-19.pdf.

Rayner-Canham:1999:BWC

[221] Marelene F. Rayner-Canham and Geoffrey W. Rayner-Canham. British women chemists and the First World War. Bulletin for the History of Chemistry, (23):20–27, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num23/num23%20p20-27.pdf.

Riethmiller:1999:EBA

[222] Steven Riethmiller. Erlich, Bertheim, and atoxyl: the origins of modern chemotherapy. Bulletin for the History of Chemistry, (23):28–33, ???? 1999. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num23/num23%20p28-33.pdf.

Saltzman:1999:AIW

[223] Martin D. Saltzman. Academia and industry: What should their relationship be? The Levinstein–Roscoe dialog. Bulletin for the History of Chemistry, (23):34–41, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num23/num23%20p34-41.pdf.

Stock:1999:KRP

[224] John T. Stock. The key role played by sugar in early experiments in kinetics and equilibria. Bulletin for the History of Chemistry, (23):42–46, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num23/num23%20p42-46.pdf.

Williams:1999:GHL

[225] William D. Williams. Gustavus Hinrichs and the Lavoisier Monument. Bulletin for the History of Chemistry, (23):47–49, ???? 1999. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num23/num23%20p47-49.pdf.

Anonymous:1999:BNC

[226] Anonymous. Book notes: From Caveman to Chemist: Circumstances and Achievements; Cavendish; Episodes from the History of the Rare Earth Elements; Paracelsus: Das Werk – die Rezeption; John Dalton, 1776– 1844: a Bibliography of Works by and About Him, with an Annotated List of his Surviving Apparatus and Personal Effects. Bulletin for the History of Chemistry, (23):50–56, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num23/num23%20p50-56.pdf.

Mauskopf:1999:DAA

[227] Seymour H. Mauskopf. The 1998 Dexter Award Address: “From an instrument of war to an instrument of the laboratory”, the affinities certainly do not change chemists and the development of munitions, 1785– 1885. Bulletin for the History of Chemistry, (24):1–15, ???? 1999. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num24/num24%20p1-15.pdf.

Adams:1999:SPM

[228] David L. Adams. Samuel Parsons Mulliken: Pioneer in organic qualitative analysis. Bulletin for the History of Chemistry, (24):16–23, ???? 1999. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/num24/num24%20p16-23.pdf.

Giunta:1999:JRN

[229] Carmen J. Giunta. J. A. R. Newlands’ classification of the elements: Periodicity, but no system. Bulletin for the History of Chemistry, (24):24–31, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num24/num24%20p24-31.pdf. See commentary [276] and

response [277].

Habashi:1999:CEG

[230] Fathi Habashi. Christlieb Ehregott Gellert and his metallurgic chymistry. Bulletin for the History of Chemistry, (24):32–39, ???? 1999. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num24/num24%20p32-39.pdf.

Rosenfeld:1999:OFD

[231] Louis Rosenfeld. Otto Folin and Donald D. Van Slyke: Pioneers of clinical chemistry. Bulletin for the History of Chemistry, (24):40–47, ???? 1999. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num24/num24%20p40-47.pdf.

Saltzman:1999:FPA

[232] Martin D. Saltzman. François-Pierre Ami Argand: Let there be light. Bulletin for the History of Chemistry, (24):48–52, ???? 1999. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/num24/num24%20p48-52.pdf.

Saltzman:1999:ADD

[233] Martin D. Saltzman. The art of distillation and the dawn of the hydrocarbon society. Bulletin for the History of Chemistry, (24):53–60, ???? 1999. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/num24/num24%20p53-60.pdf.

Weller:1999:NBF

[234] Sol W. Weller. Napoleon Bonaparte, French scientists, chemical equilibrium and mass action. Bulletin for the History of Chemistry, (24):61–65, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num24/num24%20p61-65.pdf.

Williams:1999:DHL

[235] William D. Williams. Dobereiner’s hydrogen lighter. Bulletin for the History of Chemistry, (24):66–68, ???? 1999. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/num24/num24%20p66-68.pdf.

Anonymous:1999:BNH

[236] Anonymous. Book notes: A History of Chemistry, Originally published as Histoire de la chemie, Editions La Découverte, 1993. The Making of the Chemist: the Social History of Chemistry in Europe, 1789–1914; The Aspiring Adept: Robert Boyle and His Alchemical Quest; Women in Chemistry: Their Changing Roles from Alchemical Times to the Mid-Twentieth Century; Fritz Haber: Chemiker, Nobelpreisträger, Deutscher, Jude. Bulletin for the History of Chemistry, (24):69–78, ???? 1999. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/num24/num24%20p69-78.pdf.

Gortler:2000:MAF

[237] Leon Gortler. Merck in America the first 70 years: from fine chemicals to pharmaceutical giant. Bulletin for the History of Chemistry, 25(1):1–9, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-1/v25-1%20p1-9.pdf.

Lombardino:2000:BHP

[238] Joseph G. Lombardino. A brief history of Pfizer Central Research. Bulletin for the History of Chemistry, 25(1):10–15, ???? 2000. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v25-1/v25-1%20p10-15.pdf.

Ullyot:2000:MSK

[239] Glenn E. Ullyot, Barbara Hodsdon Ullyot, and Leo B. Slater. The metamorphosis of Smith–Kline & French Laboratories to Smith Kline Beecham: 1925–1998. Bulletin for the History of Chemistry, 25(1):16– 20, ???? 2000. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v25-
1/v25-1%20p16-20.pdf.

Collins:2000:LCS

[240] Joseph C. Collins and John R. Gwilt. The life cycle of Sterling Drug, Inc. Bulletin for the History of Chemistry, 25(1):22–27, ???? 2000. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v25-1/v25-1%20p22-27.pdf.

Hoefle:2000:EHP

[241] Milton L. Hoefle. The early history of Parke–Davis and Company. Bulletin for the History of Chemistry, 25(1):28–34, ???? 2000. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v25-1/v25-1%20p28-34.pdf.

Pallo:2000:HPI

[242] Gabor Pallo. The Hungarian phenomenon in Israeli science. Bulletin for the History of Chemistry, 25(1):35–42, ???? 2000. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v25-1/v25-1%20p35-42.pdf.

Berneking:2000:CHB

[243] Carolyn Bailey Berneking. The contributions of E. H. S. Bailey to the development of pure food and water laws in Kansas. Bulletin for the History of Chemistry, 25(1):43–49, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-1/v25-1%20p43-49.pdf.

Karpenko:2000:EJA

[244] Vladimir Karpenko. Die Edelgeborne Jungfer Alchymia: the final stage of European alchemy. Bulletin for the History of Chemistry, 25(1):50–63, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-1/v25-1%20p50-63.pdf.

Anonymous:2000:ASR

[245] Anonymous. Author Saltzman responds. Bulletin for the History of Chemistry, 25(1):67, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-1/v25-1%20p67.pdf.

Anonymous:2000:BRa

[246] Anonymous. Book reviews. Bulletin for the History of Chemistry, 25(1):68–71, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-1/v25-1%20p68-71.pdf.

Nye:2000:DAA

[247] Mary Jo Nye. The 1999 Dexter Award Address: a place in history: Was Linus Pauling a revolutionary chemist? Bulletin for the History of Chemistry, 25(2):73–82, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p73-82.pdf.

Kragh:2000:CHP

[248] Helge Kragh and Malene M. Bak. Christopher H. Pfaff and the controversy over voltaic electricity. Bulletin for the History of Chemistry, 25(2):83–90, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p83-90.pdf.

Stock:2000:GAH

[249] John T. Stock. George Augustus Hulett: from liquid crystals to standard cell. Bulletin for the History of Chemistry, 25(2):91–97, ???? 2000. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v25-2/v25-2%20p91-97.pdf.

Anonymous:2000:LE

[250] Anonymous. Letter to the Editor. Bulletin for the History of Chemistry, 25(2):98, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p98.pdf.

Stock:2000:MLW

[251] John T. Stock. Morris Loeb, Walther Nernst, and the transference number. Bulletin for the History of Chemistry, 25(2):99–102, ???? 2000. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v25-2/v25-2%20p99-102.pdf.

Rayner-Canham:2000:SHE

[252] Marelene F. Rayner-Canham and Geoffrey W. Rayner-Canham. Stefanie Horovitz, Ellen Gleditsch, Ada Hitchins and the discovery of isotopes. Bulletin for the History of Chemistry, 25(2):103–108, ???? 2000. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v25-2/v25-2%20p103-108.pdf.

Habashi:2000:ZTF

[253] Fathi Habashi. Zoroaster and the theory of four elements. Bulletin for the History of Chemistry, 25(2):109–115, ???? 2000. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v25-2/v25-2%20p109-115.pdf.

Saltzman:2000:SBC

[254] Martin D. Saltzman. Is science a brotherhood? The case of Siegried Ruhemann. Bulletin for the History of Chemistry, 25(2):116–121, ???? 2000. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v25-
2/v25-2%20p116-121.pdf.

Anonymous:2000:DST

[255] Anonymous. D. Stanley Tarbell (1913–1999): a memorial tribute. Bulletin for the History of Chemistry, 25(2):122, ???? 2000. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v25-2/v25-2%20p122.pdf.

Weininger:2000:WND

[256] Stephen J. Weininger. “What’s in a name?” from designation to denunciation: the nonclassical cation controversy. Bulletin for the History of Chemistry, 25(2):123–131, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p123-131.pdf.

Anonymous:2000:BRb

[257] Anonymous. Book reviews. Bulletin for the History of Chemistry, 25(2):132–141, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p132-141.pdf.

Anonymous:2000:AI

[258] Anonymous. Author index, 1–25. Bulletin for the History of Chemistry, 25(2):142–143, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p142-143.pdf.

Anonymous:2000:SI

[259] Anonymous. Subject index, 1–25. Bulletin for the History of Chemistry, 25(2):144–168, ???? 2000. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v25-2/v25-2%20p144-168.pdf.

Anonymous:2001:AJI

[260] Anonymous. Aaron John Ihde 1909–2000. Bulletin for the History of Chemistry, 26(1):1–2, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-1/v26-1%20p1-2.pdf.

Bohning:2001:AIL

[261] James J. Bohning. Aaron Ihde: a life from Bascom’s Hill. Bulletin for the History of Chemistry, 26(1):3–14, ???? 2001. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v26-1/v26-1%20p3-14.pdf.

Rocke:2001:AIH

[262] Alan J. Rocke. Aaron Ihde and his students (1). Bulletin for the History of Chemistry, 26(1):15–23, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-1/v26-1%20p15-23.pdf.

Jensen:2001:AIC

[263] William B. Jensen. Aaron Ihde’s contributions to the history of chemistry. Bulletin for the History of Chemistry, 26(1):24–32, ???? 2001. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v26-1/v26-1%20p24-32.pdf.

Siegfried:2001:MAI

[264] Robert Siegfried. Memories of Aaron Ihde. Bulletin for the History of Chemistry, 26(1):33–35, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-1/v26-1%20p33-35.pdf.

Ihde:2001:CGR

[265] Aaron J. Ihde. Criteria for genealogical roots (1). Bulletin for the History of Chemistry, 26(1):36–39, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-1/v26-1%20p36-39.pdf.

Rubin:2001:HOS

[266] Mordecai B. Rubin. The history of ozone: the Schonbein period, 1839– 1868. Bulletin for the History of Chemistry, 26(1):40–56, ???? 2001. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v26-1/v26-1%20p40-56.pdf.

Oltra:2001:CPL

[267] Lluis garrigos Oltra, Carles-Millan Verdu, and Georgina-Blanes Nadal. The contributions of Payen and Labillardiere to the development of colorimetry. Bulletin for the History of Chemistry, 26(1):57–65, ???? 2001. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v26-
1/v26-1%20p57-65.pdf.

Page:2001:BTW

[268] Frederick G. Page. The birth of titrimetry: William Lewis and the analysis of American potashes. Bulletin for the History of Chemistry, 26(1):66–72, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-1/v26-1%20p66-72.pdf.

Anonymous:2001:BRa

[269] Anonymous. Book reviews. Bulletin for the History of Chemistry, 26(1):73–80, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-1/v26-1%20p73-80.pdf.

Rocke:2001:DAA

[270] Alan J. Rocke. The 2000 Dexter Award Address: Celebrity culture in Parisian chemistry. Bulletin for the History of Chemistry, 26(2):81– 91, ???? 2001. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v26-
2/v26-2%20p81-91.pdf.

Bohning:2001:OFA

[271] James J. Bohning. Oppposition to the formation of the American Chemical Society. Bulletin for the History of Chemistry, 26(2):92– 103, ???? 2001. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v26-
2/v26-2%20p92-103.pdf. See note [287].

Habashi:2001:NSV

[272] Fathi Habashi. Niepce De Saint-Victor and the discovery of radioactivity. Bulletin for the History of Chemistry, 26(2):104–105, ???? 2001. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v26-2/v26-2%20p104-105.pdf.

Saltzman:2001:HCC

[273] Martin D. Saltzman. The Hare–Clarke controversy over the invention of the improved gas blowpipe. Bulletin for the History of Chemistry, 26(2):106–111, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-2/v26-2%20p106-111.pdf.

Saltzman:2001:CCM

[274] Martin D. Saltzman. The contributions of Cyrus Moors Warren to the analysis of hydrocarbons. Bulletin for the History of Chemistry, 26(2):112–117, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-2/v26-2%20p112-117.pdf.

Stock:2001:RSI

[275] John T. Stock. Reciprocal solubility influence in salt mixtures: the contributions of Walther Nernst and of Arthur Noyes. Bulletin for the History of Chemistry, 26(2):118–123, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-2/v26-2%20p118-123.pdf.

Scerri:2001:PCG

[276] Eric R. Scerri. A philosophical commentary on Giunta’s critique of Newlands’ classification of the elements. Bulletin for the History of Chemistry, 26(2):124–129, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-2/v26-2%20p124-129.pdf. See [229, 277].

Giunta:2001:RSC

[277] Carmen J. Giunta. A response to Scerri’s commentary. Bulletin for the History of Chemistry, 26(2):130–132, ???? 2001. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v26-2/v26-2%20p130-132.pdf. See [276, 229].

Anonymous:2001:CNE

[278] Anonymous. Call for nominations for the Edelstein Award 2002. Bulletin for the History of Chemistry, 26(2):133, ???? 2001. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v26-2/v26-2%20p133.pdf.

Anonymous:2001:BRb

[279] Anonymous. Book reviews. Bulletin for the History of Chemistry, 26(2):134–142, ???? 2001. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v26-2/v26-2%20p134-142.pdf.

Anonymous:2002:DAA

[280] Anonymous. Dexter Award 2001: Awarded posthumouly to William A. Smeaton. Bulletin for the History of Chemistry, 27(1):1, ???? 2002. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v27-1/v27-1%20p1.pdf.

Rice:2002:CEC

[281] Richard E. Rice. Cutting-edge chemistry: Some 19th-century Russian contributions. Bulletin for the History of Chemistry, 27(1):2–3, ???? 2002. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/v27-1/v27-1%20p2-3.pdf.

See commentary [286].

Kaji:2002:DMC

[282] Masanori Kaji. D. I. Mendeleev’s concept of chemical elements and the principles of chemistry. Bulletin for the History of Chemistry, 27(1):4–16, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-1/v27-1%20p4-16.pdf.

Rice:2002:HIS

[283] Richard E. Rice. Hydrating ions in St. Petersburg and Moscow: Ignoring them in Leipzig and Baltimore. Bulletin for the History of Chemistry, 27(1):17–25, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-1/v27-1%20p17-25.pdf.

Brooks:2002:NZS

[284] Nathan M. Brooks. Nikolai Zinin and synthetic dyes: the road not taken. Bulletin for the History of Chemistry, 27(1):26–36, ???? 2002. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v27-1/v27-1%20p26-36.pdf.

Lewis:2002:BSO

[285] David L. Lewis. The beginnings of synthetic organic chemistry: Zinc alkyls and the Kazan’ School. Bulletin for the History of Chemistry, 27(1):37–42, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-1/v27-1%20p37-42.pdf.

Mauskopf:2002:CEC

[286] Seymour H. Mauskopf. Cutting-edge chemistry: Some 19th-century Russian contributions: a commentary. Bulletin for the History of Chemistry, 27(1):43–45, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-1/v27-1%20p43-45.pdf. See [281].

Bohning:2002:OFA

[287] James J. Bohning. Opposition to the formation of the American Chemical Society: Note added in proof. Bulletin for the History of Chemistry, 27(1):46–47, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-1/v27-1%20p46-47.pdf. See [271].

Craig:2002:FGC

[288] Norman C. Craig. Four giants in a college chemistry department: Oberlin College, 1880–1966. Bulletin for the History of Chemistry, 27(1): 48–56, ???? 2002. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v27-
1/v27-1%20p48-56.pdf.

Stock:2002:EBF

[289] John T. Stock. Edgar Buckingham: Fluorescence of quinine salts. Bulletin for the History of Chemistry, 27(1):57–61, ???? 2002. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v27-1/v27-1%20p57-61.pdf.

Giunta:2002:DPC

[290] Carmen J. Giunta. Dulong and Petit: a case of data fabrication? Bulletin for the History of Chemistry, 27(1):62–71, ???? 2002. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v27-1/v27-1%20p62-71.pdf.

Anonymous:2002:BRa

[291] Anonymous. Book reviews. Bulletin for the History of Chemistry, 27(1):72–80, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-1/v27-1%20p72-80.pdf.

Rubin:2002:HOI

[292] Mordecai B. Rubin. The history of ozone. II. 1869–1899 (1). Bulletin for the History of Chemistry, 27(2):81–106, ???? 2002. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v27-2/v27-2%20p81-106.pdf.

Page:2002:FHJ

[293] Frederick G. Page. Francis Home and Joseph Black: the chemistry and testing of alkaline salts in the early bleaching and alkali trade. Bulletin for the History of Chemistry, 27(2):107–113, ???? 2002. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v27-2/v27-2%20p107-113.pdf. See photograph

caption erratum [305].

Moore:2002:OGC

[294] Carl E. Moore, Alfred von Smolinski, and Bruno Jaselskis. The Ostwald–Gibbs correspondence: an interesting component in the history of the energy concept. Bulletin for the History of Chemistry, 27(2):114–127, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-2/v27-2%20p114-127.pdf.

Ede:2002:NDS

[295] Andrew Ede. The natural defense of a scientific people: the public debate over chemical warfare in post-WWI America. Bulletin for the History of Chemistry, 27(2):128–135, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-2/v27-2%20p128-135.pdf.

Lewis:2002:CR

[296] Edward S. Lewis. Chemistry at rice, 1912 to 1998. Bulletin for the History of Chemistry, 27(2):136–144, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-2/v27-2%20p136-144.pdf.

Anonymous:2002:BRb

[297] Anonymous. Book reviews. Bulletin for the History of Chemistry, 27(2):145–152, ???? 2002. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v27-2/v27-2%20p145-152.pdf.

Parascandola:2003:EAA

[298] John Parascandola. The 2002 Edelstein Award Address: To bond or not to bond: Chemical versus physical theories of drug action. Bulletin for the History of Chemistry, 28(1):1–8, ???? 2003. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v28-1/v28-1%20p1-8.pdf.

Higby:2003:CCA

[299] Gregory J. Higby. Chemistry and the 19th-century American pharmacist. Bulletin for the History of Chemistry, 28(1):9–17, ???? 2003. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v28-1/v28-1%20p9-17.pdf.

May:2003:EDC

[300] Leopold May. The early days of chemistry at Catholic University. Bulletin for the History of Chemistry, 28(1):18–25, ???? 2003. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v28-1/v28-1%20p18-25.pdf.

Takacs:2003:MCL

[301] Laszlo Takacs. M. Carey Lea, the father of mechanochemistry. Bulletin for the History of Chemistry, 28(1):26–34, ???? 2003. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v28-1/v28-1%20p26-34.pdf.

Caswell:2003:ADR

[302] Lyman R. Caswell. Andres del Rio, Alexander von Humboldt, and the twice-discovered element. Bulletin for the History of Chemistry, 28(1):35–41, ???? 2003. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v28-1/v28-1%20p35-41.pdf.

Burk:2003:FAH

[303] Lale Aka Burk. Fritz Arndt and his chemistry books in the Turkish language. Bulletin for the History of Chemistry, 28(1):42–53, ???? 2003. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v28-1/v28-1%20p42-53.pdf.

Anonymous:2003:BRa

[304] Anonymous. Book reviews. Bulletin for the History of Chemistry, 28(1):54–63, ???? 2003. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v28-1/v28-1%20p54-63.pdf.

Anonymous:2003:E

[305] Anonymous. Erratum. Bulletin for the History of Chemistry, 28(1): 63, ???? 2003. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v28-
1/v28-1%20p65.pdf. See [293].

Cintas:2003:FBA

[306] Pedro Cintas. Francis Bacon: an alchemical odyssey through the Novum Organum. Bulletin for the History of Chemistry, 28(2):65– 75, ???? 2003. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v28-
2/v28-2%20p65-75.pdf.

Marshall:2003:ERT

[307] James L. Marshall and Virginia R. Marshall. Ernest Rutherford, the “true discoverer” of radon. Bulletin for the History of Chemistry, 28(2):76–83, ???? 2003. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v28-2/v28-2%20p76-83.pdf. See comments and response

[321, 322].

Saltzman:2003:JBC

[308] Martin Saltzman. James Bryant Conant: the making of an iconoclastic chemist. Bulletin for the History of Chemistry, 28(2):84–94, ???? 2003. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v28-
2/v28-2%20p84-94.pdf.

Stock:2003:ASC

[309] John T. Stock. Arthur Slator and the chlorination of benzene. Bulletin for the History of Chemistry, 28(2):95–100, ???? 2003. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v28-2/v28-2%20p95-100.pdf.

Larson:2003:RCO

[310] Clarence E. Larson. The role of chemistry in the Oak Ridge Electromagnetic Project. Bulletin for the History of Chemistry, 28(2):101– 109, ???? 2003. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v28-
2/v28-2%20p101-109.pdf.

Rayner-Canham:2003:PDF

[311] Marlene F. Rayner-Canham and Geoffrey W. Rayner-Canham. Pounding on the doors: the fight for acceptance of British women chemists. Bulletin for the History of Chemistry, 28(2):110–119, ???? 2003. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v28-2/v28-2%20p110-119.pdf.

Anonymous:2003:BRb

[312] Anonymous. Book reviews. Bulletin for the History of Chemistry, 28(2):120–128, ???? 2003. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v28-2/v28-2%20p120-128.pdf.

Knight:2004:EAA

[313] David Knight. The 2003 Edelstein Award Address: Making chemistry popular. Bulletin for the History of Chemistry, 29(1):1–8, ???? 2004. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/v29-1/v29-1%20p1-8.pdf.

Sourkes:2004:DLF

[314] Theodore L. Sourkes. The discovery of lecithin, the first phospholipid. Bulletin for the History of Chemistry, 29(1):9–15, ???? 2004. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v29-1/v29-1%20p9-15.pdf.

Stock:2004:GLC

[315] John T. Stock. Gabriel Lippmann and the capillary electrometer. Bulletin for the History of Chemistry, 29(1):16–20, ???? 2004. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v29-1/v29-1%20p16-20.pdf.

Cohen:2004:KCL

[316] Stephen M. Cohen. Khemye: Chemical literature in Yiddish. Bulletin for the History of Chemistry, 29(1):21–29, ???? 2004. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v29-1/v29-1%20p21-29.pdf.

Shine:2004:EHC

[317] Henry J. Shine. An early history of chemistry at Texas Tech University: 1925–1970*. Bulletin for the History of Chemistry, 29(1):30– 44, ???? 2004. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v29-
1/v29-1%20p30-44.pdf.

McGrayne:2004:NLA

[318] Sharon Bertsch McGrayne. Noyes Laboratory, an ACS national chemical landmark: 100 years of chemistry at the University of Illinois. Bulletin for the History of Chemistry, 29(1):45–51, ???? 2004. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v29-1/v29-1%20p45-51.pdf.

Anonymous:2004:BRa

[319] Anonymous. Book reviews. Bulletin for the History of Chemistry, 29(1):52–63, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-1/v29-1%20p52-63.pdf.

Kurzer:2004:CLD

[320] Frederick Kurzer. Chemistry in the life of Dr. Samuel Johnson. Bulletin for the History of Chemistry, 29(2):65–88, ???? 2004. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v29-2/v29-2%20p65-88.pdf.

Rayner-Canham:2004:RTD

[321] Marlene F. Rayner-Canham and Geoffrey W. Rayner-Canham. Rutherford, the “true discoverer of radon”. Bulletin for the History of Chemistry, 29(2):89–90, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-2/v29-2%2520p89-90.pdf. See [307, 322].

Marshall:2004:R

[322] James L. Marshall and Virginia R. Marshall. Reply. Bulletin for the History of Chemistry, 29(2):90, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-2/v29-2/v29-2%20p90.pdf. See [321, 307].

Stock:2004:FGD

[323] John T. Stock. Frederick George Donnan and the relationship between electrolytic dissociation and light absorption. Bulletin for the History of Chemistry, 29(2):91–96, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-2/v29-2%20p91-96.pdf.

Habashi:2004:CFE

[324] Fathi Habashi. Cambodia’s four elements. Bulletin for the History of Chemistry, 29(2):97–98, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-2/v29-2%20p97-98.pdf.

Rubin:2004:HOI

[325] Mordecai B. Rubin. The history of ozone. IV. The isolation of pure ozone and determination of its physical properties. Bulletin for the History of Chemistry, 29(2):99–106, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-2/v29-2%20p99-106.pdf.

Anonymous:2004:HCC

[326] Anonymous. Hist citations for chemical breakthroughs. Bulletin for the History of Chemistry, 29(2):107, ???? 2004. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v29-2/v29-2%20p107.pdf.

Anonymous:2004:BRb

[327] Anonymous. Book reviews. Bulletin for the History of Chemistry, 29(2):108–119, ???? 2004. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v29-2/v29-2%20p108-119.pdf.

Anonymous:2004:CNE

[328] Anonymous. Call for nominations for the Edelstein Award for 2005. Bulletin for the History of Chemistry, 29(2):120, ???? 2004. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v29-2/v29-2%20p120.pdf.

Lambert:2005:EAA

[329] Jospeh B. Lambert. The 2004 Edelstein Award Address: the deep history of chemistry. Bulletin for the History of Chemistry, 30(1):1–9, ???? 2005. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/v30-1/v30-1%20p1-9.pdf.

Rae:2005:GPL

[330] Ian D. Rae. Granville Perkins and leprosy chemotherapy in the Phillipines. Bulletin for the History of Chemistry, 30(1):10–18, ???? 2005. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v30-1/v30-1%20p10-18.pdf.

Behrman:2005:KE

[331] E. J. Behrman. Karl Elbs, 1858–1933. Bulletin for the History of Chemistry, 30(1):19–22, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-1/v30-1%20p19-22.pdf.

Russo:2005:HSM

[332] Ruth Russo. The heart of steel: a metallurgical interpretation of iron in Homer. Bulletin for the History of Chemistry, 30(1):23–29, ???? 2005. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v30-
1/v30-1%20p23-29.pdf.

Schwenk:2005:FRH

[333] Ernst F. Schwenk. Friedlieb Runge and his capillary designs. Bulletin for the History of Chemistry, 30(1):30–34, ???? 2005. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v30-1/v30-1%20p30-34.pdf.

Stock:2005:JHD

[334] John T. Stock. John Hughes Davies. The effect of the silent electrical discharge on the ammonia system. Bulletin for the History of Chemistry, 30(1):35–40, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-1/v30-1%20p35-40.pdf.

Egolf:2005:HCE

[335] Roger A. Egolf. The history of chemical education at Lafayette College. Bulletin for the History of Chemistry, 30(1):41–50, ???? 2005. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v30-1/v30-1%20p41-50.pdf.

Anonymous:2005:BRa

[336] Anonymous. Book reviews. Bulletin for the History of Chemistry, 30(1):51–56, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-1/v30-1%20p51-56.pdf.

Schofield:2005:JPN

[337] Robert E. Schofield. Joseph Priestley, natural philosopher. Bulletin for the History of Chemistry, 30(2):57–62, ???? 2005. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v30-2/v30-2%20p57-62.pdf.

Woodcock:2005:PTC

[338] Leslie V. Woodcock. Phlogistion theory and chemical revolutions. Bulletin for the History of Chemistry, 30(2):63–69, ???? 2005. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v30-2/v30-2%20p63-69.pdf.

White:2005:PPP

[339] J. Edmund White. Pantisocracy and Pennsylvania: Plans of Coleridge and Southey and of Cooper and Priestley. Bulletin for the History of Chemistry, 30(2):70–76, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-2/v30-2%20p70-76.pdf.

Neeley:2005:ERR

[340] Kathleen L. Neeley and M. Andrea Bashore. Esteem, regard, and respect for rationality: Joseph Priestley’s female connections. Bulletin for the History of Chemistry, 30(2):77–90, ???? 2005. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v30-2/v30-2%20p77-90.pdf.

Ginsberg:2005:PJA

[341] Judah B. Ginsberg. Priestley, Jefferson, and Adams: the emigré and American politics. Bulletin for the History of Chemistry, 30(2):91– 102, ???? 2005. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v30-
2/v30-2%20p91-102.pdf.

Swain:2005:BCF

[342] Patricia A. Swain. Bernard Courtois (1777–1838), famed for discovering iodine (1811) and his life in Paris from 1798. Bulletin for the History of Chemistry, 30(2):103–111, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-2/v30-2%20p103-111.pdf.

Anonymous:2005:CNE

[343] Anonymous. Call for nominations for the Edelstein Award for 2006. Bulletin for the History of Chemistry, 30(2):112, ???? 2005. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v30-2/v30-2%20p112.pdf.

Stock:2005:DHP

[344] John T. Stock and James D. Stuart. The decomposition of hydrogen peroxide by blood: George Senter’s discovery of the enzyme involved. Bulletin for the History of Chemistry, 30(2):113–117, ???? 2005. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v30-2/v30-2%20p113-117.pdf.

Anonymous:2005:BRb

[345] Anonymous. Book reviews. Bulletin for the History of Chemistry, 30(2):118–124, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-2/v30-2%20p118-124.pdf.

Anonymous:2005:AI

[346] Anonymous. Author index, 1–30. Bulletin for the History of Chemistry, 30(2):125–126, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-2/v30-2%20p125-126.pdf.

Anonymous:2005:SI

[347] Anonymous. Subject index, 1–30. Bulletin for the History of Chemistry, 30(2):126–152, ???? 2005. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v30-2/v30-2%20p126-152.pdf.

Jensen:2006:EAA

[348] William B. Jensen. The 2005 Edelstein Award Address: Textbooks and the future of the history of chemistry as an academic discipline. Bulletin for the History of Chemistry, 31(1):1–8, ???? 2006. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v31-1/v31-1%20p1-8.pdf.

Murphy:2006:CJB

[349] Clarence J. Murphy. Charles James, B. Smith Hopkins, and the tangled web of element 61 [Pm (promethium)]. Bulletin for the History of Chemistry, 31(1):9–18, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-1/v31-1%20p9-18.pdf.

Saltzman:2006:CST

[350] Martin D. Saltzman. Chemical societies and their journals: What can be learned about the state of chemistry from an analysis of the first volumes of these journals. Bulletin for the History of Chemistry, 31(1):19–27, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-1/v31-1%20p19-27.pdf.

Dayan:2006:CKM

[351] Anthony D. Dayan. The circumstances of Kekulé’s molecular dream in London in 1854. Bulletin for the History of Chemistry, 31(1):28– 30, ???? 2006. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v31-
1/v31-1%20p28-30.pdf.

Anonymous:2006:LE

[352] Anonymous. Letter to the Editor. Bulletin for the History of Chemistry, 31(1):31, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-1/v31-1%20p31%20letter%20to%20editor.pdf.

Anonymous:2006:MJS

[353] Anonymous. Message from Jeffrey I. Seeman: Chair of Hist. Bulletin for the History of Chemistry, 31(1):32, ???? 2006. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v31-1/v31-1%20p32.pdf.

Anonymous:2006:BRa

[354] Anonymous. Book reviews. Bulletin for the History of Chemistry, 31(1):33–42, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-1/v31-1%20p33-42.pdf.

Quilez:2006:RTE

[355] Juan Quilez. The role of theories in early studies of chemical equilibria. Bulletin for the History of Chemistry, 31(2):45–57, ???? 2006. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v31-2/v31-2%20p45-57.pdf.

Jensen:2006:EA

[356] William B. Jensen. Erasmus on alchemy. Bulletin for the History of Chemistry, 31(2):58–65, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-2/v31-2%20p58-65.pdf.

Giunta:2006:TMJ

[357] Carmen J. Giunta. Thomas Midgley, Jr., and the invention of chlorofluorocarbon refrigerants: It ain’t necessarily so. Bulletin for the History of Chemistry, 31(2):66–74, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-2/v31-2%20p66-74.pdf.

Slater:2006:CNE

[358] Leo B. Slater. Chemists and national emergency: NIH’s unit of chemotherapy during World War II. Bulletin for the History of Chemistry, 31(2):75–80, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-2/v31-2%20p75-80.pdf.

Togeas:2006:ERD

[359] James Togeas. Element and radical: the divergence of synonyms. Bulletin for the History of Chemistry, 31(2):81–84, ???? 2006. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v31-2/v31-2%20p81-84.pdf.

Anonymous:2006:CNE

[360] Anonymous. Call for nominations for the 2007 Edelstein Award. Bulletin for the History of Chemistry, 31(2):85, ???? 2006. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v31-2/v31-2%20p85.pdf.

Anonymous:2006:BRb

[361] Anonymous. Book reviews. Bulletin for the History of Chemistry, 31(2):86–91, ???? 2006. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v31-2/v31-2%20p86-91.pdf.

Anonymous:2007:FYD

[362] Anonymous. Fifty years of the Dexter and Edelstein Awards. Bulletin for the History of Chemistry, 32(1):1, ???? 2007. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v32-1/v32-1%20p1.pdf.

Morris:2007:EAA

[363] Peter J. T. Morris. The 2006 Edelstein Award Address: Writing the history of modern chemistry. Bulletin for the History of Chemistry, 32(1):2–9, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-1/v32-1%20p2-9.pdf.

Morris:2007:PHC

[364] Peter J. T. Morris. A personal historiography of the chemical industry since 1956. Bulletin for the History of Chemistry, 32(1):10– 20, ???? 2007. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v32-
1/v32-1%20p10-20.pdf.

Nye:2007:SBH

[365] Mary Jo Nye. Scientific biography in the history of chemistry: the role of Dexter and Edelstein Award winners in the last fifty years. Bulletin for the History of Chemistry, 32(1):21–26, ???? 2007. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v32-1/v32-1%20p21-26.pdf.

Travis:2007:UTT

[366] Anthony S. Travis. Unintended technology transfer: Acetylene chemistry in the United States. Bulletin for the History of Chemistry, 32(1): 27–34, ???? 2007. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v32-
1/v32-1%20p27-34.pdf.

Travis:2007:MA

[367] Anthony S. Travis. Mauve and its anniversaries. Bulletin for the History of Chemistry, 32(1):35–44, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-1/v32-1%20p35-44.pdf.

Anonymous:2007:BRa

[368] Anonymous. Book reviews. Bulletin for the History of Chemistry, 32(1):57–64, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-1/v32-1%20p57-64.pdf.

Anonymous:2007:WIB

[369] Anonymous. Welcome to this issue of the Bulletin. Bulletin for the History of Chemistry, 32(2):65, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p65.pdf.

Bohning:2007:LBE

[370] James J. Bohning. Looking back: Eighty-five years of chemists and their history. Bulletin for the History of Chemistry, 32(2):66–81, ???? 2007. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v32-
2/v32-2%20p66-81.pdf.

Anonymous:2007:HMS

[371] Anonymous. Hist mission statement. Bulletin for the History of Chemistry, 32(2):81, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p81.pdf.

Jones:2007:CHC

[372] Paul R. Jones. Communicating the history of chemistry. Bulletin for the History of Chemistry, 32(2):82–86, ???? 2007. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v32-2/v32-2%20p82-86.pdf.

Seeman:2007:IHH

[373] Jeffrey I. Seeman. Influences of hist and the history of chemisty on the course of chemistry, examples of synergy. Bulletin for the History of Chemistry, 32(2):87–96, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p87-96.pdf.

Anonymous:2007:HSU

[374] Anonymous. Hist 2007 state-of-the-union. Bulletin for the History of Chemistry, 32(2):97, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p97.pdf.

Giunta:2007:LAK

[375] Carmen J. Giunta. Looking ahead: Keeping history of chemistry relevant to the future of chemistry. Bulletin for the History of Chemistry, 32(2):98–103, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p98-103.pdf.

Jensen:2007:LE

[376] William B. Jensen. Letter to the Editor. Bulletin for the History of Chemistry, 32(2):104, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p104.pdf.

Deichmann:2007:MVC

[377] Ute Deichmann. “Molecular” versus “colloidal”: Controversies in biology and biochemistry, 1900–1940. Bulletin for the History of Chemistry, 32(2):105–118, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p105-118.pdf.

Jaselskis:2007:TGT

[378] Bruno Jaselskis, Carl E. Moore, and Alfred von Smolinsk. Theodor Von Grotthuss (1785–1822) – a trail blazer. Bulletin for the History of Chemistry, 32(2):119–128, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p119-128.pdf.

Olsen:2007:CWG

[379] Kevin Olsen. Clear waters and a green gas: a history of chlorine as a swimming pool sanitizer in the United States. Bulletin for the History of Chemistry, 32(2):129–140, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p129-140.pdf.

Anonymous:2007:BRb

[380] Anonymous. Book reviews. Bulletin for the History of Chemistry, 32(2):141–147, ???? 2007. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v32-2/v32-2%20p141-147%20book%20reviews.pdf.

Travis:2008:EAP

[381] Anthony S. Travis. 2007 Edelstein Award paper: What a wonderful empire is the organic chemistry. Bulletin for the History of Chemistry, 33(1):1–11, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-1/v33-1%20p1-11.pdf.

Raos:2008:LSA

[382] Nenad Raos. Letters of Svante Arrhenius to his former Croatian student. Bulletin for the History of Chemistry, 33(1):12–16, ???? 2008. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v33-1/v33-1%20p12-16.pdf.

Martin:2008:ANP

[383] Dean F. Martin, Barbara B. Martin, and Robert Alldredge. Arsenic, nitrate, and perchlorate in water: Dangers, distribution, and removal. Bulletin for the History of Chemistry, 33(1):17–24, ???? 2008. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v33-1/v33-1%20p17-24.pdf.

Martin:2008:JPP

[384] Dean F. Martin and Barbara B. Martin. The joint papers of Paul Karrer and Alfred Werner. Bulletin for the History of Chemistry, 33(1): 25–27, ???? 2008. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v33-
1/v33-1%20p25-27.pdf.

Rasmussen:2008:ACG

[385] Seth C. Rasmussen. Advances in 13th Century glass manufacturing and their effect on chemical progress. Bulletin for the History of Chemistry, 33(1):28–34, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-1/v33-1%20p28-34.pdf.

May:2008:LKC

[386] Leopold May. The lesser known chemist-composers, past and present. Bulletin for the History of Chemistry, 33(1):35–43, ???? 2008. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v33-1/v33-1%20p35-43.pdf.

Anonymous:2008:BRa

[387] Anonymous. Book reviews. Bulletin for the History of Chemistry, 33(1):45–58, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-1/v33-1%20p45-58.pdf.

Anonymous:2008:IAa

[388] Anonymous. Instructions to authors. Bulletin for the History of Chemistry, 33(1):59, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-1/v33-1%20p59.pdf.

Kraft:2008:DHP

[389] Alexander Kraft. On the discovery and history of Prussian blue. Bulletin for the History of Chemistry, 33(2):61–67, ???? 2008. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v33-2/v33-2%20p61-67.pdf.

Rubin:2008:HOV

[390] Mordecai B. Rubin. The history of ozone. VI. Ozone on silica gel (“dry ozone”). Bulletin for the History of Chemistry, 33(2):68–75, ???? 2008. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v33-
2/v33-2%20p68-75.pdf.

Plater:2008:CEC

[391] M. John Plater. The crucial early contributions of F. R. Japp to a general synthesis of imidazole derivatives. Bulletin for the History of Chemistry, 33(2):76–81, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-2/v33-2%20p76-81.pdf.

Kragh:2008:DAI

[392] Helge Kragh. From Disulfiram to Anabuse: the invention of a drug. Bulletin for the History of Chemistry, 33(2):82–88, ???? 2008. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v33-2/v33-2%20p82-88.pdf.

Kosteka:2008:ADS

[393] Keith Kosteka. Americium — from discovery to the smoke detector and beyond. Bulletin for the History of Chemistry, 33(2):89–93, ???? 2008. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v33-
2/v33-2%20p89-93.pdf.

Daemmrich:2008:GRC

[394] Arthur Daemmrich and Leah Shaper. The Gordon Research Conferences as scientific infrastructure. Bulletin for the History of Chemistry, 33(2):94–102, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-2/v33-2%20p94-102.pdf.

Cote:2008:HCR

[395] Gregory L. Cote and Victoria L. Finkenstadt. A history of carbohydrate research at the USDA Laboratory in Peoria, Illinois. Bulletin for the History of Chemistry, 33(2):103–111, ???? 2008. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v33-2/v33-2%20p103-111.pdf.

Anonymous:2008:BRb

[396] Anonymous. Book reviews. Bulletin for the History of Chemistry, 33(2):112–124, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-2/v33-2%20p112-124.pdf.

Anonymous:2008:CNE

[397] Anonymous. Call for nominations for the 2009 Edelstein Award. Bulletin for the History of Chemistry, 33(2):125, ???? 2008. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v33-2/v33-2%20p125.pdf.

Anonymous:2008:IAb

[398] Anonymous. Instructions to authors. Bulletin for the History of Chemistry, 33(2):126, ???? 2008. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v33-2/v33-2%20p126.pdf.

Rowlinson:2009:EAP

[399] J. S. Rowlinson. 2008 Edelstein Award paper: the border between physics and chemistry. Bulletin for the History of Chemistry, 34(1):1–10, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-1/v34-1%20p1-10.pdf.

Brock:2009:JRP

[400] William H. Brock. J. R. Partington (1886–1965): Physical chemistry in deed and word. Bulletin for the History of Chemistry, 34(1):11– 20, ???? 2009. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v34-
1/v34-1%20p11-20.pdf.

Wilson:2009:HMA

[401] Anne M. Wilson. Harry S. Mosher and Arthur C. Cope: Early organic chemists who mentored women. Bulletin for the History of Chemistry, 34(1):21–29, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-1/v34-1%20p21-29.pdf.

Habashi:2009:GHH

[402] Fathi Habashi. Gmelin and his Handbuch. Bulletin for the History of Chemistry, 34(1):30–31, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-1/v34-1%20p30-31.pdf.

Sourkes:2009:DEH

[403] Theodore L. Sourkes. The discovery and early history of carotene. Bulletin for the History of Chemistry, 34(1):32–38, ???? 2009. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v34-1/v34-1%20p32-38.pdf.

Rubin:2009:HOV

[404] Mordecai B. Rubin. The history of ozone. VII. The mythical spawn of ozone: Antozone, oxozone, and ozohydrogen. Bulletin for the History of Chemistry, 34(1):39–49, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-1/v34-1%20p39-49.pdf.

Orna:2009:WCN

[405] Mary Virginia Orna. Women chemists in the National Inventors’ Hall of Fame: Their remarkable lives and their award-winning research. Bulletin for the History of Chemistry, 34(1):50–60, ???? 2009. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v34-1/v34-1%20p50-60.pdf.

Anonymous:2009:BRa

[406] Anonymous. Book reviews. Bulletin for the History of Chemistry, 34(1):62–72, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-1/v34-1%20p62-72.pdf.

Jensen:2009:IET

[407] William B. Jensen. Introduction to the English translation of Theory of Dissociation [by August Horstmann]. Bulletin for the History of Chemistry, 34(2):73–75, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-2/v34-2%20p73-75.pdf.

Horstmann:2009:PDT

[408] A. Horstmann. Primary documents — “The theory of dissociation”. Bulletin for the History of Chemistry, 34(2):76–82, ???? 2009. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v34-2/v34-2%20p76-82.pdf.

Jensen:2009:AHO

[409] William B. Jensen. August Horstmann and the origins of chemical thermodynamics. Bulletin for the History of Chemistry, 34(2):83–91, ???? 2009. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v34-
2/v34-2%20p83-91.pdf.

Travis:2009:WBT

[410] Anthony S. Travis. A woman in biochemistry and toxicology: the Polish–British refugee Regina Schoental. Bulletin for the History of Chemistry, 34(2):92–104, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-2/v34-2%20p92-104.pdf.

Karpenko:2009:SNE

[411] Vladimir Karpenko. Some notes on the early history of nitric acid: 1300–1700. Bulletin for the History of Chemistry, 34(2):105–116, ???? 2009. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v34-
2/v34-2%20p105-116.pdf.

Rayner-Canham:2009:BWC

[412] Marelene Rayner-Canham and Geoff Rayner-Canham. British women and chemistry from the 16th to the mid-19th century. Bulletin for the History of Chemistry, 34(2):117–123, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-2/v34-2%20p117-123.pdf.

Page:2009:PCC

[413] Frederick G. Page. An 1815 perspective of chlorine as a chemical agent used in bleaching: a section from James Rennie’s essay on bleaching. Bulletin for the History of Chemistry, 34(2):124–133, ???? 2009. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v34-2/v34-2%20p124-133.pdf.

Kraft:2009:TLC

[414] Alexander Kraft. On two letters from Caspar Neumann to John Woodward: Revealing the secret method for the preparation of Prussian blue. Bulletin for the History of Chemistry, 34(2):134–140, ???? 2009. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v34-2/v34-2%20p134-140.pdf.

Benfey:2009:BPS

[415] Theodor Benfey. The biography of a periodic spiral. Bulletin for the History of Chemistry, 34(2):141–145, ???? 2009. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v34-2/v34-2%20p141-145.pdf.

Anonymous:2009:BRb

[416] Anonymous. Book reviews. Bulletin for the History of Chemistry, 34(2):146–159, ???? 2009. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v34-2/v34-2%20p146-159.pdf.

Levere:2010:EAS

[417] Trevor Harvey Levere. 2009 Edelstein Address: Sons of genius: Chemical manipulation and its shifting norms from Joseph Black to Michael Faraday. Bulletin for the History of Chemistry, 35(1):1–6, ???? 2010. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/v35-1/v35-1%20p1-6.pdf.

Raffa:2010:AHD

[418] R. B. Raffa and R. J. Tallarida. ’Afinity’: Historical development in chemistry and pharmacology. Bulletin for the History of Chemistry, 35(1):7–16, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-1/v35-1%20p7-16.pdf.

Halldahl:2010:RPM

[419] Lars Hälldahl. Revisiting Pharmacy Morianen: Revealing first traces of elemental silicon in a laboratory environment. Bulletin for the History of Chemistry, 35(1):17–23, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-1/v35-1%20p17-23.pdf.

Zumbulyadis:2010:BEN

[420] Independent Scholar Nicholas Zumbulyadis. Böttger’s eureka!: New insights into the European reinvention of porcelain. Bulletin for the History of Chemistry, 35(1):24–32, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-1/v35-1%20p24-32.pdf.

Marshall:2010:RVO

[421] James L. Marshall and Virginia R. Marshall. Reinvestigating vestium, one of the spurious platinum metals. Bulletin for the History of Chemistry, 35(1):33–39, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-1/v35-1%20p33-39.pdf.

Plater:2010:TCS

[422] M. John Plater. Three crucial scientific observations from mistaken hypotheses. Bulletin for the History of Chemistry, 35(1):40– 45, ???? 2010. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v35-
1/v35-1%20p40-45.pdf.

Dambrowitz:2010:HEM

[423] K. A. Dambrowitz and S. M. Kuznicki. Henry Eyring: a model life. Bulletin for the History of Chemistry, 35(1):46–52, ???? 2010. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v35-1/v35-1%20p46-52.pdf.

Anonymous:2010:BRa

[424] Anonymous. Book reviews. Bulletin for the History of Chemistry, 35(1):56–61, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-1/v35-1%20p56-61.pdf.

Anonymous:2010:RKF

[425] Anonymous. Recollections — Kasimir Fajans. Bulletin for the History of Chemistry, 35(1):62–63, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-1/v35-1%20p62-63.pdf.

Anonymous:2010:NEB

[426] Anonymous. New Editor for the Bulletin. Bulletin for the History of Chemistry, 35(2):65, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-2/v35-2%20p65.pdf.

Bohning:2010:HHI

[427] James J. Bohning. History of hist. II. On probation. Bulletin for the History of Chemistry, 35(2):66–80, ???? 2010. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v35-2/v35-2%20p66-80.pdf.

Henderson:2010:HEQ

[428] Douglas Henderson. Henry Eyring: Quantum chemistry, statistical mechanics, theory of liquids, and significant structure theory. Bulletin for the History of Chemistry, 35(2):81–85, ???? 2010. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v35-2/v35-2%20p81-85.pdf.

Thornton:2010:FEI

[429] Brett F. Thornton and Shawn C. Burdette. Finding eka-iodine: Discovery priority in modern times. Bulletin for the History of Chemistry, 35(2):86–96, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-2/v35-2%20p86-96.pdf.

Kragh:2010:ACR

[430] Helge Kragh. Auroral chemistry: the riddle of the green line. Bulletin for the History of Chemistry, 35(2):97–104, ???? 2010. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v35-2/v35-2%20p97-104.pdf.

Rubin:2010:DML

[431] Mordecai B. Rubin. The development of the mercury lamp. Bulletin for the History of Chemistry, 35(2):105–110, ???? 2010. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v35-2/v35-2%20p105-110.pdf.

Habashi:2010:CBH

[432] Fathi Habashi. Carl Bosch and his museum. Bulletin for the History of Chemistry, 35(2):111–114, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-2/v35-2%20p111-114.pdf.

Lewis:2010:FRM

[433] David E. Lewis. Feuding rule makers: Aleksandr Mikhailovich Zaitsev (1841–1910) and Vladimir Vasil’evich Markovnikov (1838–1904). Bulletin for the History of Chemistry, 35(2):115–124, ???? 2010. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v35-2/v35-2%20p115-124.pdf.

Anonymous:2010:BRb

[434] Anonymous. Book reviews. Bulletin for the History of Chemistry, 35(2):125–133, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-2/v35-2%20p125-133.pdf.

Jensen:2010:LAI

[435] William B. Jensen. Lost artifacts? The Irving Langmuir film. Bulletin for the History of Chemistry, 35(2):134, ???? 2010. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v35-2/v35-2%20p134.pdf. See response [436].

Jensen:2010:RPC

[436] William B. Jensen. Response to previous column: Lost artifacts? Anna Lea Painting. Bulletin for the History of Chemistry, 35(2):134, ???? 2010. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.

edu/~mainzv/HIST/bulletin_open_access/v35-2/v35-2%20p134.pdf.

See [435].

Anonymous:2010:AI

[437] Anonymous. Author index, 1–35. Bulletin for the History of Chemistry, 35(2):135–136, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-2/v35-2%20p135-136.pdf.

Anonymous:2010:SI

[438] Anonymous. Subject index, 1–35. Bulletin for the History of Chemistry, 35(2):137–168, ???? 2010. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v35-2/v35-2%20p137-168.pdf.

Anonymous:2011:CL

[439] Anonymous. Chairs’ letter. Bulletin for the History of Chemistry, 36 (1):1, ???? 2011. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v36-
1/v36-1%20p1.pdf.

Anonymous:2011:ELa

[440] Anonymous. Editor’s letter. Bulletin for the History of Chemistry, 36 (1):2, ???? 2011. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v36-
1/v36-1%20p2.pdf.

Kraft:2011:NCB

[441] Alexander Kraft. “Notitia Coerulei Berolinensis Nuper Inventi” on the 300th anniversary of the first publication on Prussian blue. Bulletin for the History of Chemistry, 36(1):3–9, ???? 2011. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v36-1/v36-1%20p3-9.pdf.

Jensen:2011:PCB

[442] William B. Jensen. Physical chemistry before Ostwald: the textbooks of Josiah Parsons Cooke. Bulletin for the History of Chemistry, 36(1):10–21, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p10-21.pdf.

Saltzman:2011:BSJ

[443] Martin D. Saltzman. Benjamin Silliman, Jr.’s 1874 papers: American contributions to chemistry. Bulletin for the History of Chemistry, 36(1):22–34, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p22-34.pdf.

Rayner-Canham:2011:RFD

[444] Marelene Rayner-Canham, Geoff Rayner-Canham, and Newfoundland Corner Brook. The rise and fall of domestic chemistry in higher education in England during the Early 20th Century. Bulletin for the History of Chemistry, 36(1):35–42, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p35-42.pdf.

Jensen:2011:DHS

[445] William B. Jensen. Denison–Hackh structure symbols: a forgotten episode in the teaching of organic chemistry. Bulletin for the History of Chemistry, 36(1):43–50, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p43-50.pdf.

Neelameggham:2011:LVH

[446] Neale R. Neelameggham. Letter: Vedic Hinduism and the four elements. Bulletin for the History of Chemistry, 36(1):51, ???? 2011. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v36-1/v36-1%20p51.pdf.

Anonymous:2011:BRJa

[447] Anonymous. Book review: John Scarborough, 2009. Pharmacy and Drug Lore in Antiquity: Greece, Rome, Byzantium. Bulletin for the History of Chemistry, 36(1):52–53, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p52-53%20br1.pdf.

Kein:2011:MEE

[448] Anonymous. Book review: Ursula Kein and E. C. Spary, Eds., 2010. Materials and Expertise in Early Modern Europe. Bulletin for the History of Chemistry, 36(1):54–55, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p54-55%20br2.pdf.

Anonymous:2011:BRJb

[449] Anonymous. Book review: John G. McEvoy, 2010. The Historiography of the Chemical Revolution: Patterns of Interpretation in the History of Science. Bulletin for the History of Chemistry, 36(1):56, ???? 2011. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v36-1/v36-1%20p56%20br3.pdf.

Anonymous:2011:BRDa

[450] Anonymous. Book review: David E. Fisher, 2010. Much Ado about (Practically) Nothing: a History of the Noble Gases. Bulletin for the History of Chemistry, 36(1):57, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-1/v36-1%20p57%20br4.pdf.

Anonymous:2011:BRDb

[451] Anonymous. Book review: Deborah Blum, 2010. The Poisoner’s Handbook: Murder and the Birth of Forensic Medicine in Jazz Age New York. Bulletin for the History of Chemistry, 36(1):58–59, ???? 2011. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v36-1/v36-1%20p58-59%20br5.
pdf.

Anonymous:2011:ELb

[452] Anonymous. Editor’s letter. Bulletin for the History of Chemistry, 36 (2):60, ???? 2011. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v36-
2/v36-2%20p60.pdf.

Mauskopf:2011:DHC

[453] Seymour H. Mauskopf. Do historians or chemists write better: History of chemistry? Bulletin for the History of Chemistry, 36(2):61– 67, ???? 2011. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v36-
2/v36-2%20p61-67.pdf.

Rayner-Canham:2011:CEA

[454] Marelene Rayner-Canham, Geoff Rayner-Canham, and Newfoundland Corner Brook. Chemistry in English academic girls’ schools, 1880– 1930. Bulletin for the History of Chemistry, 36(2):68–74, ???? 2011. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v36-2/v36-2%20p68-74.pdf.

Laszlo:2011:HAA

[455] Pierre Laszlo. How an Anglo-American methodology took root in France. Bulletin for the History of Chemistry, 36(2):75–81, ???? 2011. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v36-2/v36-2%20p75-81.pdf.

Collins:2011:COD

[456] Sibrina N. Collins. Celebrating our diversity. The education of some pioneering African American chemists in Ohio. Bulletin for the History of Chemistry, 36(2):82–84, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-2/v36-2%20p82-84.pdf.

Jensen:2011:IET

[457] William B. Jensen and Julia Kuhlmann. Introduction to the English translation of A Contribution to Chemical Statics by Leopold Pfaundler: a forgotten classic of chemical kinetics. Bulletin for the History of Chemistry, 36(2):85–86, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-2/v36-2%20p85-86.pdf.

Pfaundler:2011:PDC

[458] Leopold Pfaundler. Primary documents “A Contribution to Chemical Statics”. Bulletin for the History of Chemistry, 36(2):87–98, ???? 2011. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v36-
2/v36-2%20p87-98.pdf.

Anonymous:2011:BRAa

[459] Anonymous. Book review: Adam Ganz, 15 June 2010. Nuclear Reactions. Bulletin for the History of Chemistry, 36(2):99–101, ???? 2011. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v36-2/v36-2%20p99-101.pdf.

Anonymous:2011:BRL

[460] Anonymous. Book review: Lauren Redniss, 2010. Radioactive: Marie & Pierre Curie, a Tale of Love and Fallout. Bulletin for the History of Chemistry, 36(2):101–102, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-2/v36-2%20p101-102.pdf.

Anonymous:2011:BRAb

[461] Anonymous. Book review: Anne-Marie Weidler Kubanek, 2010. Nothing Less Than an Adventure: Ellen Gleditsch and Her Life in Science. Bulletin for the History of Chemistry, 36(2):102–103, ???? 2011. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v36-2/v36-2%20p102-103.pdf.

Anonymous:2011:BRJc

[462] Anonymous. Book review: John E. Lesch, 2007. The First Miracle Drugs: How the Sulfa Drugs Transformed Medicine. Bulletin for the History of Chemistry, 36(2):103–104, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-2/v36-2%20p103-104.pdf.

Anonymous:2011:BRMa

[463] Anonymous. Book review: Michael Hunter; 2009. Boyle: Between God and Science. Bulletin for the History of Chemistry, 36(2):105– 106, ???? 2011. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v36-
2/v36-2%20p105-106.pdf.

Anonymous:2011:BRC

[464] Anonymous. Book review: Carmen J. Giunta, Ed., 2010. Atoms in Chemistry: from Dalton’s Predecessors to Complex Atoms and Beyond. Bulletin for the History of Chemistry, 36(2):106–107, ???? 2011. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v36-2/v36-2%20p106-107.pdf.

Anonymous:2011:BRMb

[465] Anonymous. Book review: Mark Griep and Marjorie Mikasen, 2009. Reaction! Chemistry in the Movies. Bulletin for the History of Chemistry, 36(2):108, ???? 2011. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v36-2/v36-2%20p108.pdf.

Anonymous:2011:BRH

[466] Anonymous. Book review: Hugh Aldersey-Williams, 2011. Periodic Tales: a Cultural History of the Elements from Arsenic to Zinc. Bulletin for the History of Chemistry, 36(2):109–110, ???? 2011. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v36-2/v36-2%20p109-110.pdf.

Valderrama:2012:ASM

[467] Ignacio Miguel Pascual Valderrama and Joaquín Pérez-Pariente. Alchemy at the service of mining technology in Seventeenth-Century Europe, according to the works of Martine De Bertereau and Jean Du Chastelet. Bulletin for the History of Chemistry, 37(1):1–13, ???? 2012. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v37-1/v37-1%20p1-13.pdf.

Jones:2012:CME

[468] Paul R. Jones. Contrasting mentors for English-speaking chemistry students in Germany in the Nineteenth Century: Liebig, Wöhler, and Bunsen. Bulletin for the History of Chemistry, 37(1):14–23, ???? 2012. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v37-1/v37-1%20p14-23.pdf.

Raos:2012:PSP

[469] Nenad Raos and Croatia Zagreb. Pan-Slavism and the periodic system of the elements. Bulletin for the History of Chemistry, 37(1):24– 28, ???? 2012. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v37-
1/v37-1%20p24-28.pdf.

Jensen:2012:LPO

[470] William B. Jensen and Julia Kuhlmann. Leopold Pfaundler and the origins of the kinetic theory of chemical reactions. Bulletin for the History of Chemistry, 37(1):29–41, ???? 2012. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v37-1/v37-1%20p29-41.pdf.

Kragh:2012:LVA

[471] Helge Kragh. Lars Vegard, atomic structure, and the periodic system. Bulletin for the History of Chemistry, 37(1):42–49, ???? 2012. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v37-1/v37-1%20p42-49.pdf.

Anonymous:2012:BRMa

[472] Anonymous. Book review: Mary Jo Nye. 2011. Michael Polanyi and His Generation: Origins of the Social Construction of Science. Bulletin for the History of Chemistry, 37(1):50–52, ???? 2012. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v37-1/v37-1%20p50-52.pdf.

Anonymous:2012:BRWa

[473] Anonymous. Book review: Walter W. Woodward, 2010. Prospero’s America. John Winthrop, Jr., Alchemy and the Creation of New England Culture, 1606–1676. Bulletin for the History of Chemistry, 37(1):52–53, ???? 2012. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v37-1/v37-1%20p52-53.pdf.

Anonymous:2012:BRMb

[474] Anonymous. Book review: Michael Faraday, Frank A. J. L. James, Ed. 2011. The Chemical History of a Candle, sesquicentenary edition. Bulletin for the History of Chemistry, 37(1):54–55, ???? 2012. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v37-1/v37-1%20p54-55.pdf.

Apotheker:2012:EWC

[475] Jan Apotheker and Livia Simon Sarkadi. European women in chemistry. Bulletin for the History of Chemistry, 37(1):55–56, ???? 2012. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v37-1/v37-1%20p55-56.pdf.

Anonymous:2012:BRMc

[476] Anonymous. Book review: Marjorie C. Malley, 2011. Radioactivity: a History of a Mysterious Science. Bulletin for the History of Chemistry, 37(1):56–57, ???? 2012. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v37-1/v37-1%20p56-57.pdf.

Anonymous:2012:BRWb

[477] Anonymous. Book review: William H. Brock, 2011. The Case of the Poisonous Socks: Tales from Chemistry. Bulletin for the History of Chemistry, 37(1):58–59, ???? 2012. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v37-1/v37-1%20p58-59.pdf.

Guerrero:2012:CTHa

[478] Saúl Guerrero. Chemistry as a tool for historical research: Identifying paths of historical mercury pollution in the Hispanic New World. Bulletin for the History of Chemistry, 37(2):61–70, ???? 2012. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v37-2/v37-2%20p61-70.pdf.

Guerrero:2012:CTHb

[479] Saúl Guerrero. Chemistry as a tool for historical research: Estimating the contraband of silver from Potosí and Oruro, 1576–1650. Bulletin for the History of Chemistry, 37(2):72–80, ???? 2012. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v37-2/v37-2%20p72-80.pdf.

Saltzman:2012:JWB

[480] Martin D. Saltzman. John William Baker and the origin of the Baker– Nathan effect. Bulletin for the History of Chemistry, 37(2):82–90, ???? 2012. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v37-
2/v37-2%20p82-90.pdf.

Witty:2012:PTC

[481] Michael Witty. The process for 2,3,5-triphenyltetrazolium chloride synthesis, an intellectual property seized immediately after World War II. Bulletin for the History of Chemistry, 37(2):91–95, ???? 2012. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v37-2/v37-2%20p91-95.pdf.

Martin:2012:KLC

[482] Dean F. Martin and Karina Bidani. The Kamlet Laboratories Collection at the University of South Florida. Bulletin for the History of Chemistry, 37(2):96–101, ???? 2012. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v37-2/v37-2%20p96-101.pdf.

Anonymous:2012:BRT

[483] Anonymous. Book review: E. Thomas Strom and Seth C. Rasmussen, Eds., 2011. 100+ Years of Plastics: Leo Baekeland and Beyond. Bulletin for the History of Chemistry, 37(2):102–103, ???? 2012. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v37-2/v37-2%20p102-103.pdf.

Anonymous:2012:BRK

[484] Anonymous. Book review: Kostas Gavroglu and Ana Simões, 2012. Neither Physics nor Chemistry: a History of Quantum Chemistry. Bulletin for the History of Chemistry, 37(2):103–105, ???? 2012. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v37-2/v37-2%20p103-105.pdf.

Anonymous:2012:BRJ

[485] Anonymous. Book review: Jeannette Brown, 2012. African American Women Chemists. Bulletin for the History of Chemistry, 37(2):106– 107, ???? 2012. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v37-
2/v37-2%20p106-107.pdf.

Shakhashiri:2013:YMA

[486] Bassam Z. Shakhashiri. 150 years of the Morrill Act: the promise and potential of the land-grant university. Bulletin for the History of Chemistry, 38(2):83–84, ???? 2013. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v38-2/v38-2%20p83-84.pdf.

Weininger:2013:I

[487] Stephen J. Weininger. Introduction. Bulletin for the History of Chemistry, 38(2):84–85, ???? 2013. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v38-2/v38-2%20p84-85.pdf.

Yarmey:2013:CVC

[488] Kristen A. Yarmey. Communicating the value of chemistry: Evan Pugh, Penn State, and public confidence at the time of the land grant. Bulletin for the History of Chemistry, 38(2):86–96, ???? 2013. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v38-2/v38-2%20p86-96.pdf.

Weininger:2013:CIC

[489] Stephen J. Weininger. Chemistry for the “industrial classes”: Laboratory instruction, mass education and women’s experience in midwestern land-grant colleges, 1870–1914 (1). Bulletin for the History of Chemistry, 38(2):97–108, ???? 2013. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v38-2/v38-2%20p97-108.pdf.

Marcus:2013:CUM

[490] Alan I. Marcus. Chemistry under the Morrill Act: Agency through service. Bulletin for the History of Chemistry, 38(2):109–113, ???? 2013. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v38-2/v38-2%20p109-113.pdf.

Finlay:2013:CLG

[491] Mark R. Finlay. Chemurgy and the land grant colleges: Bridging agriculture, industry and chemistry in the 1930s and beyond. Bulletin for the History of Chemistry, 38(2):114–122, ???? 2013. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v38-2/v38-2%20p114-122.pdf.

Seidel:2013:CSC

[492] Robert W. Seidel. Catalyst or synthesis? Chemical engineering in the land-grant college. Bulletin for the History of Chemistry, 38(2):123– 131, ???? 2013. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v38-
2/v38-2%20p123-131.pdf.

Bix:2013:CCC

[493] Amy Bix. Chemistry of cooking, chemistry in war: Women in nineteenth and twentieth-century land-grant science and engineering. Bulletin for the History of Chemistry, 38(2):132–139, ???? 2013. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v38-2/v38-2%20p132-139.pdf.

Page:2013:BPS

[494] Frederick G. Page. Baking powder and self-rising flour in Nineteenth-Century Britain: the carbon dioxide aerations of Henry Jones and Alfred Bird. Bulletin for the History of Chemistry, 38(2):140–??, ???? 2013. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2013:BRU

[495] Anonymous. Book review: Ulf Lagerkvist. 2012. The Periodic Table and a Missed Nobel Prize. Bulletin for the History of Chemistry, 38(2):155–??, ???? 2013. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2013:BRE

[496] Anonymous. Book review: Eric Scerri, 2013. A Tale of 7 Elements. Bulletin for the History of Chemistry, 38(2):157–??, ???? 2013. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2013:IA

[497] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 38(2):159–??, ???? 2013. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v38-2/v38-2%20p159.pdf.

Lewis:2014:CCB

[498] David E. Lewis. Citation for chemical breakthrough award: Mendeleev’s periodic system of the elements. Bulletin for the History of Chemistry, 39 (1):1–??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://

www.scs.illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-
1/v39-1%20p1-6.pdf.

Zumbulyadis:2014:DEU

[499] Nicholas Zumbulyadis. Decorating with explosives: the use of Aurum Fulminans as a purple pigment supplemental information. Bulletin for the History of Chemistry, 39(1):7–??, ???? 2014. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/%
7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%20p7-17.pdf.

Kraft:2014:HPB

[500] Alexander Kraft. On the history of Prussian blue: Thomas Everitt (1805– 1845) and Everitt’s salt. Bulletin for the History of Chemistry, 39(1):18– ??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://www.

scs.illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-1/v39-

1%20p18-25.pdf. See addendum [562].

Herold:2014:TNC

[501] Bernardo Jerosch Herold and Wolfram Bayer. A transnational network of chemical knowledge: the Preparadores at the Lisbon Polytechnic School in the 1860s and 1870s. Bulletin for the History of Chemistry, 39(1):26–??, ???? 2014. CODEN BHCHET. ISSN

1053-4385. URL http://www.scs.illinois.edu/%7Emainzv/HIST/
bulletin_open_access/v39-1/v39-1%20p26-42.pdf.

Suntsov:2014:CBP

[502] Vladislav Suntsov and David E. Lewis. A century of base-promoted decomposition of hydrazones: the early career of Nikolai Matveevich Kizhner (1867–1935). Bulletin for the History of Chemistry, 39(1):43–??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%
20p43-52.pdf.

Lavine:2014:TFR

[503] Matthew Lavine. The two faces of radium in early American nuclear culture. Bulletin for the History of Chemistry, 39(1):53–??, ???? 2014. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
%7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%20p53-63.pdf.

Rasmussen:2014:PCP

[504] Seth C. Rasmussen. The path to conductive polyacetylene. Bulletin for the History of Chemistry, 39(1):64–??, ???? 2014. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/%
7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%20p64-72.pdf.

Gal:2014:DUF

[505] Joseph Gal and Jeffrey I. Seeman. In defense of the use of the French language in scientific communication, 1965–1985: National and international deliberations and an ingeniously clever takeoff on the theme by R. B. Woodward supplemental information. Bulletin for the History of Chemistry, 39(1):73–??, ???? 2014. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/%7Emainzv/
HIST/bulletin_open_access/v39-1/v39-1%20p73-94.pdf.

Anonymous:2014:BRa

[506] Anonymous. Book reviews. Bulletin for the History of Chemistry, 39(1): ??, ???? 2014. CODEN BHCHET. ISSN 1053-4385.

Principe:2014:BRS

[507] Lawrence M. Principe. Book review: The Secrets of Alchemy. Bulletin for the History of Chemistry, 39(1):95–??, ???? 2014. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/%
7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%20p95-97.pdf.

deMenten:2014:BRD

[508] P. de Menten. Book review: Dictionnaire de chimie: Une approche étymologique et historique. (French) [Dictionary of chemistry: An etymological and historical approach]. Bulletin for the History of Chemistry, 39(1):97–??, ???? 2014. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/%7Emainzv/
HIST/bulletin_open_access/v39-1/v39-1%20p97.pdf.

Vijvers:2014:BRA

[509] William Vijvers. Book review: Alexander Borodin: Composer, Scientist, Educator. a Biography. Bulletin for the History of Chemistry, 39(1):98–??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%
20p98-99.pdf.

Green:2014:BRO

[510] Mark M. Green. Book review: Organic Chemistry Principles in Context: a Story Telling Historical Approach Comment and Response. Bulletin for the History of Chemistry, 39(1):99–??, ???? 2014. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/%
7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%20p99-100.pdf.

Davies:2014:BRU

[511] Alwyn Davies and Peter Grant. Book review: UCL Chemistry Department 1828–1974. Bulletin for the History of Chemistry, 39(1):101–??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%
20p101.pdf.

Strom:2014:BRP

[512] E. Thomas Strom and Angela K. Wilson. Book review: Pioneers of Quantum Chemistry. Bulletin for the History of Chemistry, 39(1):102–??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%
20p102-103.pdf.

Stocker:2014:BRF

[513] Jack Stocker and Natalie Foster. Book review: A Festival of Chemistry Entertainments. Bulletin for the History of Chemistry, 39(1):103–??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http://www.scs.

illinois.edu/%7Emainzv/HIST/bulletin_open_access/v39-1/v39-1%
20p103-104.pdf.

Anonymous:2014:IAa

[514] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 39(1):105–??, ???? 2014. CODEN BHCHET. ISSN

1053-4385. URL http://www.scs.illinois.edu/%7Emainzv/HIST/
bulletin_open_access/v39-1/v39-1%20p105.pdf.

Jensen:2014:KVT

[515] William B. Jensen. Kinetic versus thermodynamic control: Some historical landmarks. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p107-121.pdf.

Moore:2014:FLT

[516] Carl E. Moore, Alfred von Smolinski, Albert Claus, Daniel J. Graham, and Bruno Jaselskis. On the First Law of Thermodynamics and the contribution of Julius Robert Mayer: New translation and consideration of a rejected manuscript. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p122-130.pdf.

Rayner-Canham:2014:EPC

[517] Marelene F. Rayner-Canham and Geoffrey W. Rayner-Canham. Early practical chemistry at British private girls’ schools. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p131-137.pdf.

Rae:2014:WGM

[518] Ian D. Rae. William Gilbert Mixter (1846–1936): a Yale chemist who deserves to be remembered. Bulletin for the History of Chemistry, 39 (2):??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p138-144.pdf.

Jones:2014:ONE

[519] Paul R. Jones. Ouroboros, a New England chemists’ club. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v39-2/v39-2/v39-2%20p145-149.pdf.

Kirk:2014:HCN

[520] Kenneth L. Kirk and Kenneth A. Jacobson. History of chemistry in the National Institute of Diabetes and Digestive Kidney Diseases (NIDDK): Supplemental information. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p150-165.pdf.

Anonymous:2014:BRb

[521] Anonymous. Book reviews. Bulletin for the History of Chemistry, 39(2): ??, ???? 2014. CODEN BHCHET. ISSN 1053-4385.

Patterson:2014:CCC

[522] Gary D. Patterson and Seth C. Rasmussen. Characters in chemistry: a celebration of the humanity of chemistry. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p166-168.pdf.

Fleming:2014:TAB

[523] James Rodger Fleming and Ann Johnson. Toxic airs: Body, place, planet in historical perspective. Bulletin for the History of Chemistry, 39(2):

??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p168-170.pdf.

Steen:2014:ASC

[524] Kathryn Steen. The American synthetic chemicals industry: War and politics, 1910–1930. Bulletin for the History of Chemistry, 39(2): ??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p170-172.pdf.

Labinger:2014:GHI

[525] Jay A. Labinger. Up from generality: How inorganic chemistry finally became a respectable field. Bulletin for the History of Chemistry, 39(2): ??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p173-174.pdf.

Approach:2014:CR

[526] Anonymous. Comment and response: Review of organic chemistry principles in context: A story telling historical approach. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p175.pdf.

Anonymous:2014:CPG

[527] Anonymous. Comment by Prof. Green. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p175-178.pdf.

Anonymous:2014:RPR

[528] Anonymous. Response by Prof. Ramberg. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p178-180.pdf.

Anonymous:2014:E

[529] Anonymous. Errata. Bulletin for the History of Chemistry, 39(2): ??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p181-182.pdf.

Anonymous:2014:IAb

[530] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p183.pdf.

Anonymous:2014:BES

[531] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 39(2):??, ???? 2014. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v39-2/v39-2/v39-2%20p184.pdf.

Anonymous:2014:HO

[532] Anonymous. Hist officers 2014. Bulletin for the History of Chemistry, 39 (2):??, ???? 2014. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v39-
2/v39-2/v39-2%20p184.pdf.

Panzarasa:2015:RPA

[533] Guido Panzarasa. Rediscovering pyrotartaric acid: a chemical interpretation of the volatile salt of tartar. Bulletin for the History of Chemistry, 40(1):1–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p1-8.pdf. See [570].

Lewis:2015:IET

[534] David E. Lewis. Introduction to an English translation, “On the Different Explanations of Certain Cases of Isomerism” by Aleksandr Butlerov. Bulletin for the History of Chemistry, 40(1):9–??, ???? 2015. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v40-1/v40-1%20p9-12.pdf.

Boutlerow:2015:PDD

[535] A. Boutlerow. Primary documents “On the Different Explanations of Certain Cases of Isomerism”. Bulletin for the History of Chemistry, 40(1):13–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p13-28.pdf. Translated by David E.

Lewis.

Labinger:2015:WIN

[536] Jay A. Labinger. Why isn’t noble gas chemistry 30 years older? The failed (?) 1933 experiment of Yost and Kaye. Bulletin for the History of Chemistry, 40(1):29–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p29-36.pdf.

Collins:2015:RPB

[537] Sibrina N. Collins. Robert Percy Barnes: from Harvard to Howard University. Bulletin for the History of Chemistry, 40(1):37–??, ???? 2015. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v40-1/v40-1%20p37-39.pdf.

Raos:2015:SPC

[538] Nenad Raos. Science and politics: a case study of the Croatian Chemical Journal. Bulletin for the History of Chemistry, 40(1):40– ??, ???? 2015. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
1/v40-1%20p40-44.pdf.

Rasmussen:2015:EHP

[539] Seth C. Rasmussen. Early history of polypyrrole: the first conducting organic polymer. Bulletin for the History of Chemistry, 40(1):45– ??, ???? 2015. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
1/v40-1%20p45-55.pdf.

Anonymous:2015:BRa

[540] Anonymous. Book reviews. Bulletin for the History of Chemistry, 40(1): ??, ???? 2015. CODEN BHCHET. ISSN 1053-4385.

Fontani:2015:LEP

[541] Mariagrazia Costa Marco Fontani and Mary Virginia Orna. The lost elements: the periodic table’s shadow side. Bulletin for the History of Chemistry, 40(1):56–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p56-57.pdf.

Hargittai:2015:GMR

[542] Magdolna Hargittai Balazs Hargittai and Istvan Hargittai. Great minds: Reflections of 111 top scientists. Bulletin for the History of Chemistry, 40(1):58–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p58-59.pdf.

Anonymous:2015:IAa

[543] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 40(1):60–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p60.pdf.

Anonymous:2015:BESa

[544] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 40(1):61–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p61.pdf.

Anonymous:2015:HOa

[545] Anonymous. Hist officers 2015. Bulletin for the History of Chemistry, 40(1):61–??, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-1/v40-1%20p61.pdf.

Anonymous:2015:IET

[546] Anonymous. Introduction to an English translation (abridged) of Kizhner’s pioneering papers on deoxygenation supplemental information. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v40-2/vol40-2%20p61-64.pdf.

Anonymous:2015:PDCa

[547] Anonymous. Primary documents. 25. The catalytic decomposition of alkylidenehydrazines as a method for the preparation of hydrocarbons supplemental information. Bulletin for the History of Chemistry, 40 (2):–, ???? 2015. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
2/vol40-2%20p64-68.pdf.

Anonymous:2015:PDCb

[548] Anonymous. Primary documents. 27. On the catalytic decomposition of alkylidenehydrazines. (Second part) supplemental information. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v40-2/vol40-2%20p69-73.pdf.

Anonymous:2015:CFC

[549] Anonymous. From cosmochemistry to fuel cells: Notes on Emil Baur, physical chemist. Bulletin for the History of Chemistry, 40(2):

–, ???? 2015. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
2/vol40-2%20p74-85.pdf.

Anonymous:2015:HSS

[550] Anonymous. The history and structure of stantienite. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/vol40-2%20p86-94.pdf.

Anonymous:2015:TCD

[551] Anonymous. Tools for chemists: the Desreux–Bischoff viscosimeter. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v40-2/vol40-2%20p95-102.pdf.

Anonymous:2015:PCP

[552] Anonymous. A pioneering course in physical organic chemistry: J. W. Baker’s 1942 third-year lectures to undergraduates. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/vol40-2%20p103-108.pdf.

Anonymous:2015:CES

[553] Anonymous. A compelling example of scientific integrity. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/vol40-2%20p109-113.pdf.

Anonymous:2015:RKH

[554] Anonymous. A reverie, Kekulé and his dream: an interview. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v40-2/vol40-2%20p114-119.pdf.

Anonymous:2015:BRb

[555] Anonymous. Book reviews. Bulletin for the History of Chemistry, 40(2): ??, ???? 2015. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2015:LMC

[556] Anonymous. The limits of matter: Chemistry, mining, and enlightenment. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v40-2/vol40-2%20p120-121.pdf.

Anonymous:2015:MMI

[557] Anonymous. Medical monopoly: Intellectual property rights and the origins of the modern pharmaceutical industry. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/vol40-2%20p122-123.pdf.

Anonymous:2015:SHT

[558] Anonymous. Science history: a traveler’s guide. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/vol40-2%20p123-125.pdf.

Anonymous:2015:IAb

[559] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 40(2):??, ???? 2015. CODEN BHCHET. ISSN 1053-4385.

URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_open_
access/v40-2/vol40-2%20p126%20Instructions%20for%20Authors.pdf.

Anonymous:2015:BESb

[560] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 40(2):–, ???? 2015. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
2/vol40-2%20p127%20Editorial%20Board%20HIST%20Officers.pdf.

Anonymous:2015:HOb

[561] Anonymous. Hist officers 2015. Bulletin for the History of Chemistry, 40 (2):–, ???? 2015. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
2/vol40-2%20p127%20Editorial%20Board%20HIST%20Officers.pdf.

Kraft:2016:ADH

[562] Alexander Kraft. Addendum to “on the discovery and history of Prussian blue”. Bulletin for the History of Chemistry, 41(1–2):1–2, ???? 2016. CO-

DEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.
edu/bulletin_open_access/vol41-1-2/vol41-1-2%20p1-2.pdf. See

Lewis:2016:KKD

[563] David E. Lewis. Klaus at Kazan: the discovery of ruthenium. Bulletin for the History of Chemistry, 41(1–2):3–11, ???? 2016. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/vol41-1-2/vol41-1-2%20p3-11.pdf.

Moreland:2016:EFS

[564] David W. Moreland and Paul R. Jones. Emil Fischer’s sample collection. Bulletin for the History of Chemistry, 41(1–2):12–18, ???? 2016. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/vol41-1-2/vol41-1-2%20p12-18.pdf.

Jensen:2016:CTI

[565] William B. Jensen and Peter J. T. Morris. From chemical theory to industrial chemistry: the eclectic career of Geoffrey Martin. Bulletin for the History of Chemistry, 41(1–2):19–37, ???? 2016. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/vol41-1-2/vol41-1-2%20p19-37.pdf.

Epstein:2016:LTL

[566] Jessica L. Epstein. The legacy of tetraethyl lead. Bulletin for the History of Chemistry, 41(1–2):38–43, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p38-43.pdf.

Scheiding:2016:MTM

[567] Tom Scheiding. More than meets the eye: Chemical foundation investments in the Journal of Chemical Education. Bulletin for the History of Chemistry, 41(1–2):44–55, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p44-55.pdf.

Kovac:2016:ECW

[568] Jeffrey Kovac. Ethics of chemical weapons research. Bulletin for the History of Chemistry, 41(1–2):56–63, ???? 2016. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/vol41-1-2/vol41-1-2%20p56-63.pdf.

Rasmussen:2016:OSM

[569] Seth C. Rasmussen. On the origin of “synthetic metals”: Herbert Mc-Coy, Alfred Ubbelohde, and the development of metals from nonmetallic elements. Bulletin for the History of Chemistry, 41(1–2):64–73, ????

2016. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/vol41-1-2/vol41-1-2%20p64-
73.pdf.

Anonymous:2016:CRR

[570] Anonymous. Comment and response: Rediscovering pyrotartaric acid. Bulletin for the History of Chemistry, 41(1–2):74, ???? 2016. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/vol41-1-2/vol41-1-2%20p74.pdf. See [533].

Anonymous:2016:CL

[571] Joseph B. Lambert. Book review: A Chemical Life, Joseph B. Lambert, De Rigueur Press, North Manchester, IN, 2014, 388 pp, ISBN 978-0-9916503-0-9, $15.00. Bulletin for the History of Chemistry, 41(1–2):75–76, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p75-76.pdf.

Anonymous:2016:MFH

[572] Peter J. T. Morris. Book review: The Matter Factory: a History of the Chemical Laboratory. Bulletin for the History of Chemistry, 41(1–2):77–79, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p77-79.pdf.

Anonymous:2016:IA

[573] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 41(1–2):80, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p80%20Instructions%20for%20Authors.
pdf.

Anonymous:2016:BES

[574] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 41(1–2):82, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p82.pdf.

Anonymous:2016:HO

[575] Anonymous. HIST officers 2016. Bulletin for the History of Chemistry, 41(1–2):82, ???? 2016. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/vol41-1-2/vol41-1-2%20p82.pdf.

Klein:2017:CCG

[576] Ursula Klein. Chemists for the common good. Bulletin for the History of Chemistry, 42(1):1–6, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/v42-1/vol42-1%20p1-6.pdf.

Leigh:2017:CCC

[577] G. J. Leigh. The changing content of conversations on chemistry as a snapshot of the: Development of chemical science. Bulletin for the History of Chemistry, 42(1):7–28, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/v42-1/vol42-1%20p7-28.pdf.

Page:2017:CDS

[578] Frederick G. Page. Carbon dioxide in self-rising flour and baking powder: a study in apparatus, Scheibler to Chittick. Bulletin for the History of Chemistry, 42(1):29–45, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/v42-1/vol42-1%20p29-45.pdf.

Suntsov:2017:ARN

[579] Vladislav Suntsov and David E. Lewis. After the revolution: Nikolai Matveevich Kizhner (1867–1935) in Soviet Moscow. Bulletin for the History of Chemistry, 42(1):45–56, ???? 2017. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v40-2/v42-1/vol42-1%20p46-56.pdf.

Raos:2017:CCO

[580] Nenad Raos. Carbide chemistry and Oparin’s theory on the origin of life. Bulletin for the History of Chemistry, 42(1):57–62, ???? 2017. CO-

DEN BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/
~mainzv/HIST/bulletin_open_access/v40-2/v42-1/vol42-1%20p57-
62.pdf.

Rasmussen:2017:CHC

[581] Seth C. Rasmussen. Cuprene: a historical curiosity along the path to polyacetylene. Bulletin for the History of Chemistry, 42(1):63–78, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
2/v42-1/vol42-1%20p63-78.pdf.

Seeman:2017:CMP

[582] Jeffrey I. Seeman. Commentary: Moving past a seminal generation in the history of chemistry: a moment to reflect. Bulletin for the History of Chemistry, 42(1):79–80, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/v42-1/vol42-1%20p79-80.pdf.

MacLeod:2017:BRC

[583] Roy MacLeod. Book review: The Chemist’s War, 1914–1918. Bulletin for the History of Chemistry, 42(1):81–82, ???? 2017. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v40-2/v42-1/vol42-1%20p81-82.pdf.

Anonymous:2017:IAa

[584] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 42(1):83–84, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/v42-1/vol42-1%20p83.pdf.

Anonymous:2017:BESa

[585] Anonymous. Bulletin Editorial Staff. Bulletin for the History of Chemistry, 42(1):84, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v40-2/v42-1/vol42-1%20p84.pdf.

Anonymous:2017:HOa

[586] Anonymous. HIST officers 2017. Bulletin for the History of Chemistry, 42(1):84, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v40-
2/v42-1/vol42-1%20p84.pdf.

Leigh:2017:IPH

[587] G. J. Leigh. The international publication history of Conversations on Chemistry: the correspondence of Jane and Alexander Marcet during its writing. Bulletin for the History of Chemistry, 42(2):85–93, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v42-
2/v42-2/v42-2%20p85-93.pdf.

Martin:2017:FWW

[588] Dean F. Martin and Barbara B. Martin. Five women who worked with Alfred Werner. Bulletin for the History of Chemistry, 42(2):94– 102, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v42-
2/v42-2/v42-2%20p94-102.pdf.

Anonymous:2017:IIB

[589] Anonymous. Isotopes: Identifying the breakthrough publication. Bulletin for the History of Chemistry, 42(2):103–111, ???? 2017. CODEN

BHCHET. ISSN 1053-4385. URL http://www.scs.illinois.edu/~
mainzv/HIST/bulletin_open_access/v42-2/v42-2/v42-2%20p103-111.
pdf.

Lewis:2017:YAF

[590] David E. Lewis. A. Ye. Arbuzov: Father of organophosphorus chemistry in Russia. Bulletin for the History of Chemistry, 42(2):112–125, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v42-
2/v42-2/v42-2%20p112-125.pdf.

Rayner-Canham:2017:WCL

[591] Marelene Rayner-Canham and Geoff Rayner-Canham. Women chemists of the London School of Medicine for Women, 1874–1947. Bulletin for the History of Chemistry, 42(2):126–132, ???? 2017. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v42-2/v42-2/v42-2%20p126-132.pdf.

Paolieri:2017:FME

[592] Matteo Paolieri. Ferdinand Münz: EDTA and 40 years of inventions. Bulletin for the History of Chemistry, 42(2):133–140, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v42-
2/v42-2/v42-2%20p133-140.pdf.

Kragh:2017:BRT

[593] Helge Kragh. Book review: A Tale of Seven Scientists and a New Philosophy of Science. Bulletin for the History of Chemistry, 42(2):141– 143, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:/

/www.scs.illinois.edu/~mainzv/HIST/bulletin_open_access/v42-
2/v42-2/v42-2%20p141-143.pdf. See response [617].

Giunta:2017:BRC

[594] Carmen J. Giunta. Book review: A Chemical Passion. Bulletin for the History of Chemistry, 42(2):144–145, ???? 2017. CODEN BHCHET.

ISSN 1053-4385. URL http://www.scs.illinois.edu/~mainzv/HIST/
bulletin_open_access/v42-2/v42-2/v42-2%20p144-145.pdf.

Anonymous:2017:IAb

[595] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 42(2):146, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v42-2/v42-2/v42-2%20p146.pdf.

Anonymous:2017:BESb

[596] Anonymous. Bulletin Editorial Staff. Bulletin for the History of Chemistry, 42(2):147, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v42-2/v42-2/v42-2%20p147.pdf.

Anonymous:2017:HOb

[597] Anonymous. HIST officers 2017. Bulletin for the History of Chemistry, 42(2):147, ???? 2017. CODEN BHCHET. ISSN 1053-

4385. URL http://www.scs.illinois.edu/~mainzv/HIST/bulletin_
open_access/v42-2/v42-2/v42-2%20p147.pdf.

Anonymous:2017:IA

[598] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 42(2):146, ???? 2017. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v42-2/v42-2%20p146.pdf.

Anonymous:2017:BES

[599] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 42(2):147, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v42-2/v42-2%
20p147.pdf.

Anonymous:2017:HOc

[600] Anonymous. HIST officers 2017. Bulletin for the History of Chemistry, 42(2):147, ???? 2017. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v42-2/v42-2%
20p147.pdf.

Anonymous:2017:BRa

[601] Anonymous. Book reviews. Bulletin for the History of Chemistry, 42(2): ??, ???? 2017. CODEN BHCHET. ISSN 1053-4385.

Leigh:2018:AMC

[602] G. J. Leigh. Alexander Marcet, chemist, physician and geologist, a neglected figure in British science from 1797 to 1822. Bulletin for the History of Chemistry, 43(1):1–13, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-1/v43-1%20p1-13.pdf.

Takacs:2018:WSH

[603] Laszlo Takacs. Walthère Spring and his rivalry with M. Carey Lea. Bulletin for the History of Chemistry, 43(1):14–20, ???? 2018. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v43-1/v43-1%20p14-20.pdf.

Anonymous:2018:IET

[604] Anonymous. Introduction to an English translation of Markovnikov’s first paper describing “Markovnikov’s rule”. Bulletin for the History of Chemistry, 43(1):21–23, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-1/v43-1%20p21-23.pdf.

Morkovnikov:2018:PDQ

[605] V. Morkovnikov. Primary documents: on the question of the mutal influence of atoms in chemical compounds. Bulletin for the History of Chemistry, 43(1):24–26, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-1/v43-1%20p24-26.pdf.

Griep:2018:FCT

[606] Mark A. Griep. Forgotten chemistry time capsule revealed the stories of two early female chemistry professors. Bulletin for the History of Chemistry, 43(1):27–40, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-1/v43-1%20p27-40.pdf.

Giunta:2018:BRC

[607] Carmen J. Giunta. Book review: Cradle of Chemistry. the Early Years of Chemistry at the University of Edinburgh. Bulletin for the History of Chemistry, 43(1):41–42, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-1/v43-1%20p41-42.pdf.

Anonymous:2018:IAa

[608] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 43(1):44, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL

http://acshist.scs.illinois.edu/v43-1/v43-1%20p44.pdf.

Anonymous:2018:BESa

[609] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 43(1):45, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http://

acshist.scs.illinois.edu/v43-1/v43-1%20p45.pdf.

Anonymous:2018:HOa

[610] Anonymous. HIST officers 2017. Bulletin for the History of Chemistry, 43(1):45, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http://

acshist.scs.illinois.edu/v43-1/v43-1%20p45.pdf.

Seeman:2018:PPD

[611] Jeffrey I. Seeman. Profiles, pathways and dreams: from naïveté to the HIST Award. Bulletin for the History of Chemistry, 43(2):45–60, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v43-2/v43-2%20p45-60.pdf.

Leigh:2018:SPA

[612] G. J. Leigh and Carmen J. Giunta. The scientific publications of Alexander marcet supplemental information. Bulletin for the History of Chemistry, 43(2):61–78, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-2/v43-2%20p61-78.pdf.

Anonymous:2018:FAI

[613] Anonymous. Frederick Accum: an important nineteenth-century chemist fallen into oblivion. Bulletin for the History of Chemistry, 43(2):79– 89, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v43-2/v43-2%
20p79-89.pdf.

Delegard:2018:CRM

[614] C. H. Delegard, V. F. Peretrukhin, and S. I. Rovny. The contributions of radiochemistry to mastering atomic energy for weapons. Bulletin for the History of Chemistry, 43(2):90–101, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-2/v43-2%20p90-101.pdf.

Epstein:2018:DSF

[615] Jessica Epstein. Drugs that shaped the FDA: from elixir sulfanilamide to thalidomide. Bulletin for the History of Chemistry, 43(2):102–110, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v43-2/v43-2%20p102-110.pdf.

Sahyun:2018:MSL

[616] M. R. V. Sahyun. Melville Sahyun: a life in biochemistry. Bulletin for the History of Chemistry, 43(2):111–124, ???? 2018. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v43-2/v43-2%20p111-124.pdf.

Scerri:2018:RRT

[617] Eric R. Scerri. Response to review of A Tale of Seven Scientists. Bulletin for the History of Chemistry, 43(2):125–127, ???? 2018. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v43-2/v43-2%20p125-127.pdf. See [593].

Lambert:2018:BRF

[618] Joseph B. Lambert. Book review: The Foundations of Physical Organic Chemistry: Fifty Years of the James Flack Norris Award. Bulletin for the History of Chemistry, 43(2):128–130, ???? 2018. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v43-2/v43-2%20p128-130.pdf.

Anonymous:2018:IAc

[619] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 43(2):132, ???? 2018. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v43-2/v43-2%20p132.pdf.

Anonymous:2018:BESb

[620] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 43(2):133, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v43-2/v43-2%
20p133.pdf.

Anonymous:2018:HOb

[621] Anonymous. HIST Officers 2018. Bulletin for the History of Chemistry, 43(2):133, ???? 2018. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v43-2/v43-2%
20p133.pdf.

Anonymous:2018:BRb

[622] Anonymous. Book reviews. Bulletin for the History of Chemistry, 43(2): ??, ???? 2018. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2018:BRc

[623] Anonymous. Book reviews. Bulletin for the History of Chemistry, 43(2): ??, ???? 2018. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2018:BRd

[624] Anonymous. Book reviews. Bulletin for the History of Chemistry, 43(2): ??, ???? 2018. CODEN BHCHET. ISSN 1053-4385.

Poole-Sawyer:2019:CCP

[625] Julianna Poole-Sawyer. A changing curriculum: Pharmacological texts at the University of Paris in the Twelfth and Thirteenth Centuries. Bulletin for the History of Chemistry, 44(1):1–9, ???? 2019. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v44-1/v44-1%20p1-9.pdf.

Bowden:2019:NMS

[626] Mary Ellen Bowden and Dee Ann Castel. Note: a modern scientific interpretation of Joseph Priestley’s discovery of CO. Bulletin for the History of Chemistry, 44(1):10–17, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-1/v44-1%20p10-17.pdf.

Anonymous:2019:SHC

[627] Anonymous. A survey of history of chemistry by chemists. Bulletin for the History of Chemistry, 44(1):18–31, ???? 2019. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v44-1/v44-1%20p18-31.pdf.

Raos:2019:OTB

[628] Nenad Raos. Oparin’s theory of biogenesis: Biocolloidal or biomolecular? Bulletin for the History of Chemistry, 44(1):32–36, ???? 2019. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v44-1/v44-1%20p32-36.pdf.

Burri:2019:SSR

[629] Kaspar F. Burri and Richard J. Friary. A school for synthesis: R. B. Woodward and the Woodward Research Institute remembered supplemental information. Bulletin for the History of Chemistry, 44(1):37–50, ????

2019. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v44-1/v44-1%20p37-50.pdf.

Perfetti:2019:RDS

[630] Thomas A. Perfetti. The recipients of the Dexter and Sidney M. Edelstein Awards: Biographies of men and women of the history of chemistry: an enjoyable journey through chemistry. Bulletin for the History of Chemistry, 44(1):51–61, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-1/v44-1%20p51-61.pdf.

Greenberg:2019:BRPa

[631] Arthur Greenberg. Book review: The Posthumous Nobel Prize in Chemistry. Volume 1. Bulletin for the History of Chemistry, 44(1):62–68, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v44-1/v44-1%20p62-68.pdf.

Seeman:2019:BRC

[632] Jeffrey I. Seeman. Book review: Classical Methods in Structure Elucidation of Natural Products. Bulletin for the History of Chemistry, 44(1): 68–73, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v44-1/v44-1%
20p68-73.pdf.

Anonymous:2019:YPT

[633] Anonymous. 150 years of the periodic table at Fall 2019 ACS Meeting. Bulletin for the History of Chemistry, 44(1):74, ???? 2019. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v44-1/v44-1%20p74.pdf.

Anonymous:2019:IAd

[634] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 44(1):75, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-1/v44-1%20p75.pdf.

Seeman:2019:BSH

[635] Jeffrey I. Seeman. The back story: Henry Carrington Bolton and James J. Bohning. Bulletin for the History of Chemistry, 44(1):76, 2019. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v44-1/v44-1%20p76.pdf.

Anonymous:2019:BESd

[636] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 44(1):77, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v44-1/v44-1%
20p77.pdf.

Anonymous:2019:HOd

[637] Anonymous. HIST officers 2019. Bulletin for the History of Chemistry, 44(1):77, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v44-1/v44-1%
20p77.pdf.

Anonymous:2019:BRd

[638] Anonymous. Book reviews. Bulletin for the History of Chemistry, 44(1): ??, ???? 2019. CODEN BHCHET. ISSN 1053-4385.

Lewis:2019:MYD

[639] David E. Lewis. 1860–1861: Magic years in the development of the structural theory of organic chemistry. Bulletin for the History of Chemistry, 44(2):77–91, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-2/v44-2%20p77-91.pdf.

Brooks:2019:KSC

[640] Nathan M. Brooks. The Kazan School of Chemistry: a re-interpretation. Bulletin for the History of Chemistry, 44(2):92–99, ???? 2019. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v44-2/v44-2%20p92-99.pdf.

Anonymous:2019:MMA

[641] Anonymous. Mendeleev, Meyer, and atomic volumes: an introduction to an English translation of Mendeleev’s 1869 article supplemental information. Bulletin for the History of Chemistry, 44(2):100–108, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v44-2/v44-2%20p100-108.pdf.

Girolami:2019:PDA

[642] D. I. Mendeleev Translated by Gregory S. Girolami and Vera V. Mainz. Primary documents: on the atomic volume of simple bodies. Bulletin for the History of Chemistry, 44(2):109–115, ???? 2019. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v44-2/v44-2%20p109-115.pdf.

Nicholas:2019:TTF

[643] Christopher P. Nicholas. Terpene transformations and family relations: Vladimir Ipatieff. Bulletin for the History of Chemistry, 44(2):116– 122, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v44-2/v44-2%
20p116-122.pdf.

Rasmussen:2019:EHP

[644] Seth C. Rasmussen. Early history of polyaniline — revisited: Russian contributions of Fritzsche and Zinin. Bulletin for the History of Chemistry, 44(2):123–133, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-2/v44-2%20p123-133.pdf.

Martin:2019:CWW

[645] Dean F. Martin and Marwa Elkharsity. Chemist at war: World war ii roles of jonas kamlet, consulting chemist. Bulletin for the History of Chemistry, 44(2):134–138, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-2/v44-2%20p134-138.pdf.

Greenberg:2019:BRPb

[646] Arthur Greenberg. Book review: The Posthumous Nobel Prize in Chemistry. Volume 2. Bulletin for the History of Chemistry, 44(2):139–147, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v44-2/v44-2%20p139-147.pdf.

Strom:2019:AAW

[647] E. Thomas Strom. African American Women Chemists in the Modern Era. Bulletin for the History of Chemistry, 44(2):147–148, ???? 2019. CO-

DEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.
edu/bulletin_open_access/v44-2/v44-2%20p147-148.pdf.

Hendrickson:2019:BRL

[648] Connie Hendrickson. Book review: A Lab of One’s Own: Science and Suffrage in the First World War. Bulletin for the History of Chemistry, 44(2):149, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-2/v44-2%20p149.pdf.

Karol:2019:BRT

[649] Paul J. Karol. Book review: From Transuranic to Superheavy Elements: a Story of Dispute and Creation. Bulletin for the History of Chemistry, 44(2):150–151, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-2/v44-2%20p150-151.pdf.

Anonymous:2019:IAe

[650] Anonymous. Instructions for authors. Bulletin for the History of Chemistry, 44(2):151, ???? 2019. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v44-2/v44-2%20p151.pdf.

Seeman:2019:BSK

[651] Jeffrey I. Seeman. The back story Koji Nakanishi (may 11, 1925–march 28, 2019), magician supreme. Bulletin for the History of Chemistry, 44(2):152, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.

scs.illinois.edu/bulletin_open_access/v44-2/v44-2%20p152.pdf.

Anonymous:2019:BESe

[652] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 44(2):153, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v44-2/v44-2%
20p153.pdf.

Anonymous:2019:HOe

[653] Anonymous. HIST officers 2019. Bulletin for the History of Chemistry, 44(2):153, ???? 2019. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v44-2/v44-2%
20p153.pdf.

Anonymous:2019:BRe

[654] Anonymous. Book reviews. Bulletin for the History of Chemistry, 44(2): ??, ???? 2019. CODEN BHCHET. ISSN 1053-4385.

Hoffmann:2022:F

[655] Roald Hoffmann. Foreword. Bulletin for the History of Chemistry, 47 (1):1–2, ???? 2022. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v47-1/v47-1%
20p1-2.pdf.

Anonymous:2022:PHC

[656] Anonymous. Preface: Happy centennial to the ACS Division of the History of Chemistry. Bulletin for the History of Chemistry, 47(1):3–4, ???? 2022. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v47-1/v47-1%20p3-4.pdf.

Anderson:2022:ECC

[657] Robert G. W. Anderson. Epitomizing chemistry for changing audiences in Britain, 1820–2020 supplemental information. Bulletin for the History of Chemistry, 47(1):5–14, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p5-14.pdf.

Greenberg:2022:MPM

[658] Arthur Greenberg. Mendeleev’s “Problems:” a means to engage students and teachers in the history of chemistry. Bulletin for the History of Chemistry, 47(1):15–28, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p15-28.pdf.

Orna:2022:ACP

[659] Mary Virginia Orna. Archaeological chemistry: Past, present, future. Bulletin for the History of Chemistry, 47(1):29–41, ???? 2022. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v47-1/v47-1%20p29-41.pdf.

Cole:2022:CWB

[660] David Allen Cole. Can we bring chemistry back? Exploring the potential of “Gateway artifacts” at the Science History Institute. Bulletin for the History of Chemistry, 47(1):43–49, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p43-49.pdf.

Travis:2022:HCI

[661] Anthony S. Travis. Historiography of the chemical industry: Technologies and products versus corporate history. Bulletin for the History of Chemistry, 47(1):50–61, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p50-61.pdf.

Lewis:2022:FHS

[662] David E. Lewis. A future history of selectivity in organic chemistry: Whence, where, and whither? Bulletin for the History of Chemistry, 47(1):62–76, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p62-76.pdf.

Parascandola:2022:DMC

[663] John Parascandola. The development of medicinal chemistry as a disciplines: a topic ripe for historical exploration. Bulletin for the History of Chemistry, 47(1):77–84, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p77-84.pdf.

Rasmussen:2022:MBI

[664] Seth C. Rasmussen. Moving beyond the intersection of chemistry and history: Evolving multidisciplinary approaches to the historical study of chemistry. Bulletin for the History of Chemistry, 47(1):85–90, ???? 2022. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v47-1/v47-1%20p85-90.pdf.

Restrepo:2022:CHC

[665] Guillermo Restrepo. Computational history of chemistry. Bulletin for the History of Chemistry, 47(1):91–106, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p91-106.pdf.

Collins:2022:HCT

[666] Sibrina N. Collins. History of chemistry as a tool for the engagement of underrepresented students in chemistry. Bulletin for the History of Chemistry, 47(1):107–110, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p107-110.pdf.

Rayner-Canham:2022:OCF

[667] Marelene F. Rayner-Canham and Geoffrey W. Rayner-Canham. Out of obscurity: Contextualizing forgotten women chemists. Bulletin for the History of Chemistry, 47(1):111–118, ???? 2022. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v47-1/v47-1%20p111-118.pdf.

Weininger:2022:PSC

[668] Stephen J. Weininger. “The poor sister:” coming to grips with recent and contemporary chemistry. Bulletin for the History of Chemistry, 47(1):119–123, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p119-123.pdf.

Morris:2022:IPM

[669] Peter J. T. Morris and Jeffrey I. Seeman. The importance of plurality and mutual respect in the practice of the history of chemistry. Bulletin for the History of Chemistry, 47(1):124–137, ???? 2022. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v47-1/v47-1%20p124-137.pdf.

Anonymous:2022:LSI

[670] Anonymous. The long and short of it: the future writing of history of chemistry. Bulletin for the History of Chemistry, 47(1):138–145, ???? 2022. CODEN BHCHET. ISSN 1053-4385. URL http://acshist.scs.

illinois.edu/bulletin_open_access/v47-1/v47-1%20p138-145.pdf.

Jensen:2022:DHC

[671] William B. Jensen. Does history of chemistry have a future? Bulletin for the History of Chemistry, 47(1):146–152, ???? 2022. CODEN BHCHET.

ISSN 1053-4385. URL http://acshist.scs.illinois.edu/bulletin_
open_access/v47-1/v47-1%20p146-152.pdf.

Seeman:2022:RIH

[672] Jeffrey I. Seeman. Remote interviewing and the history of chemistry. Bulletin for the History of Chemistry, 47(1):153–162, ???? 2022. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v47-1/v47-1%20p153-162.pdf.

Anonymous:2022:TRP

[673] Anonymous. Is there room for the present in the history of science? Bulletin for the History of Chemistry, 47(1):163–170, ???? 2022. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v47-1/v47-1%20p163-170.pdf.

Rocke:2022:RLN

[674] Alan J. Rocke. Reflections on the last and the next hundred years. Bulletin for the History of Chemistry, 47(1):171–175, ???? 2022. CODEN

BHCHET. ISSN 1053-4385. URL http://acshist.scs.illinois.edu/
bulletin_open_access/v47-1/v47-1%20p171-175.pdf.

Seeman:2022:BSS

[675] Jeffrey I. Seeman. The back story: Steve Weininger, comfortable in many disciplines supplemental information. Bulletin for the History of Chemistry, 47(1):176, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20p176.pdf.

Anonymous:2022:BESf

[676] Anonymous. Bulletin editorial staff. Bulletin for the History of Chemistry, 47(1):177, ???? 2022. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v47-1/v47-1%
20p177.pdf.

Anonymous:2022:HOf

[677] Anonymous. HIST officers 2022. Bulletin for the History of Chemistry, 47(1):177, ???? 2022. CODEN BHCHET. ISSN 1053-4385. URL http:

//acshist.scs.illinois.edu/bulletin_open_access/v47-1/v47-1%
20p177.pdf.

Anonymous:2022:CKC

[678] Anonymous. Cover, key and credits. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN

1053-4385. URL http://acshist.scs.illinois.edu/bulletin_open_
access/v47-1/v47-1%20cover%20key%20credits.pdf.

Anonymous:2022:EAH

[679] Anonymous. Expansive approaches to the history of chemistry. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2022:FST

[680] Anonymous. Foci on specific topics. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2022:MAT

[681] Anonymous. Multidisciplinary approaches and tools. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2022:PPF

[682] Anonymous. The past, present and future of history of chemistry. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2022:RHC

[683] Anonymous. Relationships of historians and chemist-historians. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN 1053-4385.

Anonymous:2022:TID

[684] Anonymous. Thriving, inclusivity, diversity, and equity and the history of chemistry. Bulletin for the History of Chemistry, 47(1):??, ???? 2022. CODEN BHCHET. ISSN 1053-4385.

Abbri:1991:ADI

[685] F. Abbri and F. Crispini. Atti del III convegno nazionale di storia e fondamenti della chimica. Edizione Brenner, Cosenza, Italia, 1991. 402 pp. LCCN ????

Aftalion:1991:HIC

[686] Fred Aftalion. A history of the international chemical industry. The Chemical sciences in society series. University of Pennsylvania Press, Philadelphia, 1991. ISBN 0-8122-8207-8 (hardcover), 0-8122-1297-5 (paper). xix + 411 pp. LCCN HD9650.5 .A3813 1991.

Aldersey-Williams:2011:PTCa

[687] Hugh Aldersey-Williams. Periodic tales: a cultural history of the elements, from arsenic to zinc. Ecco, New York, NY, USA, 2011. ISBN 0-06-182472-0. xvii + 428 pp. LCCN QD467.

Barton:1991:SRG

[688] Sir Derek Barton. Some recollections of gap jumping. Profiles, pathways, and dreams. American Chemical Society, Washington, DC, USA, 1991. ISBN 0-8412-1770-X, 0-8412-1796-3 (paperback). ISSN 1047-8329. xxiv + 143 pp. LCCN QD22.B257 A3 1990. URL http://www.

loc.gov/catdir/enhancements/fy0604/90000872-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0604/90000872-t.html.

Benninga:1990:HLA

[689] H. (Harm) Benninga. A history of lactic acid making: a chapter in the history of biotechnology. Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0625-2 (U.S.). xxi + 478 pp. LCCN TP248.L3 B46 1990.

Bensaude-Vincent:1996:HC

[690] Bernadette Bensaude-Vincent and Isabelle Stengers. A History of Chemistry. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-39659-6. 305 pp. LCCN QD11 .B4413 1996. Translation by Deborah van Dam of Histoire de la chemie, Editions La Découverte, 1993.

Blum:2011:PHMb

[691] Deborah Blum. The poisoner’s handbook: murder and the birth of forensic medicine in Jazz Age New York. Penguin Group USA, New York, NY, USA, 2011. ISBN 1-101-52488-X (e-book), 1-101-52489-8 (e-book), 0-14-311882-X. LCCN HV6555.U62 N373 2011eb. URL

http://excerpts.cdn.overdrive.com/FormatType-410/1523-1/C8E/
0FB/2B/PoisonersHandbook9781101524893.epub;   http://images.
contentreserve.com/ImageType-100/1523-1/{C8E0FB2B-3755-4C80-
B4FA-7C70876CA705}Img100.jpg.

Brock:1997:JLC

[692] W. H. (William Hodson) Brock. Justus von Liebig: the chemical gatekeeper. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56224-4 (hardcover). xiv + 374 pp. LCCN QD22.L7 B76 1997.

URL http://www.loc.gov/catdir/description/cam027/96025468.
html;   http://www.loc.gov/catdir/samples/cam031/96025468.html;
 http://www.loc.gov/catdir/toc/cam031/96025468.html.

Brock:2011:CPS

[693] William H. Brock. The Case of the Poisonous Socks: Tales from Chemistry. Royal Society of Chemistry Publishing, Cambridge, UK, 2011. ISBN 1-84973-324-4 (paperback). xiv + 348 pp. LCCN ????

Brown:2012:AAW

[694] Jeannette E. (Jeannette Elizabeth) Brown. African American women chemists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-974288-X. xii + 252 pp. LCCN QD21

.B69 2012. URL http://www.loc.gov/catdir/enhancements/fy1205/
2011013662-b.html;   http://www.loc.gov/catdir/enhancements/
fy1205/2011013662-d.html;   http://www.loc.gov/catdir/enhancements/
fy1205/2011013662-t.html.

Bud:1998:ISH

[695] Robert Bud and Deborah Jean Warner, editors. Instruments of science: an historical encyclopedia, volume 2 (GEHS), 936 (GRLSS) of Garland encyclopedias in the history of science; Garland reference library of social science. Science Museum, London, and National Museum of American History, Smithsonian Institution, in association with Garland Publishing, London, UK and New York, NY, USA, 1998. ISBN 0-8153-1561-9. xxv + 709 pp. LCCN Q184.5 .I57 1998. URL http://www.loc.gov/catdir/

enhancements/fy0652/97015296-d.html.

Crawford:1996:AIT

[696] Elisabeth T. Crawford. Arrhenius: from ionic theory to the greenhouse effect, volume 23 of Uppsala studies in history of science. Science History Publications/USA, Canton, MA, USA, 1996. ISBN 0-88135-166-0. xiii + 320 + 9 pp. LCCN Q143.A77 C68 1996.

Davies:2013:UCD

[697] Alwyn George Davies and Peter J. (Peter Joseph) Garratt. UCL Chemistry Department: 1828–1974. Science Reviews 2000, St. Albans, Hertfordshire, UK, 2013. ISBN 1-900814-46-3 (paperback). vii + 287 pp. LCCN LF425 .D38 2013.

deMenten:2013:DCA

[698] Pierre de Menten and Brigitte Van Tiggelen. Dictionnaire de chimie: une approche étymologique et historique. (French) [Dictionary of Chemistry: An Etymological and Historical Approach]. De Boeck, Bruxelles, Belgium, 2013. ISBN 2-8041-8175-8 (paperback). 395 pp. LCCN QD5 .M46 2013.

Djerassi:1990:SMI

[699] Carl Djerassi. Steroids made it possible. Profiles, pathways, and dreams. American Chemical Society, Washington, DC, USA, 1990. ISBN 0-8412-1773-4, 0-8412-1799-8 (paperback). ISSN 1047-8329. xxiv + 205 pp. LCCN QD22.D63 A3 1990. URL http://www.

loc.gov/catdir/enhancements/fy0604/90000906-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0604/90000906-t.html.

Faraday:1861:CSL

[700] Michael Faraday and William Crookes, editors. A course of six lectures on the chemical history of a candle: to which is added a lecture on platinum. Griffin, Bohn and Co., London, UK, 1861. viii + 208 pp. LCCN QD39 .F2x 1861.

Faraday:1874:CHC

[701] Michael Faraday and William Crookes. The chemical history of a candle, a course of lectures delivered before a juvenile audience at the Royal Institution. Chatto and Windus, London, UK, 1874. 226 pp. LCCN QD39 .F2 1874.

Faraday:2011:CHC

[702] Michael Faraday and Frank A. J. L. James. The chemical history of a candle. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, sesquicentenary edition, 2011. ISBN 0-19-969491-5 (hardcover). xliv + 152 pp. LCCN QD39 .F2 2011.

Farrar:1997:CCI

[703] Wilfred Vernon Farrar, W. H. (William Hodson) Brock, and Richard Leslie Hills. Chemistry and the chemical industry in the 19th century: the Henrys of Manchester and other studies, volume 575 of Variorum collected studies series. Variorum, Aldershot, UK, 1997. ISBN 0-86078-630-7 (hardcover). 296 pp. LCCN QD18.E85 F37 1997.

Firmenich:1990:CNF

[704] Roger Firmenich. From Chuit and Naef to Firmenich S.A. [publisher not identified], [Switzerland], 1990. 139 pp. LCCN HD9999.P3934 F54713 1990. Translation by Clive Norris of De Chuit and Naef à Firmenich SA.

Fisher:2010:MAA

[705] David E. Fisher. Much ado about (practically) nothing: a history of the noble gases. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-539396-1 (hardcover). x + 264 pp. LCCN QD162 .F57 2010.

Fruton:1990:CSS

[706] Joseph S. (Joseph Stewart) Fruton. Contrasts in scientific style: research groups in the chemical and biochemical sciences, volume 191 of Memoirs series / American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1990. ISBN 0-87169-191-4. ISSN 0065-9738. ix + 473 pp. LCCN Q11 .P612 vol. 191 QP511.

Ganz:2010:NR

[707] Adam Ganz. Nuclear reactions. 45-minute radio drama based on the Farm Hall Transcripts. Performed on the BBC Radio Afternoon The-

ater., June 15, 2010. URL http://www.rhul.ac.uk/mediaarts/news/
nuclearreactions.aspx.

Gavroglou:2012:NPN

[708] Kostas Gavroglou and Ana Simões. Neither physics nor chemistry: a history of quantum chemistry. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01618-4 (hardcover). xiv + 351 pp. LCCN QD462 .G38 2012.

Giunta:2010:ACD

[709] Carmen Giunta, editor. Atoms in chemistry: from Dalton’s predecessors to complex atoms and beyond, volume 1044 of ACS symposium series. American Chemical Society, Washington, DC, USA, 2010. ISBN 0-8412-2557-5. vii + 116 pp. LCCN QD461 .A863 2010.

Green:2012:OCP

[710] Mark M. Green. Organic chemistry principles in context: a story telling historical approach. ScienceFromAway Publishing, New York, NY, USA, 2012. ISBN 0-615-70271-6. xxiv + 452 pp. LCCN QD251.3 .G742 2012.

Griep:2009:RCM

[711] Mark Griep and Marjorie Mikasen. Reaction!: chemistry in the movies. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-532692-X (hardcover). viii + 340 pp. LCCN PN1995.9.S265 G75 2009.

GuytondeMorveau:1994:MNC

[712] Louis-Bernard Guyton de Morveau, Bernadette Bensaude-Vincent, et al. Méthode de Nomenclature Chimique. Editions du Seuil, Paris, France, 1994. ISBN 2-02-019253-5. 252 pp. LCCN QD7 .M44 1994. Le présent ouvrage comprend le texte de la Méthode de nomenclature chimique dans sa version originale publiée en 1787 par la librairie Cuchet.

Habashi:1994:HM

[713] Fathi Habashi, editor. A history of metallurgy. Métallurgie Extractive Québec, Sainte-Foy, Québec, Canada, 1994. ISBN 2-9803247-1-X. vi + 322 pp. LCCN TN615 .H47 1994.

Hamlin:1990:SIW

[714] Christopher Hamlin. A science of impurity: water analysis in nineteenth century Britain. University of California Press, Berkeley, CA, USA, 1990. ISBN 0-520-07088-7. xiii + 342 pp. LCCN TD257 .H35

1990. URL http://ark.cdlib.org/ark:/13030/ft667nb43t;   http:
//www.loc.gov/catdir/bios/ucal051/90010827.html.

Hermes:1996:EOL

[715] Matthew E. Hermes. Enough for one lifetime: Wallace Carothers, inventor of nylon. History of modern chemical sciences. American Chemical Society and the Chemical Heritage Foundation, Washington, DC, USA, 1996. ISBN 0-8412-3331-4. ISSN 1069-2452. xvii + 345 pp. LCCN QD22.C35

H47 1996. URL http://www.loc.gov/catdir/enhancements/fy0630/
96005560-d.html.

Hollander:1996:LL

[716] Leonore Hollander. A Life as Lived. Raspberry Press, Sebastopol, CA, USA, 1996. xii + 445 pp. LCCN ????

Hunter:2009:BBG

[717] Michael Hunter. Boyle: Between God and Science. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-12381-7 (hardcover). xii + 366 pp. LCCN Q143.B69 H86 2009.

Ihde:1990:CVB

[718] Aaron J. Ihde. Chemistry, as viewed from Bascom’s Hill: a history of the Chemistry Department at the University of Wisconsin at Madison. Department of Chemistry, University of Wisconsin-Madison, Madison, WI, USA, 1990. xvi + 688 pp. LCCN QD47.5.W63 U54 1990.

Johnston:2007:MCB

[719] John Johnston. A Manual of Chemistry, on the Basis of Turner’s Elements of chemistry. Gardners Books, 2007. ISBN 1-4255-5894-1. LCCN ????

Knight:1992:ICH

[720] David M. Knight. Ideas in chemistry: a history of the science. Rutgers University Press, New Brunswick, NJ, USA, 1992. ISBN 0-8135-1835-0 (hardcover), 0-8135-1836-9 (paperback). vi + 213 pp. LCCN QD11 .K53 1992.

Knight:1998:MCS

[721] David M. Knight and Helge Kragh, editors. The making of the chemist: the social history of chemistry in Europe, 1789–1914. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58351-9 (hardcover). xxi + 353 pp. LCCN QD39.7 .M34 1998. URL http://

www.loc.gov/catdir/description/cam028/97043009.html;   http:
//www.loc.gov/catdir/toc/cam021/97043009.html.

Kubanek:2010:NLT

[722] Anne-Marie Weidler Kubanek. Nothing less than an adventure: Ellen Gleditsch and her life in science. Crossfield Publishing, St. Marys, ON, Canada, 2010. ISBN 1-4528-4213-2. iv + 185 pp. LCCN ????

Lambert:1997:TPU

[723] Joseph B. Lambert. Traces of the past: unraveling the secrets of archaeology through chemistry. Helix books. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-40928-3. xiii + 319 pp. LCCN CC75

.L297 1997. URL http://www.loc.gov/catdir/enhancements/fy0830/
97011454-d.html.

Lambert:2014:CL

[724] Joseph B. Lambert. A Chemical Life. De Rigueur Press, North Manchester, IN, USA, 2014. ISBN 0-9916503-0-1. 388 pp. LCCN QD22.L36 L36 2014.

Lesch:2007:FMD

[725] John E. Lesch. The first miracle drugs: how the sulfa drugs transformed medicine. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-518775-X (hardcover). x + 364 pp. LCCN RM666.S88 L47 2007. URL http://www.loc.gov/catdir/

enhancements/fy0726/2005036136-d.html;   http://www.loc.gov/
catdir/toc/ecip065/2005036136.html.

Loschmidt:1913:KFO

[726] J. (Johann Josef) Loschmidt and Richard Anschütz. Konstitutionsformeln der organischen chemie in graphischer darstellung. (German) [Constitutional formulas of organic chemistry in graphical representation], volume 190 of Ostwalds Klassiker der exakten Wissenschaften. W. Engelmann, Leipzig, Germany, 1913. 154 pp. LCCN QD471 .L7.

Malley:2011:RHM

[727] Marjorie Caroline Malley. Radioactivity: a history of a mysterious science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-976641-X (hardcover). xxi + 267 pp. LCCN QC794.6.R3 M35 2011.

McEvoy:2010:HCR

[728] John G. McEvoy. The historiography of the chemical revolution: patterns of interpretation in the history of science. Pickering and Chatto, London, UK, 2010. ISBN 1-84893-030-5 (hardcover), 1-84893-031-3 (e-book). xiii + 328 pp. LCCN QD15 .M34 2010. URL http://catalogue.bnf.fr/

ark:/12148/cb424066387.

Mierzecki:1991:HDC

[729] Roman Mierzecki. The historical development of chemical concepts, volume 12 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-0915-4. xi + 281 pp. LCCN QD11 .M6613 1990.

Morris:1989:ASR

[730] Peter John Turnbull Morris. The American synthetic rubber research program. The chemical sciences in society series. University of Pennsylvania Press, Philadelphia, 1989. ISBN 0-8122-8205-1. xii + 191 pp. LCCN TS1925 .M58 1989.

Morris:2015:MFH

[731] Peter J. T. Morris. The Matter Factory: a History of the Chemistry Laboratory. Reaktion Books, London, UK, 2015. ISBN 1-78023-442-2. 416 pp. LCCN QD51 .M68 2015.

Mulliken:1989:LSA

[732] Robert S. Mulliken and Bernard J. Ransil. Life of a scientist: an autobiographical account of the development of molecular orbital theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-50375-7, 0-387-50375-7. xv + 256 pp. LCCN ????

Nye:2011:MPH

[733] Mary Jo Nye. Michael Polanyi and his generation: origins of the social construction of science. University of Chicago Press, Chicago, IL, USA, 2011. ISBN 0-226-61063-2 (hardcover). xxi + 405 pp. LCCN B945.P584 N94 2011.

Pfaundler:1873:KUD

[734] L. Pfaundler. Der “Kampf um’s Dasein” unter den Molekülen: ein weiterer Beitrag zur chemischen Statik. (German) [The “struggle for existence” among the molecules: Another contribution to the chemical statics]. Barth, Leipzig, Germany, 1873. 182–198. pp. LCCN ????

Prelog:1991:MSC

[735] V. Prelog. My 132 semesters of chemistry studies: studium chymiae nec nisi cum morte finitur. Profiles, pathways, and dreams. American Chemical Society, Washington, DC, USA, 1991. ISBN 0-8412-1772-6 (hardcover), 0-8412-1798-X (paperback). ISSN 1047-8329. xxiv + 120 pp. LCCN QD22.P73 A3 1991. URL http://www.

loc.gov/catdir/enhancements/fy0604/90000951-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0604/90000951-t.html.

Principe:1998:AAR

[736] Lawrence Principe. The aspiring adept: Robert Boyle and his alchemical quest: including Boyle’s “lost” Dialogue on the transmutation of metals. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-01678-X (hardcover). xiv + 339 pp. LCCN QD24.B685

P75 1998. URL http://www.h-net.org/review/hrev-a0a8s4-aa;
http://www.loc.gov/catdir/description/prin021/97041793.html;
 http://www.loc.gov/catdir/toc/prin031/97041793.html.

Principe:2013:SA

[737] Lawrence Principe. The secrets of alchemy. Synthesis. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-68295-1 (hardcover), 0-226-92378-9 (e-book). v + 281 + 8 pp. LCCN QD13 .P75 2013.

Rayner-Canham:1998:WCT

[738] Marelene F. Rayner-Canham and Geoffrey Rayner-Canham. Women in chemistry: their changing roles from alchemical times to the mid-twentieth century. History of modern chemical sciences. American Chemical Society, Washington, DC, USA, 1998. ISBN 0-8412-3522-8. ISSN 1069-2452. xiv + 284 pp. LCCN QD20 .R39 1998.

Redniss:2010:RMP

[739] Lauren Redniss. Radioactive: Marie and Pierre Curie, a tale of love and fallout. IT Books/Harper Collins, New York, NY, USA, 2010. ISBN 0-06-135132-6. 205 pp. LCCN QD22.C8

R395 2010. URL http://www.loc.gov/catdir/enhancements/fy1111/
2011281920-b.html;   http://www.loc.gov/catdir/enhancements/
fy1111/2011281920-d.html.

Redniss:2015:RMP

[740] Lauren Redniss. Radioactive: Marie and Pierre Curie, a tale of love and fallout. Dey Street Books, an imprint of HarperCollins Publishers, London, UK, 2015. ISBN 0-06-241616-2 (paperback). 205 pp. LCCN ????

Rocke:1993:QRH

[741] Alan J. Rocke. The quiet revolution: Hermann Kolbe and the science of organic chemistry. California studies in the history of science. University of California Press, Berkeley, CA, USA, 1993. ISBN 0-520-08110-2 (hardcover). xiii + 501 pp. LCCN QD22.K578 R63 1993. URL http:

//ark.cdlib.org/ark:/13030/ft5g500723;   http://www.loc.gov/
catdir/bios/ucal051/92028190.html;   http://www.loc.gov/catdir/
description/ucal041/92028190.html.

Ross:1991:NCA

[742] Sydney Ross. Nineteenth-century attitudes: men of science, volume 13 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 0-7923-1308-9 (HB). xi + 235 pp. LCCN Q174.8 .R67 1991.

Russell:1996:EFC

[743] Colin Archibald Russell. Edward Frankland: chemistry, controversy, and conspiracy in Victorian England. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49636-5 (hardcover). xx + 535 pp. LCCN

QD22.F65 R86 1996. URL http://www.loc.gov/catdir/description/
cam027/95040319.html;   http://www.loc.gov/catdir/toc/cam021/
95040319.html.

Salzberg:1991:CCC

[744] Hugh W. Salzberg. From caveman to chemist: circumstances and achievements. American Chemical Society, Washington, DC, USA, 1991. ISBN 0-8412-1786-6 (hardcover), 0-8412-1787-4 (paper). xix + 294 pp. LCCN

QD11 .S23 1990. URL http://www.loc.gov/catdir/enhancements/
fy0604/90044612-d.html;   http://www.loc.gov/catdir/enhancements/
fy0604/90044612-t.html.

Scarborough:2010:PDL

[745] John Scarborough. Pharmacy and drug lore in antiquity: Greece, Rome, Byzantium, volume CS904 of Variorum collected studies series. Ashgate Publishers, Farnham, Surrey, UK, 2010. ISBN 0-7546-5954-2 (hardcover). ???? pp. LCCN RS63 .S28 2010.

Scerri:2013:TE

[746] Eric R. Scerri. A Tale of 7 Elements. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-539131-4 (hardcover). xxxiii + 270 pp. LCCN QD467 .S3475 2013. URL http://catalogue.

bnf.fr/ark:/12148/cb436625450. Preface by Oliver Sacks.

Schutt:1997:EMP

[747] Hans-Werner Schütt. Eilhard Mitscherlich, prince of Prussian chemistry. History of modern chemical sciences. American Chemical Society and the Chemical Heritage Foundation, Washington, DC, USA, 1997. ISBN 0-8412-3345-4. xvi + 239 pp. LCCN QD22.M58 S3813 1997. URL http:/

/www.loc.gov/catdir/enhancements/fy0630/96033322-d.html.

Schwartz:1990:MTP

[748] A. Truman (Albert Truman) Schwartz and John G. McEvoy, editors. Motion toward perfection: the achievement of Joseph Priestley. Skinner House Books, Boston, MA, USA, 1990. ISBN 1-55896-010-4. xxvi + 277 pp. LCCN B1383.P75 M67 1990.

Schweitzer:1988:CUH

[749] George Keene Schweitzer. Chemistry at UTK: a history of chemistry at the University of Tennessee, Knoxville from 1794 through 1987. University of Tennessee, Knoxville, Dept. of Chemistry, Knoxville, TN, USA, 1988. ISBN 0-913857-49-1. 193 pp. LCCN LD5299.5.C45 S84 1988.

Smyth:1998:JDB

[750] A. L. (Albert Leslie) Smyth. John Dalton, 1776–1844: a Bibliography of Works by and About Him, with an Annotated List of his Surviving Apparatus and Personal Effects. Manchester University Press, Manchester, UK, 1998. ISBN 1-85928-438-8. xxii + 167 pp. LCCN Z8212.7 .S5 1998 QD22.D2.

Solovyov:1990:SA

[751] Yu. I. Solovyov. Svante Arrhenius: 1859–1927. Nauka, Moscow, Russia, 1990. ISBN 5-02-001467-2. 319 pp. LCCN ????

Spitz:1988:PRI

[752] Peter H. Spitz. Petrochemicals: the rise of an industry. John Wiley, New York, NY, USA, 1988. ISBN 0-471-85985-0. xxvii + 588 pp.

LCCN HD9579.C32 S65 1988. URL http://catdir.loc.gov/catdir/
description/wiley032/87023019.html;   http://catdir.loc.gov/
catdir/toc/onix02/87023019.html;   http://www.gbv.de/dms/bowker/
toc/9780471859857.pdf.

Stocker:2013:FCE

[753] Jack H. Stocker and Natalie Foster, editors. A festival of chemistry entertainments, volume 1153 of ACS symposium series. American Chemical Society, Washington, DC, USA, 2013. ISBN 0-8412-2907-4 (hardcover). xiii + 118 pp. LCCN QD37 .F39 2013.

Strom:2012:PQC

[754] E. Thomas Strom and Angela K. Wilson, editors. Pioneers of Quantum Chemistry, volume 1122 of ACS symposium series. American Chemical Society, Washington, DC, USA, 2012. ISBN 0-8412-2716-0 (hardcover), 0-8412-2717-9. 330 pp. LCCN QD462.A1 P56 2012.

Thomas:1990:LSL

[755] J. M. (John Meurig) Thomas and D. (David) Phillips, editors. The Legacy of Sir Lawrence Bragg: selections and reflections. Science Reviews Ltd., Northwood, Middlesex, UK, 1990. ISBN 0-905927-43-5. x + 308 pp. LCCN QC16.B66 L44 1990.

Thomas:1991:MFR

[756] J. M. (John Meurig) Thomas. Michael Faraday and the Royal Institution: the genius of man and place. A. Hilger, Bristol, UK, 1991. ISBN 0-7503-0145-7. xii + 234 pp. LCCN QC16.F2 T38 1991. URL http://www.loc.

gov/catdir/enhancements/fy0666/91017248-d.html.

Tylecote:1976:HM

[757] R. F. Tylecote. A history of metallurgy. Metals Society, London, UK, 1976. ISBN 0-904357-06-6. ix + 182 pp. LCCN TN615 .T94.

Vijvers:2013:ABC

[758] Willem G. Vijvers. Alexander Borodin: composer, scientist, educator: a biography. American Book Center, Amsterdam, The Netherlands, 2013. ISBN 90-812269-0-8. 365 pp. LCCN ????

Watanabe:1990:JWS

[759] Masao Watanabe. The Japanese and Western science. University of Pennsylvania Press, Philadelphia, 1990. ISBN 0-8122-8252-3. xiv + 141 pp. LCCN Q127.J3 W3713 1990. Translated by Otto Theodor Benfey, with a foreword by Edwin O. Reischauer and a new epilogue by the author.

Wayman:1989:AGR

[760] Michael L. (Michael Lash) Wayman, editor. All that glitters: readings in historical metallurgy. Metallurgical Society of the Canadian Institute of Mining and Metallurgy, Montréal, Québec, Canada, 1989. ISBN 0-919086-24-1. x + 197 pp. LCCN TN675 .A44 1989.

Williams:1990:RRC

[761] Trevor Illtyd Williams. Robert Robinson, chemist extraordinary. Clarendon Press, New York, NY, USA, 1990. ISBN 0-19-858180-7. viii + 201 + 7 pp. LCCN QD22.R63 W54 1990. URL http://www.

loc.gov/catdir/enhancements/fy0635/89022123-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0635/89022123-t.html.

Woodward:2010:PAJ

[762] Walter William Woodward. Prospero’s America: John Winthrop, Jr., alchemy, and the creation of New England culture, 1606–1676. Published for the Omohundro Institute of Early American History and Culture by the University of North Carolina Press, Chapel Hill, NC, USA, 2010. ISBN 0-8078-3301-0 (hardcover). viii + 317 pp. LCCN QD24.W56 W66 2010.

Evans:1996:EHR

[763] C. H. Evans, editor. Episodes from the History of the Rare Earth Elements, volume 15 of Chemists and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4101-5. xxv + 240 pp. LCCN QD172.R2 E65 1996. URL http://www.

loc.gov/catdir/enhancements/fy1006/96210275-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy1006/96210275-t.html.

Klein:2010:MEE

[764] Ursula Klein and E. C. (Emma C.) Spary, editors. Materials and expertise in early modern Europe: between market and laboratory. University of Chicago Press, Chicago, IL, USA, 2010. ISBN 0-226-43968-2. LCCN TA403.6 .M369 2010.

Lagerkvist:2012:PTM

[765] Ulf Lagerkvist, editor. The periodic table and a missed Nobel Prize. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2012. ISBN 981-4295-95-7 (paperback), 981-4295-96-5 (e-book). xii + 122 pp. LCCN AS911.N9 .L34 2012; QD11 .L34 2012; QD467

.L34 2012. URL http://www.worldscientific.com/worldscibooks/

10.1142/7658. Edited by Erling Norrby.

Mauskopf:1993:CSM

[766] Seymour H. Mauskopf, editor. Chemical sciences in the modern world, The Chemical sciences in society series. University of Pennsylvania Press, Philadelphia, 1993. ISBN 0-8122-3156-2. LCCN QD11 .C48 1993.

Miles:1976:ACC

[767] Wyndham D. Miles and Robert F. Gould, editors. American chemists and chemical engineers. American Chemical Society, Washington, DC, USA, 1976, 1994. ISBN 0-8412-0278-8. x + 544 (vol. 1), ix + 365 (vol. 2) pp. LCCN QD21 .A43.

Strom:2011:YPH

[768] E. Thomas Strom and Seth C. Rasmussen, editors. 100+ years of plastics: Leo Baekeland and beyond, volume 1080 of ACS symposium series. American Chemical Society, Washington, DC, USA, 2011. ISBN 0-8412-2678-4. LCCN ????

Traynham:1987:EHO

[769] James G. Traynham, editor. Essays on the history of organic chemistry. Louisiana State University Press, Baton Rouge, LA, USA, 1987. ISBN 0-8071-1293-3. LCCN QD248 .E87 1987.

Zimmermann:1995:PWR

[770] Volker Zimmermann, editor. Paracelsus: das Werk, die Rezeption: Beiträge des Symposiums zum 500. Geburtstag von Theophrastus Bombastus von Hohenheim, genannt Paracelsus (1493-1541) an der Universität Basel am 3. and 4. Dezember 1993. (German) [Paracelsus: the work, the reception: contributions to the Symposium on the 500th birthday of Theophrastus Bombastus von Hohenheim, known as Paracelsus (1493– 1541) at the University of Basel on 3 and 4 December 1993]. Franz Steiner Verlag, Stuttgart, Germany, 1995. ISBN 3-515-06704-3. LCCN R147.P2 P38 1995.